A Complete Bibliography of Publications in Fish Physiology and Biochemistry: 2000–2009

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

20 May 2025 Version 1.09

Title word cross-reference

+3− [86]. 3n − 3 [178]. 4n − 3 [197]. 4n − 6 [197]. ^{∗} [118]. ^{+} [100]. ^{+2} [491]. +3 [491]. ^{−} [86]. ^{14} [304, 178]. ^{2+} [474, 473]. _{1} [288]. _{2} [417, 360, 367]. _{2α} [197]. _{3} [273]. _{3}^{−} [86]. _{4} [457]. _{o} [288]. α [131, 160, 291, 262, 409, 546, 246, 421, 355, 8, 396]. α_{1} [540]. β [128, 529, 258, 234, 147, 241, 262, 293, 550, 566, 245, 196, 120, 41, 206, 352, 368, 471, 355, 386, 377]. † [82]. E_{2} [223]. γ [8]. λ [503]. ′ [273, 3]. × [174, 34, 418].

-actin [377]. -adrenoceptor [206]. -amylase [409, 546]. -ATPase [50, 288]. -binding [417, 360]. -carotene [196]. -cyhalothrin [503]. -dihydroxy-4-pregenen-3-one [41]. -estradiol [529, 550]. -gustducin [396]. -hydroxylase [566, 355]. -hydroxylase/c17 [355]. -hydroxysteroid [368, 355]. -like [131]. -linolenic [8]. -Lipoic [421]. -MSH [131]. -N- [120]. -napthoflavone [128]. -proteinase [540]. -subunit [246]. -tetrachlorobiphenyl [3]. -triiodo-L-thyronine [273].

/HCO [86].

1 [362]. 1- [178]. 1-14C [57, 8]. 11-ketotestosterone [529, 250]. 126 [126]. 14C [57, 8]. 17-hydroxyprogesterone [78].

2 [57]. 2-Methylisoborneol [65]. 21-hydroxylase [78]. 26S [291]. 2N [518].

3 [57, 8]. 3n [8]. 3n-3 [8].

4-7-1942 [215]. 43 [95].

6-n-propyl-2-thiouracil [273].

7813 [95].

8-5-2003 [215].

A. [572]. ability [70, 586]. abnormal [438]. abnormalities [294, 577, 544]. Absence [340, 65]. Absorption [4, 136]. abu [476]. abundance [222]. acanthias [38]. Acanthopagrus [487, 255]. acanthopterygian [295]. acclimation [487, 501, 299, 288, 90, 407, 30]. acclimatized [388]. account [455]. Accumulation [344, 489, 284]. Accuracy [163, 574]. acetylcholinesterase [385, 426, 41]. Acetylhexosaminidase [120]. Acid [328, 340, 279, 59, 72, 430, 53, 166, 280, 420, 167, 416, 197, 198, 199, 288, 480, 33, 84, 76, 451, 146, 556, 421, 497, 542, 8, 48, 113, 165, 285]. acid-enrichment [421]. acid/vitamin [76]. acidic [150]. Acids [318, 66, 4, 304, 420, 89, 503, 25, 431, 133, 27, 39, 57, 149, 8]. acids** [80]. Acipenser [391, 91, 382, 507, 97, 40, 134, 539, 571, 449, 177, 182, 389, 74, 445]. Acipenseriformes [138]. Acquisition [269, 53]. across [578]. ACTH [35, 78, 69]. actin [377]. actinomycin [7]. Action [211, 208, 436, 516, 63, 180]. actions [216]. activating [63]. activation [26, 135]. activation/fertilization [26]. Activin [234, 240]. activities [387, 59, 416, 272, 435, 409, 510, 39, 185, 149, 551]. activity [508, 145, 408, 78, 50, 82, 403, 174, 175, 28, 234, 554, 119, 288, 385, 223, 243, 117, 538, 423, 126]. aculeatus [147, 159, 241]. Acute [545, 489, 490, 170, 341, 410, 390, 37]. adaptation [539, 55, 453, 277, 189, 60]. Adenine [317]. adenohypophyseal [520]. Adenosine [26]. adenylate [63]. administered [54]. Administration [432, 128, 457]. ADP [383]. adrenal [35]. Adrenergic [203]. adrenoceptor [206]. adult [548, 130, 148, 558]. advances [12, 358, 462]. aeglefinus [193]. aerobic [94]. Aeromonas [572]. affect [451]. affected [574, 167, 542]. affects [583, 59, 556]. African [275, 28, 423, 317]. after [72, 73, 424, 565, 407, 126, 144]. against [249, 491]. age [435, 148]. agents [77, 115]. ages [39]. aggressive [230]. agonist [82, 244, 9]. air [511, 464, 348, 359, 353, 347, 453, 349]. air-breather [511]. air-breathing [464, 348, 359, 347, 453, 349]. akaara [393]. alanine [139]. albacares [306]. albinism [437]. albinism-related [437]. albino [272]. Albumin [340]. albus [160, 138, 566]. Alcohol [293]. alevins [95]. alewives [277]. alginate [99, 13]. alginate-based [13]. algorithms [574]. alimentary [40, 418]. allocation [479]. allogynogenetic [551]. Alloxan [324]. Alosa [277]. alpha [131, 246]. alpha-melanocyte-stimulating [131]. alpinus [172, 25]. already [374]. alter [82]. Alteration [424, 84, 384, 347]. Alterations [387, 561, 338, 348, 503]. Altered [128]. Alters [328, 170]. Aluminum [319, 150]. Amazon [524]. Amazonian [511, 520]. amberjack [71]. ambient [69, 306]. America [519, 522]. americana [183]. Ameriurus [68]. amino [4, 72, 430, 280, 416, 198, 199, 133, 27, 39, 542, 149]. ammonia [453]. among [497]. AMP [77, 383, 560]. AMP-deaminase [560]. Amur [389]. amylase [418, 409, 546]. Anabas [557]. anadromous [300]. Anaesthetic [336]. anaesthetics [191]. analogue [557]. Analysis [160, 327, 342, 508, 564, 411, 375, 104, 246, 346, 443, 493, 422, 447, 377]. analytes [397]. analyzers [397]. Androgen [243, 124]. androgens [393, 200]. Anesthesia [335]. angiotensin [140]. Anguilla [270, 243, 227, 33, 407, 107]. anguillicaudatus [115, 559]. Anguilliformes [107]. anhydrase [38]. Annual [339, 31, 42]. annulatus [525]. anomalies [404, 18]. Antarctic [534]. Anti [329, 483]. Anti-Estrogen [329]. Anti-oxidant [483]. antibody [572, 541]. antioxidant [76]. Antioxidative [491]. antisera [249, 295]. antithrombin [540]. apical [86]. ApoC [504]. ApoC-I [504]. apoptosis [53, 500]. Apparent [133]. Appetite [321]. application [459, 41]. Applied [569, 519, 358, 576]. approach [252]. Approaches [569]. Aquaculture [461, 574, 525, 358, 463, 405]. aquaculture-affected [574]. aquaporin [487]. aquatic [163]. AQUI [191]. AQUI-S [191]. Arachidonic [328, 197, 89]. Arapaima [511, 520, 529]. Arctic [172, 25]. area [404]. areas [298]. argentiventris [96]. arginine [487, 88, 98]. Ariosoma [242]. Aromatase [329, 222, 566, 356, 257, 255, 260, 83, 375, 231, 369]. aromatases [370]. aromatizable [393]. arrest [115]. Arsenic [466]. Arsenic-induced [466]. Artemia [28, 477, 510, 195]. artery [553, 311]. artificial [427, 484]. asamushi [120]. ascorbic [421]. Asian [280, 483, 345, 1]. Assay [329, 83, 345, 246, 41, 109, 478, 265]. assess [533]. assessed [573]. assessment [158, 570]. assimilation [27]. associated [549, 222, 90, 81, 173]. Association [19]. astaxanthin [196]. asymmetry [429]. Atatürk [476]. Atheriniformes [259, 64]. Atlantic [4, 161, 279, 381, 45, 417, 72, 110, 494, 420, 477, 320, 136, 159, 143, 308, 379, 505, 196, 170, 267, 62, 39, 394, 542, 57, 247, 470, 553, 321, 149, 331, 48, 165, 285, 9, 498]. ATP [383, 317]. ATPase [50, 288, 269]. Atractosteus [403]. aurata [19, 10, 545, 489, 17, 18, 299, 480, 14, 125, 168, 568, 401, 84, 76, 16, 69, 43, 276, 544]. auratus [551, 304, 426, 502]. aureus [418]. Author [302, 334]. Authors [214, 313, 188]. Authors-index [214]. auto [13]. auto-fluorescence [13]. automated [397]. autotransgene [512]. autotransgenic [346]. autumn [123]. availability [133]. axis [43, 456]. Ayres [134].

Bacillus [562]. bacterial [448]. bacterium [562]. baeri [40, 177]. baerii [507, 182]. barbatus [401]. barbells [396]. Barbus [585]. Barfin [330, 219]. basal [520]. based [13, 409, 363]. Basic [569, 519, 358, 459]. Basin [268]. basis [94]. basolateral [86]. bass [508, 224, 312, 174, 34, 89, 221, 31, 178, 1, 81, 142, 56, 47, 478]. batrachus [345, 362]. Beagle [534]. beginning [561]. behavior [157, 230]. behavioural [82]. Beluga [517]. bengalense [160]. benzo [545]. beta [164]. beta-naphthoflavone [164]. better [463]. Between [344, 323, 494, 524, 571, 379, 449, 55, 575, 282, 287]. bifenthrin [580]. bigeye [306]. bile [80]. binding [381, 351, 417, 67, 119, 360, 127, 121, 400]. Bioaccumulation [475, 471]. Biochemical [339, 570, 270, 272, 177, 93, 450, 278, 391, 475, 474, 305, 414, 472, 384, 52, 510, 485, 468, 576, 499, 438, 580, 502, 508]. Biochemistry [569, 232, 389, 306, 559]. Biodiversity [519]. Bioenergetics [581, 517]. biological [571]. biology [535, 358]. biomarkers [164, 490]. biometrical [449]. biosynthesis [234]. biotransformation [65]. black [487, 414, 255, 40, 292]. black-brown [40]. blacodes [552]. bleached [283]. Bloch [387, 280, 457, 557, 1, 93, 563, 79]. Blood [327, 342, 87, 417, 428, 574, 23, 150, 424, 266, 540, 543, 397, 389, 100, 311, 445]. blue [189]. bluefin [417, 101, 105, 446]. blueprint [152]. Body [318, 328, 444, 428, 34, 85, 516, 441, 563, 442]. Bohr [87]. bonariensis [259, 411, 225, 532, 245, 527, 64, 231]. bonasus [388]. bone [193, 107]. bonnethead [186]. border [59, 174]. bovine [137]. Brachionus [510]. Brachymystax [239]. brachysoma [509]. brackish [539, 434]. brain [131, 561, 532, 83, 426, 556, 412, 51, 231, 370, 369]. brain-like [83]. brains [222]. Branchial [38, 191]. Brandet [517]. Brazilian [565, 530]. bream [545, 489, 15, 402, 18, 299, 14, 125, 168, 264, 13, 568, 76, 16, 69, 43, 21, 276, 544, 274, 292]. breather [511]. breathing [464, 348, 359, 353, 347, 453, 349]. breeding [147]. brevirostrum [382]. Bromage [215]. bromocriptine [15]. brood [571, 449]. broods [307]. broodstock [132]. brook [109, 148]. brown [95, 103, 40, 586, 68, 46, 2]. brush [59, 174]. buds [396]. buffering [112]. bullhead [68]. bullseye [525]. burbot [479]. Burchell [423]. butterflyfishes [518].

C [178, 304, 421, 495]. C-devoid [421]. C-labelled [304]. c17 [355]. c19 [549]. Ca [473]. cabrilla [401]. cadmium [475, 45, 489, 37, 385, 485]. caecal [25]. caged [224]. calcarifer [280, 483, 1]. calcium [77]. California [343, 29]. californicus [343]. calpain [395]. canadum [469]. canal [40, 92]. candidate [252]. canicula [425]. canthaxanthin [196]. capacity [161, 124, 112, 434, 400]. captive [513, 528, 31, 431]. capture [44, 448, 143]. captured [366]. carangid [237]. Caranx [101]. Carassius [304, 426, 551, 502]. carbamate [426]. carbamoyl [482]. Carbohydrate [401, 416, 516]. carbohydrates [54]. Carbon [70]. carbonic [38]. Carcharhiniformes [419]. Carcharias [513]. Cardiac [319, 380, 95, 150, 501, 92, 191, 306]. Cardiomyocytes [316, 139]. carnitine [34]. carotene [196]. carp [444, 408, 87, 240, 124, 570, 106, 288, 572, 90, 357, 459, 540, 575, 492, 577, 462, 185, 364, 471, 576, 8, 580, 100, 473, 180, 551, 144, 432, 502]. carpio [408, 87, 124, 570, 106, 572, 90, 540, 575, 492, 577, 185, 576, 580, 432]. carps [583, 486]. Cartilaginous [340, 419]. case [523]. Caspian [488, 586, 278]. caspius [488, 586]. castration [147, 159]. Catabolism [317]. catalase [408, 410]. Catecholaminergic [223]. Catfish [317, 275, 35, 475, 387, 351, 88, 436, 28, 85, 554, 348, 223, 222, 360, 359, 345, 516, 353, 121, 423, 24, 352, 122, 509, 368, 347, 65, 471, 478, 363, 349, 79, 386, 447, 396]. catfishes [453]. Catla [562, 357, 459, 485, 499]. caudal [472]. causing [500]. Cd [473, 491]. cDNA [17, 259, 225, 83, 90, 121, 407, 248, 205, 231, 365, 447]. cDNAs [160, 245, 229]. Cell [537, 95, 50, 507, 289, 262, 115, 246, 536, 366, 216, 8, 292]. cell-cycle [115]. cells [520, 472, 197, 20, 55, 433, 266, 176, 140, 512, 30, 249, 94, 327, 100, 63, 152, 445]. Cellular [53, 415, 75, 283]. Central [569]. Centropomus [521]. cephalus [229, 367]. certain [464, 424]. Cetti [404]. CGRP [553]. Chaetodon [518]. chain [71, 116]. Chalcalburnus [438, 514]. Challenge [328]. challenged [572]. change [257, 255, 260, 36, 261, 566, 371]. Changes [50, 72, 288, 412, 506, 173, 515, 10, 417, 472, 131, 570, 141, 92, 241, 222, 568, 42, 130, 251, 510, 267, 185, 93, 189, 427]. changing [375]. Channa [457, 503, 93, 355, 365]. Channel [35, 88, 85, 554, 222, 121, 24, 122, 478, 386, 447, 534]. channels [77]. char [25]. Characidae [528]. character [404]. Characterisation [78, 176]. Characteristics [332, 339, 417, 414, 552, 571, 438]. Characterization [145, 235, 225, 138, 120, 419, 3, 564, 174, 83, 119, 90, 541, 540, 177, 441, 440, 392, 116, 350, 406, 437, 504, 447, 446]. characterized [512]. characterizing [52]. charr [172]. Chemical [332, 141, 315]. chemicals [222]. chemistry [428, 574]. Chen [436]. Chinese [539, 389, 445]. Chinook [82, 268, 118, 191, 123, 336, 450, 287, 26]. Chirostoma [431]. chloride [50, 73, 390]. chlorinated [304]. CHO [512]. cholinergic [77]. chondrostean [579]. chromosomes [526]. Chronic [319, 424]. chronically [282]. chrysops [174, 34]. chuatsi [564]. chymotrypsin [145, 394]. Circadian [10, 312]. circulans [562]. circulating [449, 98, 400]. Circulations [323]. circulatory [494]. cirrhatus [61]. Cirrhinus [364]. citrulline [88]. Cl [86]. Clarias [275, 201, 475, 28, 348, 359, 345, 324, 353, 362, 423, 352, 368, 347, 349]. classes [431, 148]. Classification [574, 183]. cleavage [115, 473]. Cloning [487, 381, 17, 564, 83, 160, 241, 443, 205, 422, 437, 53, 258, 259, 310, 90, 245, 407, 229, 248, 369, 365, 274, 350, 504, 386, 447]. closed [417]. closed-system [417]. CNSS [472]. coastal [497, 307]. coated [136]. cobia [469]. Cod [553, 321, 161, 494, 477, 141, 308, 379, 62]. coding [443]. coeliac [553]. Coho [5, 273]. coioides [50, 384, 439, 314, 205, 377]. Cold [277, 299, 515]. coli [402, 205]. Colisa [226]. Collection [335, 23]. Colored [322]. Colossoma [87]. colour [189]. combined [498]. comments [513]. commercial [524, 572, 62]. common [240, 124, 106, 572, 132, 492, 577, 185, 576, 580]. Comparative [101, 550, 389, 91, 455]. compared [586]. Comparison [298, 397, 323, 559, 5, 175, 67, 571]. compensatory [161]. competence [53, 135]. competition [14]. competitive [83, 168]. complete [257]. component [554]. components [496, 97, 515]. Composition [339, 586, 444, 279, 59, 428, 270, 477, 34, 85, 89, 288, 480, 33, 84, 575, 146, 284, 497, 276, 563, 113, 278]. compositions [510]. compound [195]. concentration [47]. Concentrations [329, 72, 268, 186, 298, 173, 9, 390]. condition [282]. conditions [570, 31, 527, 238, 132, 2, 515, 446]. Confinement [126]. conflicts [513]. conger [242]. consequences [49]. conservation [467]. conspecifics [315]. constant [132, 399]. constitutive [512]. constraints [21]. constructs [512]. consumption [72, 299, 388]. containing [436, 117, 165]. Content [317, 398, 221, 192, 575, 556, 578]. Contents [213, 333, 187, 301, 503, 516]. continuous [23]. contractile [77]. Control [207, 223, 264, 157, 512]. Conversion [196, 379, 293, 149]. copper [481, 45, 541, 419, 94, 410]. copper-inducible [481]. coral [518, 44, 371]. Corals [374]. corn [436, 409]. corn-based [409]. correlated [238]. Correlation [282, 53]. Correlations [449]. corresponds [112]. Corrigendum [202, 271, 290]. corticotropin [35, 227]. corticotropin-releasing [227]. Cortisol [124, 208, 192, 128, 184, 449, 140, 43, 442, 478, 190]. Coryphaenoides [51]. could [94]. counts [404]. cownose [388]. Cr [491]. CRH [82]. croaker [555]. crossbred [575]. crucian [486, 100, 473, 551, 502]. crude [133, 502]. Cryopreservation [332, 462, 530]. crysoleucas [192, 442]. Ctenopharyngodon [180, 144]. Cu [474]. Cuban [403]. cucumber [441]. cues [315]. culture [507, 197, 62, 21, 8, 181, 446]. cultured [399, 300, 559]. Cultures [316, 289]. Curcumin [557]. Current [461, 20]. CXCR1 [564]. CXCR1-like [564]. CYA [95]. cyanobacterial [502]. cyclase [162, 63]. cyclase-activating [63]. cycle [472, 241, 222, 42, 115, 31, 482, 267, 285]. cyclic [77, 106]. cycloheximide [7]. Cyclosporin [68]. cyhalothrin [503]. CYP19a [549]. CYP2 [24]. CYP2-related [24]. Cypermethrin [503]. Cyprinidae [176, 346]. Cypriniformes [176]. Cyprinus [408, 87, 124, 570, 106, 572, 90, 540, 575, 492, 577, 185, 576, 580, 432]. cysteine [441]. cysteine-like [441]. cysts [28]. Cytochrome [368, 518, 549, 231, 370]. Cytological [294].

D [430, 7, 367]. D-amino [430]. Dahlgren [472]. daily [224]. Dam [476]. damage [545, 500]. damaged [119]. Danio [329, 119, 253, 113]. Danish [307]. danricus [410]. Dasyatis [267]. data [574]. day [329]. deaminase [560]. decapsulated [28]. deep [55, 51]. deep-hole [55]. deep-sea [51]. deethylase [554]. defence [76]. defences [309]. deficiency [384, 8]. deficient [84]. deformity [19]. degeneration [254]. dehydrogenase [368, 281, 355]. deiodinase [297]. delivery [311]. delta [480]. demands [94]. density [534, 448, 84]. dentex [132, 132]. dependence [10]. dependent [496, 254, 365, 135]. dephosphorylation [383]. depletion [359, 421]. deprivation [85]. derived [83]. desaturase [480]. desaturation [57, 113, 165]. detectable [80]. Detection [174, 119, 541, 238, 169, 478]. determinants [351]. Determination [71, 476, 514, 529, 372, 442, 254]. determining [256]. developing [297, 108]. Development [343, 508, 332, 111, 56, 442, 265, 292, 66, 403, 186, 141, 83, 528, 242, 582, 115, 256, 482, 7, 361, 109, 555, 307]. Developmental [429, 423]. devoid [421]. dextran [494]. Dicentrarchus [224, 178, 142, 337, 56, 47]. dichlorodiphenyltrichloroethane [471]. diel [41]. diet [106, 409, 546, 84, 509, 195]. Dietary [328, 59, 320, 556, 128, 45, 428, 73, 167, 416, 34, 517, 198, 199, 554, 439, 480, 546, 33, 76, 178, 451, 146, 469, 25, 142, 126, 193, 57, 563, 499, 434, 48, 285, 287]. diets [585, 562, 436, 136, 435, 13, 192, 516, 421, 142, 133, 165]. Difference [344]. Differences [572, 575, 283, 518, 450]. Different [338, 394, 585, 511, 398, 479, 330, 486, 298, 572, 383, 550, 253, 575, 132, 39, 30, 295, 484, 415, 390, 396]. Differential [240, 247]. differentially [566]. differentiation [252, 258, 554, 264, 253, 372, 251, 368, 573, 153, 370]. differently [477]. differing [175, 192, 516]. diffusion [47]. digestibility [457, 409, 133]. digestible [578]. Digestive [343, 161, 403, 195, 542, 149, 96, 508, 111, 175, 28, 435, 101, 409, 27, 555, 551, 446]. dihydroxy [41]. dimethoate [164]. dioxide [70]. dipeptides [4]. Diplodus [404]. diploid [382, 110, 109, 148, 450]. direct [168]. directions [461]. directly [190]. disadvantages [478]. Disease [320]. dismutase [419, 410]. Disparate [171]. dispersed [63]. disposition [65]. Disrupters [329]. disrupting [222]. distance [47]. distinct [228]. Distribution [396, 407, 356, 231]. disturbance [98, 492]. divergent [306]. diversity [573]. DM [256]. DMRT1 [259, 368, 411]. dmrt4 [558]. DMY [256]. DNA [545, 119, 573, 473]. docosahexaenoic [89, 556]. Does [303, 451, 82]. dogfish [131, 425, 38]. Dojo [559]. domain [256]. dominance [2]. dopamine [235, 367]. downregulate [220]. driven [522]. drop [69]. droplet [284]. drum [145]. ducts [267]. dumerilli [385]. During [318, 66, 585, 145, 252, 430, 240, 403, 570, 5, 258, 411, 186, 40, 141, 539, 241, 129, 263, 449, 42, 550, 566, 104, 105, 510, 423, 431, 412, 506, 267, 237, 361, 93, 471, 309, 355, 60, 558, 107, 250, 581]. dwarf [226]. dynamic [436, 516]. Dynamics [66, 318, 360]. dynein [269]. dysfunction [465].

Early [318, 81, 145, 430, 403, 258, 141, 7, 423, 309, 250]. eastern [404]. ecosystem [497]. eel [332, 270, 358, 243, 227, 566, 33, 407]. eelpout [496, 307]. eels [107]. Effect [444, 96, 562, 224, 414, 428, 99, 15, 457, 416, 320, 568, 359, 117, 324, 284, 142, 368, 354, 321, 341, 551, 201, 487, 475, 73, 139, 89, 14, 409, 546, 247, 555, 46, 576, 491, 269, 495]. Effects [329, 474, 279, 95, 417, 488, 150, 319, 501, 106, 164, 34, 554, 435, 448, 159, 439, 385, 273, 244, 393, 115, 6, 33, 527, 140, 76, 469, 423, 56, 336, 193, 276, 563, 61, 165, 285, 399, 484, 580, 100, 74, 545, 103, 78, 489, 82, 124, 518, 586, 147, 191, 379, 288, 582, 490, 31, 266, 171, 51, 2, 8, 48, 498, 410]. efficiency [379, 542, 563, 149]. effluent [283]. egg [280, 270, 1]. Eggs [401, 571, 52, 493, 429, 26]. eggsin [270]. Egypt [21]. Eicosanoid [328]. either [436]. electron [170]. elevated [282]. ELISA [168, 41, 109, 265]. elongation [57]. embryogenesis [66, 430, 431, 558]. Embryonic [413, 20, 289, 582, 483, 7, 108]. Embryos [401, 331, 548, 119, 115, 200]. emphasis [163, 458]. encoding [17, 378, 245, 220, 455, 512, 231, 422]. endangered [586]. endocardial [433]. Endocrine [329, 42, 222, 155]. Endocrinology [211, 153]. Energetic [317]. Energy [479, 534, 208, 337, 444, 16, 578]. enhances [82, 9, 432]. enhancing [234]. enriched [477]. enrichment [510, 421]. Enterococcus [538]. enterocytes [420, 25, 165]. entirely [300]. enveloping [433]. Environmental [62, 420, 31, 560, 47, 153]. enzymatic [510]. enzyme [508, 387, 403, 101, 409, 546, 345, 41, 56, 109, 392, 478, 265, 551]. enzyme-linked [345, 478, 265]. Enzymes [343, 325, 545, 489, 111, 416, 130, 482, 469, 423, 347, 195, 555, 446]. Epidermal [22]. epidermis [190]. Epinephelus [257, 50, 384, 439, 393, 314, 248, 205, 377]. epithelia [181]. epithelial [94]. epithelium [537, 552]. Eptatretus [61]. EROD [554, 126]. Erratum [32]. Erythrocyte [341, 146]. erythrocytes [544]. Escherichia [402, 205]. Esistance [320]. Esomus [410]. Essential [8, 279, 198, 199, 84]. Estimation [337]. estor [431, 431]. Estradiol [255, 529, 293, 550, 41, 352, 368, 93, 471]. estradiol-17 [293, 471]. estradiol-induced [368, 93]. Estrogen [329, 263, 554, 350]. Estrogen-induced [263]. estrogens [251, 200]. estuarine [497]. ethoxyresorufin [554]. ethoxyresorufin-O-deethylase [554]. ethynyl [368]. Eurasian [66, 111, 310, 492]. Europe [569]. European [332, 224, 312, 33, 178, 142, 337]. euryhaline [55]. Evaluation [329, 536, 144]. Evidence [228, 86, 190]. Evolutionary [326]. evolved [374]. exceptional [523]. exchange [494]. exchanger [86]. excretion [198, 199, 54]. Existence [226]. exogenous [546, 296]. experimentally [538]. exposed [283, 474, 511, 383, 550, 427, 282]. exposing [144]. Exposure [319, 327, 37, 150, 561, 424, 222, 568, 560, 126, 434]. express [262]. Expressed [172, 119, 228, 566, 512, 493]. Expression [548, 402, 314, 205, 355, 327, 558, 487, 252, 430, 564, 37, 413, 518, 15, 486, 179, 258, 411, 532, 83, 147, 241, 242, 480, 310, 263, 227, 222, 221, 244, 359, 117, 527, 407, 368, 347, 247, 443, 493, 370, 369, 365, 350, 504, 250, 447, 377, 396]. expressions [237, 394]. external [390]. extraction [442]. extracts [95, 518, 502]. extrahepatic [3]. Extranuclear [326]. extreme [515]. eye [311]. eyes [556].

F [288, 197, 288, 266]. Factor [274, 268, 310, 137, 449, 565, 407, 22, 171, 247]. Factor-I [274, 268, 310, 137, 449, 171]. facultative [511]. Families [338]. Family [212, 19]. farmed [575]. fasting [103, 276]. fat [521]. Fate [103, 304]. fathead [83]. Fatty [318, 328, 340, 66, 279, 304, 59, 166, 420, 288, 480, 33, 84, 76, 451, 146, 25, 431, 497, 57, 8, 48, 113, 165, 285]. fatty-acid [420]. features [91]. fed [585, 45, 436, 28, 141, 136, 192, 516, 84, 421, 195, 133]. Feed [136, 207, 279, 85, 379, 13, 405, 578, 563, 149]. feeding [398, 479, 23, 106, 525, 299]. female [387, 240, 34, 571, 129, 223, 449, 293, 221, 244, 105, 412, 506, 148, 369, 107]. females [527, 498]. ferruginea [294]. fertilization [26]. fertilizing [70, 586]. fibers [364]. fibroblast [292]. field [566, 254]. fillets [575]. final [354, 498]. finger [437]. fingerling [444]. fingerlings [448, 499]. fingerprints [429]. fins [547]. first [358]. Fish [328, 330, 523, 325, 208, 467, 569, 324, 344, 340, 209, 212, 322, 207, 323, 210, 201, 96, 481, 537, 414, 472, 564, 548, 574, 567, 259, 584, 524, 36, 86, 225, 136, 191, 67, 20, 571, 379, 12, 385, 582, 358, 263, 71, 375, 503, 245, 176, 178, 451, 372, 251, 371, 536, 556, 526, 463, 156, 356, 237, 485, 578, 440, 108, 422, 254, 232, 216, 8, 515, 530, 410, 311, 500, 581, 519]. fish-specific [422]. Fishes [522, 203, 464, 158, 87, 534, 535, 159, 101, 465, 579, 102, 458, 533, 249, 295, 51, 153]. FITC [494]. FITC-dextran [494]. five [397]. flatfish [289]. flavescens [184, 173]. flavin [117]. flavin-containing [117]. floodplains [535]. Flounder [58, 219, 428, 549, 294, 565, 412, 506, 27, 392, 265, 437, 558]. flow [311]. fluid [586, 92, 194]. fluidity [9]. fluorescence [13]. fluviatilis [66, 111, 310]. flux [100]. fluxes [511]. follicle [243, 156, 506]. follicle-stimulating [243, 506]. Follicles [401, 408, 135]. following [128, 37, 288, 189, 327, 26]. fontinalis [109, 148]. food [14, 81]. forage [497]. forebrain [362, 454]. Foreword [218, 373, 452, 569, 519]. form [246, 94]. forms [330, 183, 138]. formulated [562]. fossilis [387, 457, 223, 360, 354, 563, 79]. found [40]. Founder [330]. four [398, 67, 490]. fractions [383]. fragility [146]. franciscana [477]. Free [5, 275, 72, 246, 27, 39, 542, 149]. fresh [434]. freshwater [487, 466, 303, 73, 36, 535, 586, 424, 503, 557, 485, 453, 93, 281, 277, 8, 515, 530, 410, 181, 300]. FSH [258, 234, 236, 147, 241, 245, 249]. FSH- [147, 241, 245]. full [443]. full-length [443]. fulvidraco [396]. Function [320, 95, 380, 92, 229, 16]. Functional [97, 453, 536]. functionality [27]. functions [459]. Fundulus [37, 129]. furnieri [547, 102]. future [459, 461].

GABA [82]. gabaculine [88]. Gadus [161, 494, 477, 379, 62, 553, 321]. Gaimard [339]. gairdneri [116]. gamete [194]. gar [403]. gariepinus [275, 475, 28, 348, 359, 353, 423, 352, 368, 347, 349]. garlic [201]. Garra [176]. gas [47]. Gasterosteus [147, 159, 241]. gastrointestinal [398]. gelatinized [409]. Gender [529]. Gene [252, 327, 481, 564, 53, 486, 378, 549, 286, 263, 244, 256, 565, 220, 366, 455, 512, 350, 437, 386, 447, 377]. generating [500]. Genes [331, 252, 260, 262, 242, 375, 566, 527, 220, 422, 250]. Genetic [573, 207, 455]. genetically [572, 39]. Genetics [526, 467]. genistein [554]. genital [267]. Genomic [212, 252]. genotypes [49, 308]. genus [455, 51]. Genypterus [552]. germ [536, 152]. GH [273, 565, 314, 354, 400]. GH-binding [400]. GH-transgenic [273]. GH/IGF [354]. GH/IGF-I [354]. GHR [565]. Ghrelin [314]. gibelio [551]. gigas [511, 520, 529]. gill [50, 117, 395, 61, 94, 181]. gill-specific [395]. Gills [326]. gilthead [19, 17, 18, 299, 480, 14, 84, 76, 16, 69, 43, 276, 544]. given [88]. glass [358]. GLP [362]. GLP-1 [362]. Glucagon [362, 209, 81]. Glucagon-like [362]. Glucose [324, 4, 103, 303, 567, 436, 54]. Glutathione [325, 3]. Glycoprotein [160, 116]. glycoproteins [176]. GnRH [219, 233, 225, 532, 226, 359, 228, 527, 230, 220, 366, 412, 364, 64, 365]. GnRHa [498]. gobiid [375]. golden [192, 442, 478]. Goldfish [330, 304, 234, 291, 157, 560, 63]. Gonadal [242, 264, 221, 438, 233, 252, 158, 372, 238, 171, 56, 153, 254, 456, 79]. gonadotropes [246]. Gonadotrophin [9]. Gonadotrophin-releasing [9]. gonadotrophs [249, 295]. Gonadotropin [330, 102, 217, 64, 79, 135, 408, 351, 258, 242, 360, 221, 244, 238, 237, 247, 248, 250]. Gonadotropin-dependent [135]. Gonadotropin-II [79]. gonadotropin-like [238]. Gonadotropin-Releasing [330, 102, 217, 64, 244]. Gonadotropins [239, 376, 237, 361, 232]. gonads [370, 438, 558]. gonochoristic [486]. gourami [226]. gpcr [366]. granules [40]. grass [180, 144]. Gray [539]. green [428, 134]. greenback [265]. grey [229]. group [14]. Grouper [314, 205, 257, 50, 384, 439, 393, 248, 504, 377]. groups [516, 2]. Growth [211, 477, 85, 212, 322, 321, 274, 210, 444, 562, 161, 381, 428, 268, 5, 106, 457, 34, 448, 49, 308, 379, 480, 310, 137, 449, 273, 14, 345, 565, 407, 22, 146, 469, 156, 16, 123, 455, 512, 171, 193, 394, 542, 247, 563, 149, 495, 63, 180, 551]. GtH [527]. Guanylyl [162]. gudgeon [490]. Gulf [388]. Gunther [509]. guppy [443]. gustducin [396]. gut [77, 509]. guttatus [525]. gynogenetic [486].

habitat [528]. habits [398]. haddock [193]. Haemato [499]. Haemato-biochemical [499]. haematological [414, 485, 576, 399, 427, 484, 580]. haematology [488, 559]. haemoglobin [87, 308]. haemoglobin-oxygenation [87]. hagfish [238, 61]. Halibut [343, 505, 196, 9]. Ham. [562, 364, 495]. Hamilton [444, 409, 546, 176, 346, 512, 499]. handling [44, 143]. hatch [483]. hatching [392]. having [438]. hCG [355]. hCG-induced [355]. HCO [86, 86]. head [78, 507, 355, 365]. health [463]. heart [10, 150, 191, 433]. heat [15, 266, 282]. heat-shock [282]. heavy [582, 71, 491]. Heckel [476]. held [84, 132, 2]. helleri [481]. Hematologic [305]. hematological [474, 91]. Hematology [382, 110, 276]. hemi [159]. hemi-castration [159]. Hemisphere [399]. Hemoglobin [114, 49, 97, 386]. hemoglobin- [386]. hemopexin [481]. Hepatic [338, 336, 567, 416, 22, 3, 24, 400]. hepatocyte [113]. hepatocytes [420, 178, 352, 57, 165]. herbal [432]. herbs [584]. heteroclitus [37, 129]. Heterogeneity [212]. heterologous [162, 541]. Heteropneustes [387, 457, 223, 360, 354, 563, 79]. hexachlorocyclohexane [471]. hierarchies [2]. high [252, 45, 112, 560, 84, 60]. high-salinity [60]. highly [420, 119]. hilarii [528]. hill [176]. Hippoglossus [505, 196, 9]. Hira [486]. Hirudinea [414]. histamine [296]. Histochemical [40, 547, 552, 272]. Histological [324, 18, 267]. histology [438]. Histones [326]. Histopathological [158, 580]. Histopathology [464]. histories [415]. history [533]. hole [55]. homeostasis [16]. homeoviscous [277]. homologous [162]. homologue [481]. honeycomb [257]. Hoplosternum [511]. Horabagrus [509]. Hormonal [157]. Hormone [211, 330, 160, 212, 35, 381, 131, 5, 85, 310, 137, 449, 243, 227, 244, 359, 345, 565, 102, 407, 217, 16, 506, 455, 512, 64, 349, 374, 9, 63, 180]. hormone-encoding [512]. hormone-releasing [180]. hormones [75, 224, 513, 261, 130, 123, 171, 456, 514, 74]. horn [92]. HPI [43]. HSP70 [499]. hubbsi [541]. Hucho [323]. human [481, 351, 240]. Huso [91, 517]. hybrid [174, 34, 418, 221, 146, 440]. hydroelectrolytic [568]. hydrolase [59]. hydrolysis [175]. hydrophila [572]. hydrostatic [51]. hydroxylase [78, 223, 566]. hydroxylase/c17 [355]. hydroxyprogesterone [78]. hydroxysteroid [368, 355]. hyperosmotic [565]. Hypersaline [328, 89]. Hyperthermic [331]. hypophyseal [456]. hypophysis [131]. hypothalamo [456]. Hypoxia [341, 203, 6, 33].

Ictalurus [35, 88, 85, 554, 24, 122, 386, 447]. idella [144]. idellus [180]. Identical [486]. Identification [481, 183, 291, 395, 440, 331, 520, 40, 249, 295]. IGF [402, 139, 310, 565, 205]. IGF-I [139, 310, 565, 354, 205]. II [545, 489, 175, 140, 79]. III [540]. Immature [341, 243, 180]. immune [16, 185, 495, 432]. immunity [562]. immunoassay [56]. immunoblot [246]. Immunochemical [323]. Immunocytochemical [520, 361, 249, 295]. Immunohistochemical [238, 131]. Immunological [345]. immunoreactive [35, 364]. immunoreactivity [131, 362, 454]. immunoregulation [432]. immunosorbant [41]. immunosorbent [345, 109, 478, 265]. immunostaining [506]. immunostimulant [495]. immunostimulating [99, 584]. impact [240]. Impaired [307]. Implication [357]. implies [472]. importance [400]. improvement [405]. incidence [404]. inclusion [428]. increase [60]. Increased [565]. incubated [429]. Index [188, 334, 214, 302]. India [464, 467, 461, 471]. Indian [444, 166, 465, 357, 459, 176, 458, 463, 453, 364, 93]. indicators [149]. indices [576]. individual [133]. induce [378]. Induced [338, 327, 300, 275, 466, 348, 263, 503, 465, 1, 368, 347, 485, 173, 93, 355]. induces [257, 500]. inducible [481, 380]. inducing [514]. Induction [30, 521, 3, 352, 126, 499, 282]. inert [13]. infected [538]. Influence [178, 485, 322, 57, 47, 434, 400, 194, 146]. influences [29, 25, 62]. information [459]. ingestion [13]. ingredients [578]. Inhibition [43, 197].

Inhibitor [329, 257, 540]. inhibitors [96, 380]. inhibitory [235]. inhibits [70, 557, 7]. initiated [216]. injected [125, 54]. injections [378]. inner [297]. inner-ring [297]. innervation [553]. iNOS [380]. insecticides [426]. Insights [209, 212, 456]. Instruction [313]. Insulin [407, 209, 274, 268, 139, 310, 137, 449, 565, 81, 122, 171, 39, 247]. Insulin-like [407, 274, 268, 310, 137, 449, 565, 171, 247]. insulins [121]. Intake [207, 136, 81]. integration [346]. intensities [506]. interactions [22]. interrelation [379]. interrelationship [455]. interrenal [69]. intervention [463]. Intestinal [174, 175, 86, 59, 138, 170, 284, 165]. intestine [568, 58]. intestines [4]. Intracranial [378]. intraperitoneally [125]. inversion [393]. involved [77, 255, 416]. Involvement [330, 16, 370]. iodothyronine [297]. Ion [398, 511]. Ionoregulatory [36]. ions [150]. isoelectric [468]. isoforms [381, 226]. isolated [178, 352, 512, 470]. Isolation [367, 24, 369]. isotocin [98]. isozymes [39]. issue [569, 519]. Italian [97]. ivermectin [125, 168].

Japanese [270, 549, 243, 227, 407, 412, 506, 27, 437, 107]. japonica [270, 243, 227, 407, 107]. japonicus [441]. jellyfish [120]. juices [201]. junctions [537]. jundiá [305, 339]. Juvenile [329, 418, 283, 382, 479, 88, 521, 428, 82, 380, 539, 379, 449, 14, 546, 76, 182, 315, 282]. Juveniles [337, 322, 50, 409, 84, 196, 556, 142]. juxtaglomerular [140].

Katsuwonus [306]. ketotestosterone [529, 250]. key [546, 468]. kidney [78, 303, 507]. kinase [291, 22, 135]. kinetics [88]. kisutch [5]. kits [478]. kitsutch [273]. kraft [283]. kyphosis [19].

L [273, 457, 434, 77, 139, 298]. L-alanine [139]. L-leucine [139]. l-selenomethionine [434]. l-thyroxine [457]. L-type [77]. L. [96, 585, 4, 279, 45, 72, 99, 124, 380, 420, 106, 416, 572, 568, 505, 196, 76, 105, 178, 25, 170, 169, 62, 56, 69, 193, 39, 394, 542, 47, 470, 321, 544, 576, 149, 285, 580, 498, 406]. labelled [304]. Labeo [444, 448, 409, 546, 127, 346, 455, 512, 495]. labrax [224, 178, 142, 337, 56, 47]. Lack [180, 78, 299, 310]. lactate [281]. lacustrine [130]. Lake [476, 431]. lalia [226]. lamellae [552]. lamprey [80]. lamta [176]. landings [524]. lard [556]. largemouth [478]. Larvae [318, 343, 328, 508, 50, 99, 111, 280, 18, 28, 477, 435, 89, 12, 480, 469, 195, 555, 531]. Larval [80, 66, 145, 172, 403, 134, 141, 483, 13, 101, 309, 307]. larviculture [521]. laser [366]. laser-captured [366]. late [449]. Lates [280, 483, 1]. Latin [519]. latipes [548]. Latris [328]. lazera [201, 324]. lead [423]. lecithin [284]. lecitotrophic [40]. lectin [127]. leidyi [278]. length [443]. lenok [239]. leopardus [44]. lessons [216]. lethal [94]. leucine [139]. leucocyte [89]. Leukocyte [538]. leukocytes [35]. level [59, 546, 560, 451, 469, 142, 285]. levels [398, 45, 496, 529, 5, 89, 159, 513, 385, 449, 409, 527, 412, 171, 56, 43, 563, 309, 79, 26, 400]. LH [233, 240, 236, 147, 241, 223, 245, 249, 221]. LH- [147, 241, 245]. LHRHa [300]. life [550, 533, 509, 152]. ligand [381, 351]. ligand-binding [381, 351]. ligands [226]. Light [322, 10, 527, 542, 399]. light-dependence [10]. lighting [189]. like [564, 131, 268, 378, 83, 289, 310, 137, 449, 433, 362, 565, 407, 238, 441, 171, 247, 274, 77]. Limanda [294]. line [292]. linear [87, 578]. lineata [328]. lineatus [531]. lines [262]. lining [170]. linked [345, 41, 109, 478, 265]. Linn. [362]. Linnaeus [414, 570]. linoleic [8]. linolenic [8]. linseed [25, 48, 165]. lipase [418]. Lipid [129, 431, 203, 45, 496, 59, 166, 477, 34, 136, 12, 192, 516, 557, 575, 146, 469, 284, 142, 285, 410, 287]. Lipids [320, 105, 66, 480, 31, 469, 515]. Lipofuscin [433]. Lipofuscin-like [433]. Lipoic [421]. lipopolysaccharide [380]. Lipoproteins [340]. Lithuanian [573]. littorale [511]. Live [405]. Liver [344, 545, 78, 489, 166, 117, 104, 84, 556, 297, 171, 93, 491, 473]. Liza [476]. loach [115, 559]. loading [103, 448]. local [378]. localization [227, 361, 365]. locomotor [82]. Long [507, 289]. Long-term [507]. lordosis [19]. Loreto [524]. Low [112, 72, 117, 276]. LPS [380]. LR [95]. lubrica [414]. lucioperca [318, 435]. Lupine [324]. luteinizing [506]. Lutjanus [96, 525]. Lutra [492]. lyase [355]. lymphocytes [500]. Lysine [338].

M. [174, 34]. M3 [77]. M3-like [77]. Maccullochella [141]. macquarensis [141]. Macrobrachium [281]. macrolepidotum [541]. Macrophage [320]. macropomum [87]. maculatofasciatus [508]. Maintenance [337, 444, 26]. major [444, 357, 540, 364, 236, 264, 244, 250]. malathion [385]. malcolmsonii [281]. male [585, 583, 240, 34, 186, 256, 267, 56, 476, 369, 9, 250]. malondialdehyde [298]. Manchurian [239]. mandarin [564]. manipulation [224]. mannose [127]. mannose-binding [127]. Marine [328, 574, 525, 86, 289, 101]. masou [130, 140]. Masu [330, 219, 233, 130, 140]. material [238]. Maturation [321, 449, 357, 105, 533, 156, 412, 506, 237, 354, 355, 514, 498, 74, 107]. maturational [53, 135]. maturing [243]. maxima [167]. maximize [198]. maximizes [199]. maximus [99, 518, 272, 289, 49, 6, 406]. meal [133]. measured [397]. Measurement [122, 41, 109, 265]. measurements [400]. Measures [1]. Mechanism [473]. Mechanisms [211, 63, 76, 372]. medaka [548, 256]. Media [332]. mediated [92]. medirostris [134]. Mediterranean [10, 404]. medium [586]. meeki [242]. Megalobrama [402, 274]. melampygus [101]. melanocyte [131]. Melanogrammus [193]. melatonin [224, 312, 98, 357, 459, 62]. members [14]. membrane [59, 174, 383, 400]. membranes [22, 51]. meridionalis [516]. meridonalis [436]. Meristic [404]. merra [257]. Mesogloea [120]. mesopotamicus [482, 421]. messenger [53]. Metabolic [304, 46, 306, 546, 33, 423, 142, 429, 282]. Metabolically [316]. Metabolism [197, 338, 401, 208, 337, 296, 567, 106, 416, 134, 68, 178, 451, 469, 200, 453, 281, 8, 48, 285]. metabolites [150]. Metallothionein [37, 489, 385, 309]. metals [582, 491]. Metasystox [424]. Methods [332, 181]. methomyl [490]. methyl [368]. Methylisoborneol [65]. methylmercury [517]. metomidate [191]. Mexico [388]. Microbial [463, 409, 405]. microchemistry [254]. Microcystin [500, 95]. microcystin-LR [95]. Microcystin-RR [500]. Microcystis [95]. microinjection [27]. microparticulate [13]. Micropogonias [547, 102]. microsatellite [573]. microscopical [170]. microsomal [78]. migrating [268]. Miichthys [555]. miiuy [555, 555]. mill [283]. Milt [335, 177, 462, 9]. mimic [554]. mineralisation [193]. Minimising [94]. minimize [198]. minimizes [199]. minireview [468]. minnow [83, 550]. Misgurnus [115, 559]. Mitchill [388]. mitochondria [55]. mitochondria-rich [55]. mitochondrial [288, 500]. mitsukurii [135]. mixture [432]. Mnemiopsis [278]. Mobilization [203]. Mode [211]. model [363, 8, 292]. modulate [126, 456]. modulates [190]. Modulation [546, 230, 223, 281, 153, 349]. Molecular [258, 259, 160, 310, 245, 372, 229, 248, 456, 274, 350, 504, 386, 53, 83, 138, 450]. monoaminergic [561]. monoamines [359]. monooxygenase [117]. monophase [163]. Monopterus [160, 566]. monsoon [471]. morhua [161, 494, 477, 379, 62, 553, 321]. Morone [174, 34, 89, 183, 81]. Morphological [552, 52, 267, 445, 472]. Morphology [509, 153, 50]. mortality [448, 490]. mortem [470, 335]. mossambicus [137, 293, 55, 126, 171, 454, 60]. Motility [335, 70, 586, 112, 269]. Moxostoma [541]. Mozambique [137, 126]. mrigala [364]. mRNA [487, 37, 532, 147, 314, 171, 60, 377]. mRNAs [30]. MS222 [191]. MSH [131]. mucus [541, 440, 116]. Mugil [229, 367]. Müller [510]. mullet [229]. Mullidae [401]. Mullus [401]. multiple [183, 226]. multiplicity [114]. murrel [93, 355, 365]. Muscarinic [77]. muscimol [82]. muscle [77, 523, 298, 433, 130, 560, 84, 146, 556, 149, 306, 515, 107]. mykiss [77, 474, 252, 70, 78, 413, 335, 150, 319, 501, 179, 114, 167, 316, 198, 199, 338, 98, 117, 104, 540, 7, 41, 22, 206, 284, 297, 3, 395, 108, 296, 491, 315, 429, 341, 399, 427, 484, 282, 311, 342, 415, 269, 400, 190]. myosin [71]. Mystus [424].

N [120, 279, 273, 57]. n-3 [279, 57]. n-6 [279, 57]. naccarii [97]. naphthalene [561]. naphthoflavone [164]. napthoflavone [128]. natriuretic [162, 61]. natural [528, 2, 515]. nearshore [497]. nebulosus [68]. neighboring [220]. Neotropical [440, 526]. nerka [130]. nesting [147]. Net [511]. network [376]. Neuroanatomical [231]. neuroendocrine [217]. neuromodulators [217]. Neuronal [454, 64]. neuropeptide [363]. neurosecretory [472]. neurotransmitters [561]. Niall [215]. Nibe [135]. Nibea [135]. Nile [258, 416, 325, 538, 361, 370, 350]. niloticus [73, 258, 416, 418, 133, 370, 113, 350, 390]. nine [497]. nitrate [423]. nitric [380, 454]. nitrite [390]. Nitrogen [167, 198, 199]. Non [87, 304, 37, 409]. non-chlorinated [304]. non-gelatinized [409]. Non-linear [87]. non-spawning [37]. nonaromatizable [393]. nonessential [199]. nongenomic [216]. normal [272]. normalization [400]. Northeast [461]. northern [105, 388]. Notemigonus [192, 442]. notothenioids [523]. novel [127, 24, 395, 108, 437]. novemfasciatus [96]. NPY [363]. nuclear [544, 370]. nucleic [167, 503]. Nucleotide [377, 317]. number [140]. Nutrient [280, 457]. Nutrition [525, 12]. Nutritional [179, 567, 420, 113, 562, 495].

O [554, 417]. O. [130]. obesus [306]. obligate [511]. obscure [300]. obscurus [300]. obtaining [289]. occur [303]. ocean [450, 166]. ocean-type [450]. ocellatus [145]. ocular [543]. Odontesthes [259, 411, 225, 532, 245, 527, 64, 231]. oil [136, 178, 556, 25, 284, 165, 285]. oils [178, 48]. okinawae [375]. olfactory [169]. olivaceus [428, 549, 27, 392, 558]. olive [428, 558]. omnivorous [502]. Onchorynchus [415]. Oncorhynchus [77, 474, 252, 70, 78, 82, 268, 413, 335, 150, 319, 501, 5, 179, 114, 118, 167, 316, 198, 199, 338, 191, 273, 98, 117, 130, 266, 104, 140, 7, 41, 22, 206, 284, 297, 3, 123, 336, 395, 108, 296, 450, 491, 315, 429, 341, 399, 427, 484, 282, 311, 342, 287, 26, 269, 400, 190]. Oncorhynchuss [540]. one [41]. onion [201]. Ontogenetic [131, 555]. Ontogenic [370]. Ontogeny [318, 118, 184, 418, 253, 27, 145, 435, 101]. oocyte [53, 528, 357, 354, 355, 514, 74, 135]. oocytes [291]. open [417]. open- [417]. opercular [18]. operculum [176]. Ophisternon [160]. optimal [49]. optimum [198, 199]. oral [88, 457]. orally [54]. Orange [314, 205, 504, 377]. Orange-spotted [314, 504, 377]. orbignyanus [565]. Oreochromis [258, 73, 416, 418, 137, 293, 55, 126, 133, 171, 454, 370, 60, 113, 350, 390]. organismal [283]. Organization [212]. organizations [286]. organs [464, 387, 186]. orientalis [446]. ornithine [482]. Oryzias [548]. osmolalities [30]. osmoregulate [58]. Osmoregulation [539, 208, 182]. osmoregulatory [73, 434]. osmotic [98, 146]. Osteichthyes [76]. other [571]. otolith [254]. otter [492]. outer [297]. outer- [297]. output [150]. Ovarian [156, 408, 552, 83, 586, 194, 459, 412, 506, 349, 135]. ovarian-derived [83]. ovaries [532, 294]. ovary [53, 376, 549, 365]. over-ripening [52]. over-wintering [492]. Overview [212, 584, 536]. oviparous [533]. ovulation [156, 173, 300]. oxidant [483]. oxidase [430]. oxidation [420, 136, 165]. Oxidative [341, 287]. oxide [380, 454]. oxidised [136]. Oxygen [299, 311, 560, 388, 47, 306, 500]. oxygenation [87]. oxystats [163].

P450 [83, 549, 355, 231, 370]. P450-c19 [549]. P45011 [566]. P4501A [518]. P450arom [566]. P450aromatases [368]. P450aromB [231]. PACAP [63]. Pacific [101, 446]. pacu [482, 421, 478]. Pagrus [236, 264, 42, 244, 102, 189, 250]. Paiche [529]. Pallas [438, 514]. pallid [138]. paracrine [376]. Paralabrax [508]. Paralichthys [343, 428, 549, 565, 27, 392, 558]. parallelus [521]. Parameters [335, 391, 474, 583, 501, 73, 424, 52, 399, 427, 484, 580, 432]. parr [279, 45, 5, 48, 415]. parr-smolt [5, 48]. Partial [446, 259, 119, 541, 178, 58, 231, 365]. partially [421]. parts [575]. parva [490]. pathway [303, 135]. pathways [297, 100]. pattern [548, 198, 558].

Patterns [524, 528, 486, 359, 41, 347, 370]. Pátzcuaro [431]. Pb [474]. PB7 [562]. PCB [126]. PCC [95]. PCR [83, 331]. Pectoral [547]. pejerrey [259, 411, 225, 532, 245, 64, 231]. pelamis [306]. Pelteobagrus [396]. peptide [362, 180]. peptide-1 [362]. peptide-6 [180]. peptides [162, 249, 61]. Perca [66, 111, 184, 310, 173]. perch [66, 111, 183, 184, 310, 173]. perciform [102, 249]. Performance [319, 501, 457, 85, 29, 132, 542]. performances [551]. pericardial [92]. pericardioperitoneal [92]. period [240]. periods [510]. peripheral [445]. peroxidation [557, 410]. peroxidative [45]. Peroxiredoxin [447]. perryi [323]. Persian [571, 449]. persicus [391, 91, 571, 449]. perspectives [20, 467]. Peruvian [524]. Pesticide [465, 490]. Petroleum [275]. Petroleum-induced [275]. PGF [197]. pH [112, 27, 269, 531]. phagocytic [538]. Pharmacokinetics [168]. pharmacological [15]. phase [545, 489, 364]. phase-related [364]. pheromonal [157]. pheromone [80]. phosphatases [220]. phospholipid [469, 556]. phosphorus [193, 499, 107]. phosphorylate [291]. Photoperiod [321, 224, 459, 132, 56, 427]. photoperiods [484]. Physio [332]. Physio-Chemical [332]. Physiological [236, 324, 492, 2, 189, 201, 448, 143, 384, 439, 52, 345, 266, 29, 43, 287]. Physiology [569, 382, 501, 194, 6, 148, 232, 519]. phytase [174, 175, 133]. phytate [175]. Piaractus [482, 421]. pigment [40]. pigments [433]. Pikeperch [318, 435]. Pimelodidae [339]. Pimephales [83]. pineal [312, 364, 454]. Pirarucu [529, 511, 520]. Pisces [346]. Piscicolidae [414]. piscine [354, 152]. Pituitary [160, 314, 235, 532, 221, 362, 238, 412, 454, 63]. plant [96]. Plasma [332, 339, 88, 268, 494, 79, 502, 398, 103, 529, 5, 85, 159, 112, 31, 541, 540, 527, 127, 505, 469, 177, 122, 397, 171, 173, 56, 43, 39, 276, 9]. platyfish [433]. play [251]. Plectropomus [44]. plicatilis [510]. plumes [80]. Poecilia [164, 443]. Poeciliidae [433]. point [468]. polar [496]. Polydactylus [101]. Polymorphisms [549]. polypeptide [138, 63]. Polyunsaturated [48, 76, 113]. pools [542]. population [507]. populations [526, 573]. porcus [414]. porgy [487, 255, 42, 189]. Portal [336]. Possible [340, 203, 288]. Post [28, 147, 483, 195, 470, 335]. post-breeding [147]. post-hatch [483]. post-mortem [470, 335]. Post-prandial [28]. Potential [254, 472, 21, 269, 292]. PR [262]. PR-responsive [262]. practices [461]. prandial [28]. prawn [281]. praziquantel [576]. Pre [583, 402, 471]. pre-IGF-I [402]. pre-monsoon [471]. Pre-spawning [583]. prebiotic [551]. precerebellin [378]. precerebellin-like [378]. precision [163]. preconditioning [543]. precooked [436]. Prediction [578]. Preface [11, 151, 204]. pregenen [41]. preliminary [174, 18]. presence [553]. present [38]. pressure [51]. pretreated [133]. Primary [316, 286, 323, 494, 71, 181]. Primordial [152]. prmt5 [548]. probiotic [562]. problems [468]. procedure [442]. process [566]. Prochilodus [531]. produce [80]. Production [328, 169, 341, 197, 358, 137, 462, 285]. products [136, 220]. Professor [215]. Profile [328, 297, 123, 195]. profiles [166, 280, 242, 368, 493]. profiling [366]. progesterone [200]. progestogen [262]. progestogens [9]. prolactin [15]. promelas [83]. Promoter [375, 234, 549]. promoters [367, 512]. promotes [293]. properties [97, 560, 476]. propyl [273]. prostaglandin [197]. protandrous [255]. Protease [39, 96, 28, 418, 441, 149, 406]. proteases [96, 440]. proteasome [291]. Protein [337, 444, 99, 15, 416, 378, 67, 291, 90, 221, 409, 503, 546, 22, 220, 142, 24, 133, 563, 282, 437, 135, 292]. protein/carbohydrate [416]. proteinase [540]. proteins [579, 266, 104, 505, 422]. Proteome [104]. protogynous [257, 261, 566]. protons [87]. PrP [422]. PRS [262]. Psetta [167]. Pseudobranch [342, 543]. pseudobranchial [311]. pseudoharengus [277]. Pseudolabrus [369]. Pseudoplatystoma [556]. Pseudorasbora [490]. PTU [273]. Puberty [154, 186, 229, 56]. PUFA [76]. PUFA/tocopherol [76]. Puffer [344, 525, 300]. pufferfish [286]. punctatus [35, 88, 457, 85, 554, 503, 24, 122, 93, 386, 447]. puntazzo [404]. Purification [541, 353, 540, 441, 392, 406, 183, 121, 109]. purified [95]. pyrene [545]. Pyridoxine [384]. pyruvate [281].

quality [585, 562, 270, 1, 542, 495, 530]. quantify [83]. Quantitation [13]. quantitative [246, 25, 109]. quelen [305, 339, 383]. Quillaja [106]. Quoy [339].

R [215, 367]. Rachycentron [469]. radical [275]. Radicals [341]. radioimmunoassay [122]. Rafinesque [513, 386, 447]. Rainbow [335, 319, 338, 206, 342, 77, 474, 252, 128, 78, 303, 413, 150, 501, 179, 114, 167, 561, 198, 199, 98, 52, 117, 266, 104, 540, 29, 7, 41, 22, 284, 297, 3, 200, 395, 296, 94, 116, 491, 315, 399, 427, 484, 100, 311, 181, 400, 190]. raised [249]. raises [9]. RAP [331]. RAP-PCR [331]. rapeseed [48]. Rapid [216]. rare [550]. Rat [314]. Rate [46, 10, 150, 33, 80, 322, 149]. rates [299, 429]. rather [86]. Ratio [329, 416, 199, 76]. ration [444, 72]. raw [436]. rays [388]. reabsorption [303]. reactivation [583]. reactive [500]. reared [91, 31]. rearing [521, 415]. receptor [487, 381, 235, 228, 565, 121, 22, 367, 350]. receptor-binding [121]. receptor-protein [22]. Receptors [212, 77, 351, 255, 532, 226, 262, 217, 237, 370]. recombinant [240, 402, 137, 243]. recrudescence [349]. Red [327, 145, 53, 236, 264, 42, 244, 393, 266, 478, 189, 100, 250]. red-spotted [393]. redhorse [541]. redistribution [560]. reduced [282]. reduction [69]. reef [371]. refeeding [85]. regimes [479, 379, 542, 427]. region [524]. regression [578]. regular [106]. regulate [229]. Regulation [233, 92, 480, 206, 210, 250, 203, 179, 567, 420, 221, 459, 16, 356, 155, 363, 113]. regulators [217]. Regulatory [211, 78, 560, 123]. Related [325, 208, 283, 561, 533, 24, 364, 422, 437, 190]. Relation [335, 270, 532, 513, 394]. Relationship [571, 287, 71, 55]. relative [14]. Release [330, 314, 341, 233, 312, 87, 169, 80, 63]. released [315]. Releasing [330, 227, 244, 102, 217, 64, 9, 180]. relevance [62]. Renal [54]. repair [473]. repeated [23]. replacement [556]. reporter [262]. Reproduction [364, 472, 460, 358, 363, 374]. Reproductive [535, 132, 224, 186, 532, 549, 241, 513, 129, 222, 42, 465, 31, 229, 267, 148]. required [256]. requirement [444]. Requirements [337]. requires [135]. rerio [329, 119, 253, 113]. Research [519, 584, 525, 467, 569]. reservoirs [107]. Resistance [328, 336]. respiration [583]. respiratory [464, 6, 94]. respirometers [163]. response [75, 283, 77, 381, 50, 380, 118, 44, 477, 85, 143, 572, 184, 54, 192, 538, 81, 43, 30, 495, 315, 287, 144]. Responses [98, 209, 128, 45, 466, 36, 448, 550, 33, 492, 485, 499, 415, 502, 190]. responsive [262]. responsiveness [310, 69]. results [69, 400]. retention [279, 457, 198, 199, 563]. reticulata [164, 443]. reveal [299]. reversal [263]. reverses [421]. reversibly [70]. review [521, 582, 16, 530]. Rhamdia [305, 339, 383]. Rhinoptera [388]. Rhizoprionodon [143]. rhizostomous [120]. Rhombosolea [265]. Rhopilema [120]. rhythms [224]. ribonucleic [53]. rice [566]. rich [55]. ring [297]. ripening [52]. River [268]. riverine [80]. rivers [573]. RNA [470, 135]. roach [479]. rods [71]. rohita [444, 448, 409, 546, 127, 346, 512, 495]. rohu [448]. Role [261, 459, 363, 210, 203, 472, 460, 266, 543, 482, 251, 281, 491]. roles [236, 107]. rose [525]. rotifer [510]. Routine [134]. RR [500]. RT [83]. RT-PCR [83]. rubripes [286, 344, 493]. ruthenus [177].

S [191, 3]. S-transferases [3]. sabina [267]. Sakhalin [323]. salar [4, 279, 45, 72, 110, 380, 420, 320, 136, 159, 170, 39, 394, 542, 470, 149, 331, 48, 165, 285, 498]. sale [570]. saline [586, 144]. salinity [50, 55, 60]. Salminus [528]. Salmo [4, 279, 45, 95, 72, 110, 488, 380, 420, 139, 320, 136, 586, 159, 170, 39, 394, 542, 573, 470, 46, 2, 116, 149, 331, 48, 165, 285, 498]. Salmon [330, 320, 246, 336, 331, 283, 219, 233, 4, 279, 381, 45, 72, 82, 110, 268, 488, 5, 380, 118, 420, 136, 159, 191, 273, 130, 140, 170, 123, 39, 394, 542, 57, 247, 470, 450, 149, 48, 165, 285, 498, 287, 26]. salmonicida [572]. salmonicida [?]AJeney:2009:DAR. Salmonid [323, 197]. Salmonidae [194]. Salmonids [211, 75, 162, 112]. Salvelinus [172, 25, 109, 148]. sampling [23]. sand [508, 513]. Sander [318, 435]. sandwich [246]. Sanlıurfa [476]. saponin [106]. sarba [15, 13]. saturated [25]. saxatilis [174, 34, 89, 81]. Saxitoxins [344]. SBP [67]. scale [533]. Scaphirhynchus [138]. scenario [463]. Schlegel [107]. schlegeli [487, 255]. Schneider [552]. schrenckii [389]. Sciaenidae [135]. Sciaenops [145]. Sciences [569]. scoliosis [19]. Scophtalmus [289]. Scophthalmus [99, 272, 49, 6, 406]. Scorpaena [414]. scorpion [414]. scyliorhinid [131]. Scyliorhinus [131, 425, 419]. Sea [21, 545, 224, 312, 489, 15, 18, 299, 242, 14, 125, 168, 264, 13, 568, 76, 178, 1, 142, 16, 80, 441, 56, 69, 43, 276, 573, 47, 51, 544, 292, 404, 278]. Seabass [337, 59, 280, 483]. Seabream [365, 19, 10, 17, 53, 236, 480, 244, 84, 250]. season [585, 129, 471]. Seasonal [472, 360, 130, 185, 267]. Seasonality [544]. seasonally [388]. seawater [73, 407, 58]. second [17]. Secondary [323, 494, 144]. secrete [35, 246]. secretion [240, 86, 223, 243, 360, 62, 180]. secretory [176]. seed [462]. Seeds [324]. selected [475, 164, 576]. Selective [351]. selenite [474]. selenium [491]. selenomethionine [434]. Semen [339, 581]. Seminal [332, 339, 460, 505, 112, 348]. seminiferous [537]. Senegal [309]. senegalensis [195, 309]. Sensitive [331]. sensitivity [425, 112, 426]. sensory [570]. separation [108]. septemfasciatus [248]. Sequence [160, 172, 443, 493, 231, 377]. sequences [486, 259, 225, 121]. Sequencing [205, 241, 369, 386]. Sequential [335]. sera [323]. serial [375]. serial-sex [375]. Seriola [385]. serpins [540]. Serranidae [401]. Serranus [401]. Serum [391, 186, 513, 305, 559]. seven [166]. sevenband [248]. several [370]. Sevin [424]. Sex [283, 329, 260, 371, 252, 257, 255, 258, 554, 261, 67, 263, 449, 264, 375, 393, 566, 256, 372, 251, 169, 368, 153, 254, 370, 79]. Sex-related [283]. sexfilis [101]. Sexual [321, 513, 253, 105, 533, 237, 374, 498, 107]. sexually [180]. SF1 [259]. shallow [44]. shark [166, 425, 92, 143, 419]. sharks [186, 513]. sharpnose [143]. shiners [192, 442, 478]. shock [15, 266, 282]. short [85, 104, 16]. short-term [85]. shorthead [541]. shortnose [382]. Siberian [507, 40, 177, 182]. sieboldi [369]. sign [34]. signal [51]. Signaling [206, 376]. Silurus [436, 516]. silver [15, 13, 473]. silverside [431]. simulated [2]. sinensis [539, 389, 445]. single [366]. Siniperca [564]. Sinnularia [518]. six [455]. size [444, 379, 14, 516, 140]. skeletal [404]. skin [297, 440, 116, 190]. skins [272]. skipjack [306]. Slices [344]. smolt [5, 48]. smolts [450]. smooth [77]. SMR [46]. snake [355, 365, 268]. snapper [525]. snook [521]. social [230, 46]. sockeye [130]. sodium [474, 73]. soft [518, 501]. soft-water [501]. sole [309]. Solea [195, 309]. somatolactin [17]. somatolactin-encoding [17]. somatostatin [179]. Somatostatins [210]. Some [201, 427, 474, 91, 111, 73, 449, 347, 438, 515, 580, 432]. source [146, 287]. sources [428]. south [404, 522]. south-eastern [404]. southeastern [497]. southern [522, 436, 516, 278, 399]. SOX9 [259]. Sox9b [443]. soybean [556, 284, 133]. sp [538, 556]. Space [342]. Sparidae [42, 401, 76]. Sparus [19, 10, 545, 489, 17, 15, 18, 299, 480, 14, 125, 168, 13, 568, 401, 84, 76, 16, 69, 43, 276, 544]. Spatiotemporal [430]. spawning [585, 583, 521, 240, 37, 157]. spawns [1]. special [569, 458, 519]. Species [518, 160, 96, 574, 166, 175, 525, 36, 67, 571, 419, 122, 455, 497, 578, 306, 515, 530, 500]. species-specific [122]. Specific [516, 436, 256, 122, 395, 422]. specificity [426]. Spectral [425]. spectrum [189]. Sperm [585, 332, 579, 530, 70, 583, 112, 269]. Spermatogenesis [155, 411, 536, 250]. spermatological [476]. Spermatozoa [335, 317, 586, 177, 577, 493, 581]. Sphoeroides [525]. Sphyrna [186]. spiggin [147]. spined [147, 159, 241]. spontaneous [173]. Spotted [205, 508, 525, 393, 314, 504, 377]. spp [518]. spring [82]. Squalus [38]. Stability [470, 71]. stably [262]. stage [40, 365]. stage-dependent [365]. stages [483, 449, 550, 423, 509, 309, 498]. Standard [388, 46]. starch [436, 546]. Starvation [338, 37, 104, 510, 126, 555]. starved [141]. status [85, 532, 20, 439, 483, 467, 461, 21]. stearidonic [197, 8]. steelhead [70, 429, 282, 415, 269]. stem [289, 20]. Stereological [342, 536]. steriod [261]. sterlet [177]. steroid [255, 186, 159, 513, 67, 171, 173, 216, 74]. steroidogenic [347]. steroids [233, 527, 169, 79]. Stichopus [441]. stickleback [147, 159, 241]. stimulant [15]. Stimulates [314, 234]. stimulating [131, 243, 506]. Stimulation [408, 137, 162]. stingray [267]. stocking [84]. stocks [571, 449]. stomach [406]. Storage [335, 581]. Stored [317]. Storer [294]. story [522]. strains [95, 253, 65]. strategies [175]. strategy [260]. stream [176]. Stress [75, 320, 331, 327, 315, 283, 128, 382, 118, 44, 448, 143, 184, 98, 266, 565, 230, 170, 126, 405, 43, 442, 46, 309, 287, 415, 144, 292, 190]. Stress-Induced [327]. stress-related [190]. stressed [315]. striatus [355, 365]. Striped [328, 174, 34, 89, 221, 31, 81]. stroke [150]. Structural [577, 116, 97, 71, 568, 429]. structure [19, 71, 455]. structures [435, 286, 121]. studied [27]. Studies [409, 324, 201, 252, 254, 216, 445, 292]. Study [326, 91, 131, 547, 18, 557, 25, 170, 363]. studying [229]. sturgeon [382, 507, 97, 23, 40, 134, 539, 571, 449, 138, 579, 177, 182, 389, 434, 74, 445]. sturgeons [91]. sub [534, 94, 287]. sub-Antarctic [534]. sub-lethal [94]. sub-yearling [287]. sublethal [150]. substitution [178]. Substrate [426]. Substrates [317]. subtypes [230]. subunit [114, 291, 221, 244, 246, 250]. Subunits [160, 258, 241, 245, 247]. success [358]. sufficient [80]. suggests [370]. sulfide [100]. sulpiride [15]. superoxide [419, 410]. supplement [405]. supplementation [409]. supplemented [106]. supply [543]. suppressant [15]. suppressed [263]. surface [216]. surgeon [389]. surubim [556]. Survival [203, 572, 469, 434, 531]. Susceptibility [320, 53]. Swimming [319, 501]. swordtail [481]. Synbranchiformes [160]. synthase [380, 454]. synthesis [233, 420, 93, 135]. synthesizing [124]. synthetic [249]. system [417, 472, 114, 374]. systems [219, 494, 64].

T [457, 273]. T. [306]. tachykinins [553]. tag [493]. tags [172]. Tail [336]. Taimen [323]. Takifugu [286, 344, 493, 300]. tambacu [440]. tambaqui [87]. tamoata [511]. tapirina [265]. tarichi [438, 514]. taste [396]. taurine [60]. taurus [513]. tea [428]. technique [23, 108]. technology [358]. Teleost [330, 326, 520, 87, 175, 73, 86, 159, 67, 424, 263, 55, 557, 362, 543, 533, 81, 156, 485, 108, 153, 410]. teleostean [472, 579]. Teleostei [528, 42, 433, 135]. teleosts [398, 466, 460, 228, 154, 456]. Temminck [107]. Temperate [328, 496, 515]. Temperature [496, 29, 279, 479, 583, 417, 50, 522, 36, 49, 379, 288, 90, 117, 527, 132, 69, 47, 254, 306, 282, 498, 415]. Temperature-dependent [496, 254]. temperatures [383, 276, 429, 484]. Temporal [297]. tench [298, 169]. term [507, 85, 289, 104]. terminalis [402, 274]. terraenovae [143]. testes [243]. testicular [124]. Testing [329]. testis [348]. testosterone [293, 244, 41, 368, 56]. testosterone- [368]. testudineus [557]. tetrachlorobiphenyl [3]. Tetrodotoxin [344]. their [376, 241, 513, 288, 238, 220, 3, 62, 237, 370]. Thermal [281]. thiamin [439]. thiouracil [273]. Thiourea [347]. Thiourea-induced [347]. third [350]. threadfin [101]. threatened [509]. Three [330, 574, 175, 258, 225, 147, 159, 241, 148, 65, 306, 515, 427]. three-spined [147, 159, 241]. threshold [54]. throughout [252]. Thunnus [417, 105, 306, 446]. thynnus [105, 417, 105]. Thyroid [349, 456, 85, 359]. thyronine [273]. Thyroxine [330, 348, 5, 457, 449]. Thyroxine-induced [348]. tiburo [186]. tiger [513]. tilapia [235, 258, 416, 418, 137, 263, 293, 54, 538, 146, 126, 133, 171, 361, 370, 60, 113, 350, 390, 454]. Time [335]. Tinca [298, 169]. Tissue [45, 328, 344, 342, 158, 72, 89, 407, 421, 108]. Tissues [208, 475, 103, 570, 470, 60]. tocopherol [76, 451]. tolerance [89, 405, 94, 277]. tolerances [306]. tools [229]. topmouth [490]. torazame [131, 419]. Total [318, 66, 5]. toxic [491]. Toxicity [125, 275, 490, 390]. trabeculae [433]. tracer [41]. Trachelobdella [414]. tract [398, 28, 418, 27]. traits [549, 449, 533]. Transcript [493, 222]. transcription [378, 565]. transcripts [227, 422]. transduction [51]. transfer [73]. transferase [482]. transferases [3]. transformation [5, 48]. Transgene [346]. Transgenic [262, 110, 273]. transient [69]. transmontanus [74]. transplantation [536]. Transport [340, 44, 448, 129, 568]. transporter [60]. Treated [324, 283]. Treatment [257, 310, 293, 407, 7, 498]. trevally [101]. triacylglycerol [556]. triangular [402, 274]. Tributyltinchloride [327]. triiodo [273]. triiodothyronine [5]. Trimma [375]. triphosphate [26]. triploid [382, 387, 110, 294, 109, 148, 450, 79]. triploidy [488]. trisaccharide [116]. tristoechus [403]. Trout [335, 319, 316, 338, 206, 327, 341, 342, 77, 474, 252, 128, 95, 103, 78, 303, 239, 413, 150, 501, 179, 114, 167, 44, 139, 561, 198, 199, 141, 586, 98, 52, 117, 266, 104, 540, 29, 7, 41, 22, 284, 297, 3, 200, 30, 395, 573, 109, 148, 296, 46, 2, 94, 116, 491, 315, 429, 399, 427, 484, 282, 100, 311, 415, 181, 400, 190]. Trumpeter [328]. trutta [488, 586, 95, 139, 573, 46, 2]. trypsin [39, 394]. tshawytscha [82, 268, 118, 191, 123, 336, 450, 287, 26]. tsinlingensis [239]. tuna [417, 105, 397, 306, 446]. turbot [99, 167, 272, 49, 6, 406]. turnover [99]. Two [160, 121, 96, 381, 87, 36, 159, 286, 101, 138, 375, 409, 228, 516, 102, 80, 397, 509, 295, 429]. type [77, 546, 450, 369, 374]. Types [342, 286, 228, 542]. tyrosinase [272]. tyrosine [223, 220].

Ubiquitous [60]. UCP2 [413]. ultrastructural [547, 18, 25]. undergoing [48]. Understand [207]. understanding [468]. unique [80]. Universal [295, 533]. unsaturated [420]. uptake [139]. Urea [482, 303, 117]. ureogenesis [167, 453]. urinary [54]. urine [23]. urotensin [227]. USA [268]. Use [495, 584, 478, 254, 292]. Using [207, 529, 168, 13, 41, 249, 300]. utilisation [167]. Utilization [340, 142, 563]. UV [119]. UV-damaged-DNA [119].

vaccinated [572]. vaccine [572, 542]. vagus [191]. Val [144]. Valenciennes [528]. Validation [478, 345, 56]. values [305]. variant [17]. variants [532]. Variation [497, 207, 573]. Variations [339, 308, 309, 49, 364]. varieties [572]. variety [468]. Various [317, 428, 71, 249]. Vascular [336]. vasculature [61]. vasoactive [138]. vasoactivity [553]. vasotocin [487, 98]. vegetable [178, 285]. Vein [336]. Venezuelan [535]. ventricle [191]. Ventricular [316]. vertebrae [193]. vertebrate [374]. vertebrate-type [374]. vesicle [460]. vesicles [348]. vessels [191]. via [303]. Viable [316]. Vibrio [568]. VIP [138]. vital [387]. vitamin [196, 76, 421, 495]. Vitellogenesis [458, 561]. Vitellogenic [401]. Vitellogenin [329, 529, 183, 541, 353, 352, 171, 109, 93, 265]. vitro [408, 568, 557, 352, 354, 94, 74, 26, 312, 381, 206, 200, 247, 355, 180, 292, 190]. vittatus [424]. viviparus [307]. vivo [557, 7, 27, 103, 175, 180]. Volume [187, 213, 301, 333, 150, 92]. vulnificus [568].

Walbaum [474, 284, 491, 190]. wall [441]. Wap65 [481, 286, 90]. warm [90]. warm-temperature-acclimation-associated [90]. Water [531, 583, 501, 44, 539, 80, 47, 434, 144]. waterborne [37]. waters [511, 307]. weaning [435]. which [199, 291, 262]. white [23, 183, 560, 149, 434, 74]. Whole [328, 442, 285]. whole-body [442]. Wild [321, 158, 404, 242, 431, 412, 559]. wintering [240, 492]. within [549, 2]. world [358]. Wrasse [369, 261].

Xiphophorus [481]. XX [361]. XY [263, 361]. xylooligosaccharides [551].

Y-specific [256]. yearling [268, 287]. yellow [184, 509, 173, 396]. yellowfin [306]. yellowtail [294, 237]. yolk [108]. young [91, 563].

Zebrafish [329, 430, 376, 119, 222, 253, 113]. zinc [419, 485, 437]. Zoarces [307]. Zoarcidae [496].

References

Nocillado:2000:MEQ

[1] J. N. Nocillado, V. D. Peñaflorida, and I. G. Borlongan. Measures of egg quality in induced spawns of the Asian sea bass, Lates calcarifer Bloch. Fish Physiology and Biochemistry, 22(1):1–9, January 2000. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

     link.springer.com/article/10.1023/A:1007881231664.

Sloman:2000:PED

[2] K. A. Sloman, K. M. Gilmour, and N. B. Metcalfe. Physiological effects of dominance hierarchies within groups of brown trout, Salmo trutta, held under simulated natural conditions. Fish Physiology and Biochemistry, 22(1):11–20, January 2000. CODEN FPBIEP. ISSN 0920-1742 (print),

     1573-5168 (electronic). URL https://link.springer.com/article/10.
     1023/A:1007837400713.

Perez-Lopez:2000:CHE

[3] M. Pérez-López, P. Anglade, and P. Rouimi. Characterization of hepatic and extrahepatic glutathione S-transferases in rainbow trout (Oncorhynchus mykiss) and their induction by 3,3′, 4,4′-tetrachlorobiphenyl. Fish Physiology and Biochemistry, 22(1):21–32, January 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007885332573.

Bakke-McKellep:2000:AGA

[4] A. M. Bakke-McKellep, S. Nordrum, and R. K. Buddington. Absorption of glucose, amino acids, and dipeptides by the intestines of Atlantic salmon (Salmo salar L.). Fish Physiology and Biochemistry, 22(1):33–44, January 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007872929847.

Ebbesson:2000:FPT

[5] Lars O. E. Ebbesson, Björn Th. Björnsson, and Peter Ekström. Free plasma thyroxine levels in Coho salmon, Oncorhynchus kisutch, during parr-smolt transformation: comparison with total thyroxine, total triiodothyronine, and growth hormone levels. Fish Physiology and Biochemistry, 22(1):45–50, January 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1007841517552.

Maxime:2000:EHR

[6] Valérie Maxime, Karine Pichavant, and Guy Nonnotte. Effects of hypoxia on respiratory physiology of turbot, Scophthalmus maximus. Fish Physiology and Biochemistry, 22(1):51–59, January 2000. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007829214826.

Nagler:2000:VTC

[7] J. J. Nagler. In vivo treatment with cycloheximide or actinomycin D inhibits early embryonic development in rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 22(1):61–66, January 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1007825013917.

Tocher:2000:EFA

[8] Douglas R. Tocher and James R. Dick. Essential fatty acid deficiency in freshwater fish: the effects of linoleic, α-linolenic, γ-linolenic and stearidonic acids on the metabolism of [1-14C]18:3n-3 in a carp cell culture model. Fish Physiology and Biochemistry, 22(1):67–75, January 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1007877130756.

Vermeirssen:2000:GRH

[9] E. L. M. Vermeirssen, R. J. Shields, and A. P. Scott. Gonadotrophinreleasing hormone agonist raises plasma concentrations of progestogens and enhances milt fluidity in male Atlantic halibut (Hippoglossus hippoglossus). Fish Physiology and Biochemistry, 22(1):77–87, January 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1007820913009.

Aissaoui:2000:CHR

[10] A. Aissaoui, L. Tort, and J. Altimiras. Circadian heart rate changes and light-dependence in the Mediterranean seabream Sparus aurata. Fish Physiology and Biochemistry, 22(1):89–94, January 2000. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007861118404.

Anonymous:2000:P

[11] Anonymous. Preface. Fish Physiology and Biochemistry, 22(2):95, March 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1017247806152.

Izquierdo:2000:RAL

[12] M. S. Izquierdo, J. Socorro, and C. M. Hernández-Cruz. Recent advances in lipid nutrition in fish larvae. Fish Physiology and Biochemistry, 22 (2):97–107, March 2000. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007810506259.

Kelly:2000:QIF

[13] S. P. Kelly, S. D. Larsen, and N. Y. S. Woo. Quantitation of inert feed ingestion in larval silver sea bream (Sparus sarba) using auto-fluorescence of alginate-based microparticulate diets. Fish Physiology and Biochemistry, 22(2):109–117, March 2000. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007895429457.

Karplus:2000:EFC

[14] I. Karplus, D. Popper, and O. Goldan. The effect of food competition and relative size of group members on growth of juvenile gilthead sea bream, Sparus aurata. Fish Physiology and Biochemistry, 22(2):119– 123, March 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007885525902.

Deane:2000:EPP

[15] E. E. Deane, S. P. Kelly, and N. Y. S. Woo. Effect of a prolactin pharmacological stimulant (sulpiride) and suppressant (bromocriptine) on heat shock protein 70 expression in silver sea bream, Sparus sarba. Fish Physiology and Biochemistry, 22(2):125–133, March 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1023/A:1007807831274.

Perez-Sanchez:2000:IGH

[16] J. Pérez-Sánchez. The involvement of growth hormone in growth regulation, energy homeostasis and immune function in the gilthead sea bream (Sparus aurata): a short review. Fish Physiology and Biochemistry, 22 (2):135–144, March 2000. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007816015345.

Cavari:2000:CSV

[17] B. Cavari, B. Funkenstein, and H. Kawauchi. Cloning of a second variant of somatolactin-encoding cDNA from the gilthead seabream Sparus aurata. Fish Physiology and Biochemistry, 22(2):145–150, March 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1007826924006.

Galeotti:2000:PHU

[18] M. Galeotti, P. Beraldo, and M. Pinosa. A preliminary histological and ultrastructural study of opercular anomalies in gilthead sea bream larvae (Sparus aurata). Fish Physiology and Biochemistry, 22(2):151–157, March 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007883008076.

Afonso:2000:ALS

[19] J. M. Afonso, D. Montero, and R. Ginés. Association of a lordosis– scoliosis–kyphosis deformity in gilthead seabream (Sparus aurata) with family structure. Fish Physiology and Biochemistry, 22(2):159–163, March 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007811702624.

Hong:2000:ESC

[20] Y. Hong, S. Chen, and M. Schartl. Embryonic stem cells in fish: current status and perspectives. Fish Physiology and Biochemistry, 22(2):165– 170, March 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007871121784.

Sadek:2000:SBC

[21] S. Sadek. Sea bream culture in Egypt; status, constraints and potential. Fish Physiology and Biochemistry, 22(2):171–178, March 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007835126731.

Newsted:2000:EGF

[22] J. L. Newsted and J. P. Giesy. Epidermal growth factor receptor-protein kinase interactions in hepatic membranes of rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 22(3):181–189, April 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1007872209650.

Deng:2000:NTF

[23] D. F. Deng, S. Refstie, and S. S. O. Hung. A new technique of feeding, repeated sampling of blood and continuous collection of urine in white sturgeon. Fish Physiology and Biochemistry, 22(3):191–197, April 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1007803307617.

Perkins:2000:INH

[24] E. J. Perkins, B. C. DeBusk, and D. Schlenk. Isolation of a novel hepatic CYP2-related protein from channel catfish, Ictalurus punctatus. Fish Physiology and Biochemistry, 22(3):199–206, April 2000. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007833315734.

Olsen:2000:IDL

[25] R. E. Olsen, R. Myklebust, and T. M. Mayhew. The influences of dietary linseed oil and saturated fatty acids on caecal enterocytes in Arctic char (Salvelinus alpinus L.): a quantitative ultrastructural study. Fish Physiology and Biochemistry, 22(3):207–216, April 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/A:1007879127182.

Wendling:2000:ATL

[26] N. C. Wendling, D. C. Bencic, and R. L. Ingermann. Adenosine triphosphate levels of Chinook salmon (Oncorhynchus tshawytscha) eggs following in vitro maintenance and activation/fertilization. Fish Physiology and Biochemistry, 22(3):217–223, April 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1007870028938.

Ronnestad:2000:ODT

[27] I. Rønnestad, R. Pérez Dominguez, and M. Tanaka. Ontogeny of digestive tract functionality in Japanese flounder, Paralichthys olivaceus studied by in vivo microinjection: pH and assimilation of free amino acids. Fish Physiology and Biochemistry, 22(3):225–235, April 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007801510056.

Garcia-Ortega:2000:PPP

[28] A. García-Ortega, J. Verreth, and H. Segner. Post-prandial protease activity in the digestive tract of African catfish Clarias gariepinus larvae fed decapsulated cysts of artemia. Fish Physiology and Biochemistry, 22(3):237–244, April 2000. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007893223006.

Myrick:2000:TIC

[29] C. A. Myrick and J. J. Cech, Jr. Temperature influences on California rainbow trout physiological performance. Fish Physiology and Biochemistry, 22(3):245–254, April 2000. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007805322097.

Sakamoto:2000:IMR

[30] T. Sakamoto, N. Ojima, and M. Yamashita. Induction of mRNAs in response to acclimation of trout cells to different osmolalities. Fish Physiology and Biochemistry, 22(3):255–262, April 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/A:1007843508043.

Lund:2000:ACP

[31] E. D. Lund, C. V. Sullivan, and A. R. Place. Annual cycle of plasma lipids in captive reared striped bass: effects of environmental conditions and reproductive cycle. Fish Physiology and Biochemistry, 22(3):263– 275, April 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007818114057.

Anonymous:2000:E

[32] Anonymous. Erratum. Fish Physiology and Biochemistry, 22(3):277– 280, April 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1017370922589.

McKenzie:2000:EDF

[33] D. J. McKenzie, G. Piraccini, and E. W. Taylor. Effects of dietary fatty acid composition on metabolic rate and responses to hypoxia in the European eel (Anguilla anguilla). Fish Physiology and Biochemistry, 22(4): 281–296, May 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007865327923.

Gaylord:2000:EDC

[34] T. G. Gaylord and D. M. Gatlin III. Effects of dietary carnitine and lipid on growth and body composition of hybrid striped bass (Morone chrysops [female sign] × M. saxatilis [male sign]). Fish Physiology and Biochemistry, 22(4):297–302, May 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1007843223107.

Arnold:2000:CCI

[35] R. E. Arnold and C. D. Rice. Channel catfish, Ictalurus punctatus, leukocytes secrete immunoreactive adrenal corticotropin hormone (ACTH). Fish Physiology and Biochemistry, 22(4):303–310, May 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007884903577.

Gonzalez:2000:IRT

[36] R. J. Gonzalez and D. G. McDonald. Ionoregulatory responses to temperature change in two species of freshwater fish. Fish Physiology and Biochemistry, 22(4):311–317, May 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1007837214327.

Cleef-Toedt:2000:MME

[37] K. A. Van Cleef-Toedt, L. A. E. Kaplan, and J. F. Crivello. Metallothionein mRNA expression in spawning and non-spawning Fundulus heteroclitus following acute exposure to starvation and waterborne cadmium. Fish Physiology and Biochemistry, 22(4):319–327, May 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007854106269.

Wilson:2000:BCA

[38] J. M. Wilson, D. J. Randall, and G. K. Iwama. Branchial carbonic anhydrase is present in the dogfish, Squalus acanthias. Fish Physiology and Biochemistry, 22(4):329–336, May 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1007890000123.

Rungruangsak-Torrissen:2000:PAP

[39] K. Rungruangsak-Torrissen and A. Sundby. Protease activities, plasma free amino acids and insulin at different ages of Atlantic salmon (Salmo salar L.) with genetically different trypsin isozymes. Fish Physiology and Biochemistry, 22(4):337–347, May 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1007864413112.

Gisbert:2000:HIB

[40] E. Gisbert and C. Sarasquete. Histochemical identification of the blackbrown pigment granules found in the alimentary canal of Siberian sturgeon (Acipenser baeri) during the lecitotrophic stage. Fish Physiology and Biochemistry, 22(4):349–354, May 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1007895207177.

Nash:2000:ELI

[41] J. P. Nash, B. Davail-Cuisset, and D. E. Kime. An enzyme linked immunosorbant assay (ELISA) for testosterone, estradiol, and 17,20βdihydroxy-4-pregenen-3-one using acetylcholinesterase as tracer: application to measurement of diel patterns in rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 22(4):355–363, May 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1007850014021.

Kokokiris:2000:ECD

[42] L. Kokokiris, B. Mourot, and A. Fostier. Endocrine changes during the annual reproductive cycle of the red porgy, Pagrus pagrus (Teleostei, Sparidae). Fish Physiology and Biochemistry, 23(1):1–11, July 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007882807782.

Rotllant:2000:IHA

[43] J. Rotllant, R. J. Arends, and L. Tort. Inhibition of HPI axis response to stress in gilthead sea bream (Sparus aurata) with physiological plasma levels of cortisol. Fish Physiology and Biochemistry, 23(1):13–22, July 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1007848128968.

Frisch:2000:RCT

[44] A. J. Frisch and T. A. Anderson. The response of coral trout (plectropomus leopardus) to capture, handling and transport and shallow water stress. Fish Physiology and Biochemistry, 23(1):23–34, July 2000. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1007811302002.

Berntssen:2000:TLP

[45] Marc H. G. Berntssen, Anne-Katrine Lundebye, and Kristin Hamre. Tissue lipid peroxidative responses in Atlantic salmon (Salmo salar L.) parr fed high levels of dietary copper and cadmium. Fish Physiology and Biochemistry, 23(1):35–48, July 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1007894816114.

Sloman:2000:ESS

[46] K. A. Sloman, G. Motherwell, and A. C. Taylor. The effect of social stress on the standard metabolic rate (SMR) of brown trout, Salmo trutta. Fish Physiology and Biochemistry, 23(1):49–53, July 2000. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007855100185.

Saroglia:2000:IET

[47] M. Saroglia, S. Cecchini, and A. De Stradis. Influence of environmental temperature and water oxygen concentration on gas diffusion distance in sea bass (Dicentrarchus labrax, L.). Fish Physiology and Biochemistry, 23(1):55–58, July 2000. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007851430227.

Tocher:2000:PFA

[48] D. R. Tocher, J. G. Bell, and P. C. Morris. Polyunsaturated fatty acid metabolism in Atlantic salmon (Salmo salar) undergoing parr-smolt transformation and the effects of dietary linseed and rapeseed oils. Fish Physiology and Biochemistry, 23(1):59–73, July 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/A:1007807201093.

Imsland:2000:HGT

[49] A. K. Imsland, A. Foss, and G. Nævdal. Hemoglobin genotypes of turbot (Scophthalmus maximus): consequences for growth and variations in optimal temperature for growth. Fish Physiology and Biochemistry, 23(1): 75–81, July 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007815402911.

Caberoy:2000:CKA

[50] N. B. Caberoy and G. F. Quinitio. Changes in Na+,K+-ATPase activity and gill chloride cell morphology in the grouper Epinephelus coioides larvae and juveniles in response to salinity and temperature. Fish Physiology and Biochemistry, 23(1):83–94, July 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/A:1007827331137.

Siebenaller:2000:EHP

[51] J. F. Siebenaller. The effects of hydrostatic pressure on signal transduction in brain membranes of deep-sea fishes of the genus Coryphaenoides. Fish Physiology and Biochemistry, 23(2):99–106, August 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007830722804.

Lahnsteiner:2000:MPB

[52] F. Lahnsteiner. Morphological, physiological and biochemical parameters characterizing the over-ripening of rainbow trout eggs. Fish Physiology and Biochemistry, 23(2):107–118, August 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/A:1007839023540.

Choi:2000:CAS

[53] C. Y. Choi, A. Takemura, and F. Takashima. Cellular apoptosis susceptibility gene messenger ribonucleic acid in the ovary of red seabream: molecular cloning and correlation with acquisition of oocyte maturational competence. Fish Physiology and Biochemistry, 23(2):119–125, August 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1007826123986.

Lin:2000:RTU

[54] Shian-Chuann Lin, Chyng-Hwa Liou, and Shi-Yen Shiau. Renal threshold for urinary glucose excretion by tilapia in response to orally administered carbohydrates and injected glucose. Fish Physiology and Biochemistry, 23(2):127–132, August 2000. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007888108057.

Lee:2000:RBD

[55] T. H. Lee, P. P. Hwang, and C. H. Lin. The relationship between ‘deephole’ mitochondria-rich cells and salinity adaptation in the euryhaline teleost, Oreochromis mossambicus. Fish Physiology and Biochemistry, 23(2):133–140, August 2000. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007818631917.

Rodriguez:2000:DVE

[56] L. Rodríguez, I. Begtashi, and M. Carrillo. Development and validation of an enzyme immunoassay for testosterone: Effects of photoperiod on plasma testosterone levels and gonadal development in male sea bass (Dicentrarchus labrax, L.) at puberty. Fish Physiology and Biochemistry, 23(2):141–150, August 2000. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007871604795.

Ruyter:2000:IDF

[57] B. Ruyter, C. Røsjø, and M. S. Thomassen. Influence of dietary n-3 fatty acids on the desaturation and elongation of [1-14C] 18:2 n-6 and [1-14C] 18:3 n-3 in Atlantic salmon hepatocytes. Fish Physiology and Biochemistry, 23(2):151–158, August 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1007893317923.

Specker:2000:FPI

[58] J. L. Specker, K. Sundell, and B. Soffientino. Flounder with partial intestine osmoregulate in seawater. Fish Physiology and Biochemistry, 23 (2):159–164, August 2000. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007802911488.

Cahu:2000:DLL

[59] C. L. Cahu, J. L. Zambonino Infante, and D. Coves. Dietary lipid level affects fatty acid composition and hydrolase activities of intestinal brush border membrane in seabass. Fish Physiology and Biochemistry, 23(2):165–172, August 2000. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007807324809.

Takeuchi:2000:UIT

[60] K. Takeuchi, H. Toyohara, and M. Sakaguchi. Ubiquitous increase in taurine transporter mRNA in tissues of tilapia (Oreochromis mossambicus) during high-salinity adaptation. Fish Physiology and Biochemistry, 23(2):173–182, August 2000. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007889725718.

Simpson:2000:ENP

[61] C. W. Simpson, C. N. Glover, and M. E. Forster. Effects of natriuretic peptides on the gill vasculature of a hagfish, Eptatretus cirrhatus. Fish Physiology and Biochemistry, 23(2):183–189, August 2000. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007844623032.

Porter:2000:EIM

[62] M. J. R. Porter, S. O. Stefansson, and N. R. Bromage. Environmental influences on melatonin secretion in Atlantic cod (Gadus morhua L.) and their relevance to commercial culture. Fish Physiology and Biochemistry, 23(3):191–200, October 2000. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007816426509.

Wirachowsky:2000:MAP

[63] Noel R. Wirachowsky, Patrick Kwong, and John P. Chang. Mechanisms of action of pituitary adenylate cyclase-activating polypeptide (PACAP) on growth hormone release from dispersed goldfish pituitary cells. Fish Physiology and Biochemistry, 23(3):201–214, October 2000. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007837708880.

Stefano:2000:GRH

[64] A. V. Stefano, H. J. Aldana-Marcos, and G. M. Somoza. Gonadotropinreleasing hormone (GnRH) neuronal systems in the pejerrey, Odontesthes bonariensis (Atheriniformes). Fish Physiology and Biochemistry, 23 (3):215–223, October 2000. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007841813926.

Schlenk:2000:MDT

[65] Daniel Schlenk, Bryan DeBusk, and Everett J. Perkins. 2-methylisoborneol disposition in three strains of catfish: absence of biotransformation. Fish Physiology and Biochemistry, 23(3):225–232, October 2000. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007834720306.

Abi-Ayad:2000:DTL

[66] S.-M. E.-A. Abi-Ayad, P. Kestemont, and C. Mélard. Dynamics of total lipids and fatty acids during embryogenesis and larval development of Eurasian perch (Perca fluviatilis). Fish Physiology and Biochemistry, 23 (3):233–243, October 2000. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007891922182.

Hobby:2000:CSS

[67] A. C. Hobby, N. W. Pankhurst, and D. P. Geraghty. A comparison of sex steroid binding protein (SBP) in four species of teleost fish. Fish Physiology and Biochemistry, 23(3):245–256, October 2000. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007869713348.

Jegorov:2000:CMB

[68] A. Jegorov, P. Halada, and K. Šafarčík. Cyclosporin a metabolism in brown bullhead, ameriurus nebulosus. Fish Physiology and Biochemistry, 23(3):257–264, October 2000. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007844212439.

Rotllant:2000:DAT

[69] J. Rotllant, P. H. M. Balm, and L. Tort. A drop in ambient temperature results in a transient reduction of interrenal ACTH responsiveness in the gilthead sea bream (Sparus aurata, L.). Fish Physiology and Biochemistry, 23(3):265–273, October 2000. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007873811975.

Bencic:2000:CDR

[70] D. C. Bencic, J. G. Cloud, and R. L. Ingermann. Carbon dioxide reversibly inhibits sperm motility and fertilizing ability in steelhead (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 23(4):275–281, November 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1011157329855.

Kawabata:2000:DPS

[71] Reina Kawabata, Nobuyuki Kanzawa, and Takahide Tsuchiya. Determination of primary structure of amberjack myosin heavy chain and its relationship with structural stability of various fish myosin rods. Fish Physiology and Biochemistry, 23(4):283–294, November 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1011176105285.

Carter:2000:CTF

[72] C. G. Carter, D. F. Houlihan, and Z-Y. He. Changes in tissue free amino acid concentrations in Atlantic salmon, Salmo salar L., after consumption of a low ration. Fish Physiology and Biochemistry, 23(4):295–306, November 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1011187026997.

Fontainhas-Fernandes:2000:EDS

[73] A. Fontaínhas-Fernandes, F. Russell-Pinto, and João Coimbra. The effect of dietary sodium chloride on some osmoregulatory parameters of the teleost, Oreochromis niloticus, after transfer from freshwater to seawater. Fish Physiology and Biochemistry, 23(4):307–316, November 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1011156806524.

Webb:2000:ESH

[74] M. A. H. Webb, J. P. Van Eenennaam, and S. I. Doroshov. Effects of steroid hormones on in vitro oocyte maturation in white sturgeon (Acipenser transmontanus). Fish Physiology and Biochemistry, 23(4): 317–325, November 2000. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1011130924271.

Ackerman:2000:SHC

[75] P. A. Ackerman, R. B. Forsyth, and G. K. Iwama. Stress hormones and the cellular stress response in salmonids. Fish Physiology and Biochemistry, 23 (4):327–336, November 2000. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1011107610971.

Mourente:2000:EDP

[76] G. Mourente, E. Díaz-Salvago, and J. G. Bell. Effects of dietary polyunsaturated fatty acid/vitamin E (PUFA/tocopherol) ratio on antioxidant defence mechanisms of juvenile gilthead sea bream (Sparus aurata L., Osteichthyes, Sparidae). Fish Physiology and Biochemistry, 23(4):337–351, November 2000. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1011128510062.

Aronsson:2000:MML

[77] Ulrika Aronsson and Susanne Holmgren. Muscarinic M3-like receptors, cyclic AMP and L-type calcium channels are involved in the contractile response to cholinergic agents in gut smooth muscle of the rainbow trout, Oncorhynchus mykiss. Fish Physiology and Biochemistry, 23(4): 353–361, November 2000. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1011107630078.

Blom:2001:CHK

[78] S. Blom, L. Förlin, and T. B. Andersson. Characterisation of head kidney and liver microsomal 17-hydroxyprogesterone 21-hydroxylase activity in rainbow trout (Oncorhynchus mykiss) and the lack of regulatory effects of ACTH. Fish Physiology and Biochemistry, 24(1):1–8, January 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1011148604039.

Tiwary:2001:PLG

[79] Basant K. Tiwary, R. Kirubagaran, and Arun K. Ray. Plasma levels of gonadotropin-II and gonadal sex steroids in triploid catfish, Heteropneustes fossilis (Bloch). Fish Physiology and Biochemistry, 24(1):9–14, January 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1011159615057.

Polkinghorne:2001:LSL

[80] Christine N. Polkinghorne, Judith M. Olson, and Peter W. Sorensen. Larval sea lamprey release two unique bile acids** to the water at a rate sufficient to produce detectable riverine pheromone plumes. Fish Physiology and Biochemistry, 24(1):15–30, January 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/A:1011159313239.

Papatryphon:2001:EIG

[81] E. Papatryphon, E. Capilla, and J. H. Soares, Jr. Early insulin and glucagon response associated with food intake in a teleost, the striped bass (Morone saxatilis). Fish Physiology and Biochemistry, 24(1):31–39, January 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1011123908446.

Clements:2001:GAM

[82] S. Clements and C. B. Schreck. The GABA a agonist muscimol enhances locomotor activity, but does not alter the behavioural effects of CRH in juvenile spring Chinook salmon (Oncorhynchus tshawytscha)†. Fish Physiology and Biochemistry, 24(1):41–48, January 2001. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1011102420877.

Halm:2001:CMC

[83] S. Halm, M. Rand-Weaver, and C. R. Tyler. Cloning and molecular characterization of an ovarian-derived (brain-like) P450 aromatase cDNA and development of a competitive RT-PCR assay to quantify its expression in the fathead minnow (Pimephales promelas). Fish Physiology and Biochemistry, 24(1):49–62, January 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1011162920955.

Montero:2001:ALM

[84] D. Montero, L. E. Robaina, and M. S. Izquierdo. Alteration of liver and muscle fatty acid composition in gilthead seabream (Sparus aurata) juveniles held at high stocking density and fed an essential fatty acid deficient diet. Fish Physiology and Biochemistry, 24(1):63–72, January 2001. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1011145426543.

Gaylord:2001:GPB

[85] T. G. Gaylord, D. S. MacKenzie, and D. M. Gatlin III. Growth performance, body composition and plasma thyroid hormone status of channel catfish (Ictalurus punctatus) in response to short-term feed deprivation and refeeding. Fish Physiology and Biochemistry, 24(1):73–79, January 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1011199518135.

Grosell:2001:IHS

[86] M. Grosell, C. N. Laliberte, and C. M. Wood. Intestinal HCO_{3}^{−} secretion in marine teleost fish: evidence for an apical rather than a basolateral Cl^{−} /HCO+3− exchanger. Fish Physiology and Biochemistry, 24 (2):81–95, February 2001. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1011994129743.

Brauner:2001:NLR

[87] C. J. Brauner, T. Wang, and F. B. Jensen. Non-linear release of Bohr protons with haemoglobin-oxygenation in the blood of two teleost fishes; carp (Cyprinus carpio) and tambaqui (Colossoma macropomum). Fish Physiology and Biochemistry, 24(2):97–104, February 2001. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1011944407938.

Buentello:2001:PCA

[88] J. A. Buentello and D. M. Gatlin III. Plasma citrulline and arginine kinetics in juvenile channel catfish, Ictalurus punctatus, given oral gabaculine. Fish Physiology and Biochemistry, 24(2):105–112, February 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1011991312908.

Harel:2001:ETD

[89] M. Harel, S. Gavasso, and A. R. Place. The effect of tissue docosahexaenoic and arachidonic acids levels on hypersaline tolerance and leucocyte composition in striped bass (Morone saxatilis) larvae. Fish Physiology and Biochemistry, 24(2):113–123, February 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/A:1011924704459.

Kinoshita:2001:CCC

[90] S. Kinoshita, S. Itoi, and S. Watabe. cDNA cloning and characterization of the warm-temperature-acclimation-associated protein Wap65 from carp, Cyprinus carpio. Fish Physiology and Biochemistry, 24(2):125–134, February 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1011939321298.

Bahmani:2001:CSS

[91] Mahmoud Bahmani, Rezvan Kazemi, and P. Donskaya. A comparative study of some hematological features in young reared sturgeons (Acipenser persicus and Huso huso). Fish Physiology and Biochemistry, 24(2):135– 140, February 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1011911019155.

Head:2001:RCF

[92] Brian P. Head, Jeffrey B. Graham, and N. Chin Lai. Regulation of cardiac function in the horn shark by changes in pericardial fluid volume mediated through the pericardioperitoneal canal. Fish Physiology and Biochemistry, 24(2):141–148, February 2001. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1011915303225.

Sehgal:2001:BCL

[93] Neeta Sehgal and S. V. Goswami. Biochemical changes in the liver of the Indian freshwater murrel, Channa punctatus (Bloch) during estradiolinduced vitellogenin synthesis. Fish Physiology and Biochemistry, 24(2): 149–155, February 2001. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1011967401714.

Smith:2001:MAR

[94] Richard W. Smith, Maria Jönsson, and Peter Pärt. Minimising aerobic respiratory demands could form the basis to sub-lethal copper tolerance by rainbow trout gill epithelial cells in vitro. Fish Physiology and Biochemistry, 24(2):157–169, February 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1011932002623.

Best:2001:EPM

[95] J. H. Best, F. B. Eddy, and G. A. Codd. Effects of purified microcystin-LR and cell extracts of Microcystis strains PCC 7813 and CYA 43 on cardiac function in brown trout (Salmo trutta) alevins. Fish Physiology and Biochemistry, 24(3):171–178, April 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/A:1014081827372.

Alarcon:2001:EPP

[96] F. J. Alarcón, F. L. García-Carreño, and M. A. Navarrete del Toro. Effect of plant protease inhibitors on digestive proteases in two fish species, Lutjanus argentiventris and L. novemfasciatus. Fish Physiology and Biochemistry, 24(3):179–189, April 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1014079919461.

Clementi:2001:FSP

[97] M. E. Clementi, M. C. De Rosa, and B. Giardina. Functional and structural properties of the hemoglobin components from Italian sturgeon (Acipenser naccarii). Fish Physiology and Biochemistry, 24(3):191– 200, April 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1014088106257.

Kulczykowska:2001:RCA

[98] E. Kulczykowska. Responses of circulating arginine vasotocin, isotocin, and melatonin to osmotic and disturbance stress in rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 24(3):201–206, April 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1014079917644.

Conceicao:2001:EIA

[99] L. E. C. Conceição, J. Skjermo, and J. A. J. Verreth. Effect of an immunostimulating alginate on protein turnover of turbot (Scophthalmus maximus L.) larvae. Fish Physiology and Biochemistry, 24(3):207–212, April 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1014052611305.

Volkel:2001:ESK

[100] S. Völkel, M. Berenbrink, and M. Nikinmaa. Effects of sulfide on K^{+} flux pathways in red blood cells of crucian carp and rainbow trout. Fish Physiology and Biochemistry, 24(3):213–223, April 2001. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1014050001585.

Kim:2001:CDE

[101] Bong G. Kim, S. Divakaran, and Anthony C. Ostrowski. Comparative digestive enzyme ontogeny in two marine larval fishes: Pacific threadfin (Polydactylus sexfilis) and bluefin trevally (Caranx melampygus). Fish Physiology and Biochemistry, 24(3):225–241, April 2001. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1014054431627.

Montaner:2001:GRH

[102] A. D. Montaner, L. A. Miranda, and G. M. Somoza. Gonadotropinreleasing hormone in two perciform fishes: Micropogonias furnieri and Pagrus pagrus. Fish Physiology and Biochemistry, 24(3):243–246, April 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1014062415118.

Blasco:2001:FPG

[103] J. Blasco, I. Marimón, and J. Fernández-Borrás. Fate of plasma glucose in tissues of brown trout in vivo: effects of fasting and glucose loading. Fish Physiology and Biochemistry, 24(3):247–258, April 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1014084313207.

Martin:2001:PAR

[104] S. A. M. Martin, P. Cash, and D. F. Houlihan. Proteome analysis of rainbow trout (Oncorhynchus mykiss) liver proteins during short term starvation. Fish Physiology and Biochemistry, 24(3):259–270, April 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1014015530045.

Mourente:2001:LFN

[105] Gabriel Mourente, César Megina, and Esther Díaz-Salvago. Lipids in female northern bluefin tuna (Thunnus thynnus thynnus L.) during sexual maturation. Fish Physiology and Biochemistry, 24(4):??, March 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1015011609017.

Francis:2001:ECR

[106] G. Francis, H. P. S. Makkar, and K. Becker. Effects of cyclic and regular feeding of a Quillaja saponin supplemented diet on growth and metabolism of common carp (Cyprinus carpio L.). Fish Physiology and Biochemistry, 24(4):??, March 2001. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1015047208108.

Yamada:2001:RBM

[107] Y. Yamada, A. Okamura, and H. P. Oka. The roles of bone and muscle as phosphorus reservoirs during the sexual maturation of female Japanese eels, Anguilla japonica Temminck and Schlegel (Anguilliformes). Fish Physiology and Biochemistry, 24(4):??, March 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1023/A:1015059524947.

Shahsavarani:2001:NTS

[108] A. Shahsavarani, Z. C. Thomas, and P. A. Wright. A novel technique for the separation of yolk from the developing embryonic tissue in a teleost fish, Oncorhynchus mykiss. Fish Physiology and Biochemistry, 24(4): ??, March 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1015074405382.

Schafhauser-Smith:2001:PDQ

[109] D. Schafhauser-Smith and T. J. Benfey. The purification and development of a quantitative enzyme linked immunosorbent assay (ELISA) for the measurement of vitellogenin in diploid and triploid brook trout (Salvelinus fontinalis). Fish Physiology and Biochemistry, 24(4):??, March 2001.

CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1015085317364.

Cogswell:2001:HDT

[110] A. T. Cogswell, T. J. Benfey, and A. M. Sutterlin. The hematology of diploid and triploid transgenic Atlantic salmon (Salmo salar). Fish Physiology and Biochemistry, 24(4):271–277, March 2001. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1015020716456.

Cuvier-Peres:2001:DSD

[111] A. Cuvier-Péres and P. Kestemont. Development of some digestive enzymes in Eurasian perch larvae Perca fluviatilis. Fish Physiology and Biochemistry, 24(4):279–285, March 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1015033300526.

Ingermann:2001:LSP

[112] R. L. Ingermann, D. C. Bencic, and J. G. Gloud. Low seminal plasma buffering capacity corresponds to high pH sensitivity of sperm motility in salmonids. Fish Physiology and Biochemistry, 24(4):299–307, March 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1015037422720.

Tocher:2001:NRH

[113] D. R. Tocher, M. Agaba, and A. J. Teale. Nutritional regulation of hepatocyte fatty acid desaturation and polyunsaturated fatty acid composition in zebrafish (Danio rerio) and tilapia (Oreochromis niloticus). Fish Physiology and Biochemistry, 24(4):309–320, March 2001. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1015022406790.

Fago:2001:HSM

[114] A. Fago, E. Forest, and R. E. Weber. Hemoglobin and subunit multiplicity in the rainbow trout (Oncorhynchus mykiss) hemoglobin system. Fish Physiology and Biochemistry, 24(4):335–342, March 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1015063625855.

Liu:2001:ECC

[115] T. M. Liu, C. J. Wu, and X. Y. Li. Effects of cell-cycle arrest agents on cleavage and development of loach (Misgurnus anguillicaudatus) embryos. Fish Physiology and Biochemistry, 25(1):3–9, June 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1019731909925.

Sumi:2001:SCT

[116] Toshihisa Sumi, Yoichiro Hama, and Makio Asakawa. Structural characterization of the trisaccharide chain from the skin mucus glycoprotein of the rainbow trout, Salmo gairdneri. Fish Physiology and Biochemistry, 25(1):11–17, June 2001. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1019763214180.

Larsen:2001:EUL

[117] B. K. Larsen and D. Schlenk. Effect of urea and low temperature on the expression and activity of flavin-containing monooxygenase in the liver and gill of rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 25(1):19–29, June 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1019769026824.

Feist:2001:OSR

[118] Grant Feist and Carl B. Schreck. Ontogeny of the stress response in Chinook salmon, Oncorhynchus tshawytscha^{∗}. Fish Physiology and Biochemistry, 25(1):31–40, June 2001. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1019709323520.

Hsu:2001:DPC

[119] T. Hsu, C. S. Cheng, and F. L. Yeh. Detection and partial characterization of a UV-damaged-DNA binding activity highly expressed in zebrafish (Danio rerio) embryos. Fish Physiology and Biochemistry, 25(1): 41–51, June 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1019755900858.

Nagai:2001:CEB

[120] Takeshi Nagai, Yuko Watarai, and Nobutaka Suzuki. Characterization of β-N- acetylhexosaminidase from rhizostomous jellyfish, Rhopilema asamushi, Mesogloea. Fish Physiology and Biochemistry, 25(1):53–59, June 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1019707816787.

Mommsen:2001:TIC

[121] T. P. Mommsen, J. T. Silverstein, and J. M. Conlon. Two insulins from channel catfish: purification, structures, receptor-binding and cDNA sequences. Fish Physiology and Biochemistry, 25(1):61–71, June 2001. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1019708017696.

Plisetskaya:2001:MCC

[122] Erika M. Plisetskaya, Vera M. Bondareva, and J. T. Silverstein. Measurement of channel catfish (Ictalurus punctatus) plasma insulin in speciesspecific radioimmunoassay. Fish Physiology and Biochemistry, 25(1):73– 82, June 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1019760001766.

Pierce:2001:APG

[123] A. L. Pierce, K. D. Shearer, and W. W. Dickhoff. An autumn profile of growth regulatory hormones in Chinook salmon (Oncorhynchus tshawytscha). Fish Physiology and Biochemistry, 25(1):83–88, June 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1019717001668.

Consten:2001:CET

[124] D. Consten, E. D. Keuning, and H. J. Th. Goos. Cortisol effects on the testicular androgen synthesizing capacity in common carp, Cyprinus carpio L. Fish Physiology and Biochemistry, 25(2):91–98, September 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1020538729104.

Katharios:2001:TII

[125] P. Katharios, J. Iliopoulou-Georgudaki, and S. Spiropoulos. Toxicity of intraperitoneally injected ivermectin in sea bream, Sparus aurata. Fish Physiology and Biochemistry, 25(2):99–108, September 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1020574810332.

Quabius:2001:CSS

[126] E. S. Quabius, D. T. Nolan, and S. E. Wendelaar Bonga. Confinement stress and starvation modulate the induction of EROD activity after dietary exposure to PCB 126 in the Mozambique tilapia (Oreochromis mossambicus). Fish Physiology and Biochemistry, 25(2):109– 119, September 2001. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1020590726262.

Mitra:2001:NMB

[127] S. Mitra and H. R. Das. A novel mannose-binding lectin from plasma of Labeo rohita. Fish Physiology and Biochemistry, 25(2):121–129, September 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1020545510295.

Basu:2001:ASR

[128] N. Basu, C. J. Kennedy, and G. K. Iwama. Altered stress responses in rainbow trout following a dietary administration of cortisol and βnapthoflavone. Fish Physiology and Biochemistry, 25(2):131–140, September 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1020566721026.

Jensen:2001:LTF

[129] B. H. Jensen and M. H. Taylor. Lipid transport in female Fundulus heteroclitus during the reproductive season. Fish Physiology and Biochemistry, 25(2):141–151, September 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1020564121935.

Leonard:2001:SCH

[130] J. B. K. Leonard, M. Iwata, and H. Ueda. Seasonal changes of hormones and muscle enzymes in adult lacustrine masu (Oncorhynchus masou) and sockeye salmon (O. nerka). Fish Physiology and Biochemistry, 25(2): 153–163, September 2001. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1020512105096.

Chiba:2001:OCA

[131] A. Chiba and S. Oka. Ontogenetic changes of alpha-melanocyte-stimulating hormone (α-MSH)-like immunoreactivity in the brain and hypophysis of a scyliorhinid dogfish (Scyliorhinus torazame): An immunohistochemical study. Fish Physiology and Biochemistry, 25(2):165–170, September 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1020516206005.

Pavlidis:2001:RPC

[132] M. Pavlidis, S. Keravec, and A. P. Scott. Reproductive performance of common dentex, Dentex dentex, broodstock held under different photoperiod and constant temperature conditions. Fish Physiology and Biochemistry, 25(3):171–180, December 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1022284511903.

Riche:2001:ADC

[133] M. Riche, N. L. Trottier, and D. L. Garling. Apparent digestibility of crude protein and apparent availability of individual amino acids in tilapia (Oreochromis niloticus) fed phytase pretreated soybean meal diets. Fish Physiology and Biochemistry, 25(3):181–194, December 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1022222431004.

Gisbert:2001:RML

[134] E. Gisbert, J. J. Cech, Jr., and S. I. Doroshov. Routine metabolism of larval green sturgeon (Acipenser medirostris Ayres). Fish Physiology and Biochemistry, 25(3):195–200, December 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/A:1022260627062.

Yoshizaki:2001:GDO

[135] G. Yoshizaki, M. Shusa, and R. Patiño. Gonadotropin-dependent oocyte maturational competence requires activation of the protein kinase a pathway and synthesis of RNA and protein in ovarian follicles of nibe, Nibea mitsukurii (Teleostei, Sciaenidae). Fish Physiology and Biochemistry, 25 (3):201–208, December 2001. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1022209710224.

Hamre:2001:FIA

[136] Kristin Hamre, Kjersti Kolås, and Anders Kiessling. Feed intake and absorption of lipid oxidation products in Atlantic salmon (Salmo salar) fed diets coated with oxidised fish oil. Fish Physiology and Biochemistry, 25 (3):209–219, December 2001. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1022257928437.

Kajimura:2001:SIL

[137] S. Kajimura, K. Uchida, and E. G. Grau. Stimulation of insulin-like growth factor-i production by recombinant bovine growth hormone in Mozambique tilapia, Oreochromis mossambicus. Fish Physiology and Biochemistry, 25(3):221–230, December 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1022268811599.

Kim:2001:CTM

[138] J. B. Kim, B. A. Barton, and J. M. Conlon. Characterization of two molecular forms of vasoactive intestinal polypeptide (VIP) from the pallid sturgeon, Scaphirhynchus albus (Acipenseriformes). Fish Physiology and Biochemistry, 25(3):231–238, December 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/A:1022243304391.

Gallardo:2001:LAU

[139] M. A. Gallardo, C. Castejón, and J. Sánchez. L-leucine and L-alanine uptake by trout (Salmo trutta) cardiomyocytes: the effect of IGF-I and insulin. Fish Physiology and Biochemistry, 25(3):239–248, December 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1022285620321.

Mizuno:2001:ECA

[140] S. Mizuno, N. Misaka, and N. Kasahara. Effects of cortisol and angiotensin II on the number and size of juxtaglomerular cells in masu salmon, Oncorhynchus masou. Fish Physiology and Biochemistry, 25(3):249–254, December 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1022274327140.

Gunasekera:2001:CCF

[141] Rasanthi M. Gunasekera, Sena S. De Silva, and Brett A. Ingram. Chemical changes in fed and starved larval trout cod, Maccullochella macquarensis during early development. Fish Physiology and Biochemistry, 25(4): 255–268, December 2001. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1023247718139.

Peres:2001:EDP

[142] H. Peres and A. Oliva-Teles. Effect of dietary protein and lipid level on metabolic utilization of diets by European sea bass (Dicentrarchus labrax) juveniles. Fish Physiology and Biochemistry, 25(4):269–275, December 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1023239819048.

Hoffmayer:2001:PRC

[143] Eric R. Hoffmayer and Glenn R. Parsons. The physiological response to capture and handling stress in the Atlantic sharpnose shark, Rhizoprionodon terraenovae. Fish Physiology and Biochemistry, 25(4):277–285, December 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1023210620904.

Yavuzcan-Yildiz:2001:ESS

[144] H. Yavuzcan-Yıldız and M. Kırkavgaç-Uzbilek. The evaluation of secondary stress response of grass carp (Ctenopharyngodon idella, val. 1844) after exposing to the saline water. Fish Physiology and Biochemistry, 25 (4):287–290, December 2001. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1023279604975.

Applebaum:2001:CCA

[145] S. L. Applebaum, R. Perez, and G. J. Holt. Characterization of chymotrypsin activity during early ontogeny of larval red drum (Sciaenops ocellatus). Fish Physiology and Biochemistry, 25(4):291–300, December 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1023202219919.

Ng:2001:IDL

[146] Wing-Keong Ng, Phaik-Kin Lim, and Hadizah Sidek. The influence of a dietary lipid source on growth, muscle fatty acid composition and erythrocyte osmotic fragility of hybrid tilapia. Fish Physiology and Biochemistry, 25(4):301–310, December 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1023271901111.

Hellqvist:2001:LBF

[147] Anna Hellqvist, Monika Schmitz, and Bertil Borg. LH-β and FSH-β mRNA expression in nesting and post-breeding three-spined stickleback, Gasterosteus aculeatus, and effects of castration on expression of LH-β, FSH-β and spiggin mRNA. Fish Physiology and Biochemistry, 25(4): 311–317, December 2001. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1023210407163.

Schafhauser-Smith:2001:RPT

[148] D. Schafhauser-Smith and T. J. Benfey. The reproductive physiology of three age classes of adult female diploid and triploid brook trout (Salvelinus fontinalis). Fish Physiology and Biochemistry, 25(4):319–333, December 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1023285008072.

Sunde:2001:DPA

[149] J. Sunde, G. L. Taranger, and K. Rungruangsak-Torrissen. Digestive protease activities and free amino acids in white muscle as indicators for feed conversion efficiency and growth rate in Atlantic salmon (Salmo salar L.). Fish Physiology and Biochemistry, 25(4):335–345, December 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1023233024001.

Dussault:2001:ESA

[150] Ève B. Dussault, Richard C. Playle, and R. Scott McKinley. Effects of sublethal, acidic aluminum exposure on blood ions and metabolites, cardiac output, heart rate, and stroke volume of rainbow trout, Oncorhynchus mykiss. Fish Physiology and Biochemistry, 25(4):347–357, December 2001. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1023295413119.

Anonymous:2002:Pa

[151] Anonymous. Preface. Fish Physiology and Biochemistry, 26(1):1, January 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1023373006718.

Yoshizaki:2002:PGC

[152] Goro Yoshizaki, Yutaka Takeuchi, and Toshio Takeuchi. Primordial germ cells: the blueprint for a piscine life. Fish Physiology and Biochemistry, 26(1):3–12, January 2002. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1023388317621.

Strussmann:2002:MEE

[153] Carlos Augusto Strüssmann and Masaru Nakamura. Morphology, endocrinology, and environmental modulation of gonadal sex differentiation in teleost fishes. Fish Physiology and Biochemistry, 26(1):13–29, January 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1023343023556.

Okuzawa:2002:PT

[154] Koichi Okuzawa. Puberty in teleosts. Fish Physiology and Biochemistry, 26(1):31–41, January 2002. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1023395025374.

Schulz:2002:SER

[155] Rüdiger W. Schulz and Takeshi Miura. Spermatogenesis and its endocrine regulation. Fish Physiology and Biochemistry, 26(1):43–56, January 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1023303427191.

Patino:2002:OFG

[156] Reynaldo Patiño and Craig V. Sullivan. Ovarian follicle growth, maturation, and ovulation in teleost fish. Fish Physiology and Biochemistry, 26(1):57–70, January 2002. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1023311613987.

Kobayashi:2002:HPC

[157] Makito Kobayashi, Peter W. Sorensen, and Norm E. Stacey. Hormonal and pheromonal control of spawning behavior in the goldfish. Fish Physiology and Biochemistry, 26(1):71–84, January 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/A:1023375931734.

Blazer:2002:HAG

[158] V. S. Blazer. Histopathological assessment of gonadal tissue in wild fishes. Fish Physiology and Biochemistry, 26(1):85–101, January 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1023332216713.

Hellqvist:2002:EHC

[159] Anna Hellqvist, Ian Mayer, and Bertil Borg. Effects of hemi-castration on plasma steroid levels in two teleost fishes; the three-spined stickleback, Gasterosteus aculeatus, and the Atlantic salmon, Salmo salar. Fish Physiology and Biochemistry, 26(2):107–110, March 2002. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1025435625551.

Han:2002:MCS

[160] Yu-San Han and John Yuh-Lin Yu. Molecular cloning and sequence analysis of the cDNAs for pituitary glycoprotein hormone α subunits from two species of synbranchiformes, Monopterus albus and Ophisternon bengalense. Fish Physiology and Biochemistry, 26(2):111–120, March 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1025478031158.

Belanger:2002:DCC

[161] F. Bélanger, P. U. Blier, and J.-D. Dutil. Digestive capacity and compensatory growth in Atlantic cod (Gadus morhua). Fish Physiology and Biochemistry, 26(2):121–128, March 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1025461108348.

Callahan:2002:GCS

[162] Will Callahan, Hamish McWilliam, and Tes Toop. Guanylyl cyclase stimulation by homologous and heterologous natriuretic peptides in salmonids. Fish Physiology and Biochemistry, 26(2):129–138, March 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1025466000972.

Eriksen:2002:APA

[163] N. T. Eriksen. Accuracy and precision of aquatic respirometers with emphasis on monophase oxystats. Fish Physiology and Biochemistry, 26 (2):139–147, March 2002. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1025461900063.

Frasco:2002:EDB

[164] M. F. Frasco and L. Guilhermino. Effects of dimethoate and betanaphthoflavone on selected biomarkers of Poecilia reticulata. Fish Physiology and Biochemistry, 26(2):149–156, March 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1023/A:1025457831923.

Tocher:2002:EDC

[165] Douglas R. Tocher, Jorge Fonseca-Madrigal, and John R. Sargent. Effects of diets containing linseed oil on fatty acid desaturation and oxidation in hepatocytes and intestinal enterocytes in Atlantic salmon (Salmo salar). Fish Physiology and Biochemistry, 26(2):157–170, March 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1025416731014.

Davidson:2002:LLF

[166] Bruce Davidson and Geremy Cliff. The liver lipid fatty acid profiles of seven Indian Ocean shark species. Fish Physiology and Biochemistry, 26(2):171–175, March 2002. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1025447718625.

Fournier:2002:NUU

[167] V. Fournier, M. F. Gouillou-Coustans, and S. J. Kaushik. Nitrogen utilisation and ureogenesis as affected by dietary nucleic acid in rainbow trout (Oncorhynchus mykiss) and turbot (Psetta maxima). Fish Physiology and Biochemistry, 26(2):177–188, March 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1025465603604.

Katharios:2002:PIS

[168] P. Katharios, J. Iliopoulou-Georgudaki, and M. Pavlidis. Pharmacokinetics of ivermectin in sea bream, Sparus aurata using a direct competitive ELISA. Fish Physiology and Biochemistry, 26(2):189–195, March 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1025473904513.

Pinillos:2002:PRO

[169] M. L. Pinillos, A. I. Guijarro, and A. P. Scott. Production, release and olfactory detection of sex steroids by the tench (Tinca tinca L.). Fish Physiology and Biochemistry, 26(2):197–210, March 2002. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1025421920443.

Olsen:2002:ASA

[170] Rolf Erik Olsen, Kristina Sundell, and Einar Ringø. Acute stress alters the intestinal lining of Atlantic salmon, Salmo salar L.: an electron microscopical study. Fish Physiology and Biochemistry, 26(3):211–221, May 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1026217719534.

Riley:2002:DEG

[171] L. G. Riley, T. Hirano, and E. G. Grau. Disparate effects of gonadal steroid hormones on plasma and liver mRNA levels of insulin-like growth factor-i and vitellogenin in the tilapia, Oreochromis mossambicus. Fish Physiology and Biochemistry, 26(3):223–230, May 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/A:1026209502696.

Carton:2002:EST

[172] M. W. Carton, M. Schmitz, and L. Byrnes. Expressed sequence tags in larval Arctic charr (Salvelinus alpinus). Fish Physiology and Biochemistry, 26(3):231–238, May 2002. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1026254132468.

Rinchard:2002:CPS

[173] J. Rinchard, K. Dabrowski, and J. Ottobre. Changes of plasma steroid concentrations associated with spontaneous or induced ovulation in yellow perch Perca flavescens. Fish Physiology and Biochemistry, 26(3):239– 248, May 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1026242406796.

Ellestad:2002:IPD

[174] L. E. Ellestad, R. Angel, and J. H. Soares, Jr. Intestinal phytase i: Detection and preliminary characterization of activity in the intestinal brush border membrane of hybrid striped bass Morone saxatilis × M. chrysops. Fish Physiology and Biochemistry, 26(3):249–258, May 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1026279507705.

Ellestad:2002:IPI

[175] L. E. Ellestad, R. Angel, and J. H. Soares, Jr. Intestinal phytase II: A comparison of activity and in vivo phytate hydrolysis in three teleost species with differing digestive strategies. Fish Physiology and Biochemistry, 26(3):259–273, May 2002. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1026231624543.

Mittal:2002:CGS

[176] S. Mittal, Pinky, and A. K. Mittal. Characterisation of glycoproteins in the secretory cells in the operculum of an Indian hill stream fish Garra lamta (Hamilton) (Cyprinidae, Cypriniformes). Fish Physiology and Biochemistry, 26(3):275–288, May 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1026227523634.

Piros:2002:BCS

[177] B. Piros, J. Glogowski, and A. Ciereszko. Biochemical characterization of Siberian sturgeon Acipenser baeri and sterlet Acipenser ruthenus milt plasma and spermatozoa. Fish Physiology and Biochemistry, 26(3):289– 295, May 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1026280218957.

Mourente:2002:IPS

[178] G. Mourente and J. R. Dick. Influence of partial substitution of dietary fish oil by vegetable oils on the metabolism of [1-^{14}C] 18:3n − 3 in isolated hepatocytes of European sea bass (Dicentrarchus labrax L.). Fish Physiology and Biochemistry, 26(3):297–308, May 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/A:1026236418184.

Ehrman:2002:NRS

[179] Melissa M. Ehrman, Gregory T. Melroe, and Mark A. Sheridan. Nutritional regulation of somatostatin expression in rainbow trout, Oncorhynchus mykiss. Fish Physiology and Biochemistry, 26(4):309–314, July 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1023/B:FISH.
0000009261.30672.3a.

Xiao:2002:LGH

[180] Dong Xiao, Anderson O. L. Wong, and Hao-Ran Lin. Lack of growth hormone-releasing peptide-6 action on in vivo and in vitro growth hormone secretion in sexually immature grass carp (Ctenopharyngodon idellus). Fish Physiology and Biochemistry, 26(4):315–327, July 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1023/B:FISH.
0000009260.34930.12.

Wood:2002:NMP

[181] C. M. Wood, B. Eletti, and P. Pärt. New methods for the primary culture of gill epithelia from freshwater rainbow trout. Fish Physiology and Biochemistry, 26(4):329–344, July 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/B:FISH.0000009262.45438.79.

Rodriguez:2002:OJS

[182] A. Rodríguez, M. A. Gallardo, and F. Castelló-Orvay. Osmoregulation in juvenile Siberian sturgeon (Acipenser baerii). Fish Physiology and Biochemistry, 26(4):345–354, July 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/B:FISH.0000009263.83075.68.

Hiramatsu:2002:IPC

[183] N. Hiramatsu, T. Matsubara, and C. V. Sullivan. Identification, purification and classification of multiple forms of vitellogenin from white perch (Morone americana). Fish Physiology and Biochemistry, 26(4):355–370, July 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1023/B:FISH.
0000009266.58556.9a.

Jentoft:2002:OCS

[184] Sissel Jentoft, James A. Held, and Terence P. Barry. Ontogeny of the cortisol stress response in yellow perch (Perca flavescens). Fish Physiology and Biochemistry, 26(4):371–378, July 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/B:FISH.0000009276.05161.8d.

Saha:2002:SCI

[185] N. R. Saha, T. Usami, and Y. Suzuki. Seasonal changes in the immune activities of common carp (Cyprinus carpio). Fish Physiology and Biochemistry, 26(4):379–387, July 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/B:FISH.0000009275.25834.67.

Gelsleichter:2002:SSC

[186] J. Gelsleichter, L. E. L. Rasmussen, and L. Lombardi-Carlson. Serum steroid concentrations and development of reproductive organs during puberty in male bonnethead sharks, Sphyrna tiburo. Fish Physiology and Biochemistry, 26(4):389–401, July 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/B:FISH.0000009292.70958.65.

Anonymous:2002:VCa

[187] Anonymous. Volume contents. Fish Physiology and Biochemistry, 26(4): 403–405, July 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/B:
FISH.0000008699.12569.13.

Anonymous:2002:IA

[188] Anonymous. Index of authors. Fish Physiology and Biochemistry, 26(4): 406, July 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/B:
FISH.0000008700.32555.02.

Szisch:2002:PCC

[189] V. Szisch, A. L. van der Salm, and M. Pavlidis. Physiological colour changes in the red porgy, Pagrus pagrus, following adaptation to blue lighting spectrum. Fish Physiology and Biochemistry, 27(1-2):1–8, September 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/B:
FISH.0000021719.65318.5a.

vanderSalm:2002:EVE

[190] A. L. van der Salm, D. T. Nolan, and S. E. Wendelaar Bonga. In vitro evidence that cortisol directly modulates stress-related responses in the skin epidermis of the rainbow trout (Oncorhynchus mykiss Walbaum). Fish Physiology and Biochemistry, 27(1-2):9–18, September 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/B:FISH.0000021741.05086.55.

Hill:2002:EFA

[191] J. V. Hill, W. Davison, and M. E. Forster. The effects of fish anaesthetics (MS222, metomidate and AQUI-s) on heart ventricle, the cardiac vagus and branchial vessels from Chinook salmon (Oncorhynchus tshawytscha). Fish Physiology and Biochemistry, 27(1-2):19–28, September 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1023/B:FISH.
0000021742.30567.2d.

Lochmann:2002:CRG

[192] R. T. Lochmann, K. B. Davis, and B. A. Simco. Cortisol response of golden shiners (Notemigonus crysoleucas) fed diets differing in lipid content. Fish Physiology and Biochemistry, 27(1-2):29–34, September 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/B:FISH.0000021772.86442.08.

Roy:2002:EDP

[193] P. K. Roy, P. E. Witten, and S. P. Lall. Effects of dietary phosphorus on bone growth and mineralisation of vertebrae in haddock (Melanogrammus aeglefinus L.). Fish Physiology and Biochemistry, 27(1-2):35–48, September 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1023/B:FISH.
0000021778.99011.ce.

Lahnsteiner:2002:IOF

[194] F. Lahnsteiner. The influence of ovarian fluid on the gamete physiology in the Salmonidae. Fish Physiology and Biochemistry, 27(1-2):49–59, September 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/B:
FISH.0000021792.97913.2e.

Ribeiro:2002:DEP

[195] L. Ribeiro, J. L. Zambonino-Infante, and M. T. Dinis. Digestive enzymes profile of Solea senegalensis post larvae fed Artemia and a compound diet. Fish Physiology and Biochemistry, 27(1-2):61–69, September 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1023/B:FISH.
0000021817.98363.47.

Moren:2002:CBC

[196] Mari Moren, Tore Næss, and Kristin Hamre. Conversion of β-carotene, canthaxanthin and astaxanthin to vitamin A in Atlantic halibut (Hippoglossus hippoglossus L.) juveniles. Fish Physiology and Biochemistry, 27 (1-2):71–80, September 2002. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/B:FISH.0000021819.46235.12.

Ghioni:2002:MSA

[197] C. Ghioni, A. E. A. Porter, and D. R. Tocher. Metabolism of 18:4n − 3 (stearidonic acid) and 20:4n−3 in salmonid cells in culture and inhibition of the production of prostaglandin F_{2α} (PGF_{2α}) from 20:4n − 6 (arachidonic acid). Fish Physiology and Biochemistry, 27(1-2):81–96, September 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1023/B:FISH.
0000021866.78048.45.

Green:2002:ODEa

[198] J. A. Green and R. W. Hardy. The optimum dietary essential amino acid pattern for rainbow trout (Oncorhynchus mykiss), to maximize nitrogen retention and minimize nitrogen excretion. Fish Physiology and Biochemistry, 27(1-2):97–108, September 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/B:FISH.0000021878.81647.6e.

Green:2002:ODEb

[199] J. A. Green, R. W. Hardy, and E. L. Brannon. The optimum dietary essential: nonessential amino acid ratio for rainbow trout (Oncorhynchus mykiss), which maximizes nitrogen retention and minimizes nitrogen excretion. Fish Physiology and Biochemistry, 27(1-2):109–115, September 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1023/B:FISH.
0000021877.49848.8f.

Petkam:2002:EVM

[200] R. Petkam, R. L. Renaud, and J. F. Leatherland. In vitro metabolism of progesterone, androgens and estrogens by rainbow trout embryos. Fish Physiology and Biochemistry, 27(1-2):117–128, September 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/B:FISH.0000021867.71729.22.

Al-Salahy:2002:SPS

[201] M. B. Al-Salahy. Some physiological studies on the effect of onion and garlic juices on the fish, Clarias lazera. Fish Physiology and Biochemistry, 27(1-2):129–142, September 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/B:FISH.0000021913.60189.76.

Anonymous:2002:C

[202] Anonymous. Corrigendum. Fish Physiology and Biochemistry, 27(1-2):143, September 2002. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/B:FISH.0000022019.00098.47.

vandenThillart:2002:ARL

[203] Guido van den Thillart, Gerjanne Vianen, and Johan Zaagsma. Adrenergic regulation of lipid mobilization in fishes; a possible role in hypoxia survival. Fish Physiology and Biochemistry, 27(3-4):??, December 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1023/B:FISH.
0000032754.42056.cb.

Anonymous:2002:Pb

[204] Anonymous. Preface. Fish Physiology and Biochemistry, 27(3-4):145–146, December 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/B:
FISH.0000032838.65275.bf.

Shi:2002:IOS

[205] F. T. Shi, W. S. Li, and H. R. Lin. IGF-i of orange spotted grouper Epinephelus coioides: cDNA cloning, sequencing and expression in Escherichia coli. Fish Physiology and Biochemistry, 27(3-4):147–156, December 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1023/B:FISH.
0000032722.06647.dd.

Nickerson:2002:EVR

[206] J. G. Nickerson, G. Drouin, and T. W. Moon. In vitro regulation of βadrenoceptor signaling in the rainbow trout, Oncorhynchus mykiss. Fish Physiology and Biochemistry, 27(3-4):157–171, December 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/B:FISH.0000032723.78349.4e.

Silverstein:2002:UGV

[207] Jeffrey T. Silverstein. Using genetic variation to understand control of feed intake in fish. Fish Physiology and Biochemistry, 27(3-4):173–178, December 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/B:
FISH.0000032724.36866.ce.

Laiz-Carrion:2002:EMF

[208] Raúl Laiz-Carrión, Susana Sangiao-Alvarellos, and Juan M. Mancera. Energy metabolism in fish tissues related to osmoregulation and cortisol action. Fish Physiology and Biochemistry, 27(3-4):179–188, December 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1023/B:FISH.
0000032725.96481.b8.

Navarro:2002:IIG

[209] I. Navarro, P. Rojas, and J. Gutiérrez. Insights into insulin and glucagon responses in fish. Fish Physiology and Biochemistry, 27(3-4):205–216, December 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1023/B:FISH.
0000032726.78074.04.

Very:2002:RSR

[210] N. M. Very and M. A. Sheridan. The role of somatostatins in the regulation of growth in fish. Fish Physiology and Biochemistry, 27(3-4):217–226, December 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1023/B:FISH.
0000032727.75493.e8.

Bjornsson:2002:GHE

[211] Björn Thrandur Björnsson, Viktoria Johansson, and Elisabeth Jönsson. Growth hormone endocrinology of salmonids: Regulatory mechanisms and mode of action. Fish Physiology and Biochemistry, 27(3-4):227–242, December 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1023/B:FISH.
0000032728.91152.10.

Perez-Sanchez:2002:OFG

[212] J. Pérez-Sánchez, J. A. Calduch-Giner, and M. M. Valdivia. Overview of fish growth hormone family. New insights in genomic organization and heterogeneity of growth hormone receptors. Fish Physiology and Biochemistry, 27(3-4):243–258, December 2002. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/B:FISH.0000032729.72746.c8.

Anonymous:2002:VCb

[213] Anonymous. Volume contents. Fish Physiology and Biochemistry, 27(3-4):259–260, December 2002. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/B:FISH.0000032839.10536.38.

Anonymous:2002:AI

[214] Anonymous. Authors-index. Fish Physiology and Biochemistry, 27(3-4):261, December 2002. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/B:FISH.0000032840.82651.8e.

Anonymous:2003:PNR

[215] Anonymous. Professor Niall R. Bromage 4-7-1942 to 8-5-2003. Fish Physiology and Biochemistry, 28(1-4):1–2, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/B:FISH.0000030563.16667.90.

Thomas:2003:RNS

[216] Peter Thomas. Rapid, nongenomic steroid actions initiated at the cell surface: lessons from studies with fish. Fish Physiology and Biochemistry, 28(1-4):3–12, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/B:FISH.0000030461.35242.57.

Parhar:2003:GRH

[217] Ishwar S. Parhar. Gonadotropin-releasing hormone receptors: neuroendocrine regulators and neuromodulators. Fish Physiology and Biochemistry, 28(1-4):13–18, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/B:FISH.0000030462.10997.24.

Anonymous:2003:F

[218] Anonymous. Foreword. Fish Physiology and Biochemistry, 28(1-4):17, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1023/B:FISH.
0000030784.21024.d5.

Amano:2003:GSM

[219] Masafumi Amano, Kataaki Okubo, and Kunio Yamamori. GnRH systems in masu salmon and barfin flounder. Fish Physiology and Biochemistry, 28(1-4):19–22, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/B:FISH.0000030463.33250.d2.

Okubo:2003:GGP

[220] K. Okubo and K. Aida. GnRH gene products downregulate their neighboring genes encoding protein tyrosine phosphatases. Fish Physiology and Biochemistry, 28(1-4):23–24, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/B:FISH.0000030464.70887.bd.

Klenke:2003:GRG

[221] U. Klenke and Y. Zohar. Gonadal regulation of gonadotropin subunit expression and pituitary Lh protein content in female hybrid striped bass. Fish Physiology and Biochemistry, 28(1-4):25–27, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/B:FISH.0000030465.56876.97.

Kazeto:2003:AEZ

[222] Y. Kazeto, R. Goto-Kazeto, and J. M. Trant. Aromatase expression in zebrafish and channel catfish brains: changes in transcript abundance associated with the reproductive cycle and exposure to endocrine disrupting chemicals. Fish Physiology and Biochemistry, 28(1-4):29–32, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1023/B:FISH.
0000030466.23085.94.

Joy:2003:CCL

[223] K. P. Joy and R. Chaube. Catecholaminergic control of LH secretion: E_{2} modulation of tyrosine hydroxylase activity in female catfish Heteropneustes fossilis. Fish Physiology and Biochemistry, 28(1-4):33–35, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1023/B:FISH.
0000030467.61134.d0.

Bayarri:2003:EPM

[224] M. J. Bayarri, L. Rodríguez, and M. Carrillo. Effect of photoperiod manipulation on daily rhythms of melatonin and reproductive hormones in caged European sea bass (Dicentrarchus labrax). Fish Physiology and Biochemistry, 28(1-4):37–38, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/B:FISH.0000030468.98771.c0.

Guilgur:2003:CTG

[225] L. G. Guilgur, L. A. Miranda, and G. M. Somoza. Characterization of three GnRH cDNA sequences in the pejerrey fish Odontesthes bonariensis. Fish Physiology and Biochemistry, 28(1-4):39–40, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/B:FISH.0000030470.32151.07.

Ikemoto:2003:EMI

[226] T. Ikemoto, Y. Oka, and M. K. Park. Existence of multiple isoforms of GnRH ligands and receptors in the dwarf gourami, Colisa lalia. Fish Physiology and Biochemistry, 28(1-4):41–42, March 2003. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/B:FISH.0000030469.80158.a6.

Kawaguchi:2003:ELC

[227] N. Kawaguchi, K. Okubo, and K. Aida. The expression and localization of corticotropin-releasing hormone and urotensin I transcripts in the Japanese eel, Anguilla japonica. Fish Physiology and Biochemistry, 28 (1-4):43–44, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/B:FISH.0000030471.64294.6c.

Lethimonier:2003:ETD

[228] C. Lethimonier, O. Kah, and J.-J. Lareyre. Evidence for two distinct GnRH receptor types expressed in teleosts. Fish Physiology and Biochemistry, 28(1-4):45–46, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/B:FISH.0000030472.79953.47.

Nocillado:2003:MTS

[229] J. N. Nocillado, H. K. Chua, and A. Elizur. Molecular tools for studying puberty in the grey mullet, Mugil cephalus: cloning of cDNAs that regulate reproductive function. Fish Physiology and Biochemistry, 28(1-4): 47–48, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/B:
FISH.0000030473.61547.b0.

Ogawa:2003:MGS

[230] S. Ogawa, T. Soga, and I. S. Parhar. Modulation of GnRH subtypes by social stress and aggressive behavior. Fish Physiology and Biochemistry, 28(1-4):49–50, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/B:FISH.0000030474.32151.12.

Strobl-Mazzulla:2003:NDP

[231] P. H. Strobl-Mazzulla, G. C. López, and G. M. Somoza. Neuroanatomical distribution and partial sequence of cDNA encoding brain cytochrome P450 aromatase (P450aromB) in pejerrey Odontesthes bonariensis. Fish Physiology and Biochemistry, 28(1-4):51–52, March 2003. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/B:FISH.0000030475.35723.33.

Swanson:2003:BPF

[232] P. Swanson, J. T. Dickey, and B. Campbell. Biochemistry and physiology of fish gonadotropins. Fish Physiology and Biochemistry, 28(1-4):53–59, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1023/B:FISH.
0000030476.73360.07.

Ando:2003:RLS

[233] H. Ando, P. Swanson, and A. Urano. Regulation of LH synthesis and release by GnRH and gonadal steroids in masu salmon. Fish Physiology and Biochemistry, 28(1-4):61–63, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/B:FISH.0000030477.87920.2d.

Ge:2003:ASG

[234] Wei Ge, Nga-Ling Ko, and Katsumi Aida. Activin stimulates goldfish FSH biosynthesis by enhancing FSHβ promoter activity. Fish Physiology and Biochemistry, 28(1-4):65–71, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/B:FISH.0000030478.31396.bb.

Berta:2003:CID

[235] Levavi-Sivan Berta, Avitan Ayelet, and Kanias Tamir. Characterization of the inhibitory dopamine receptor from the pituitary of tilapia. Fish Physiology and Biochemistry, 28(1-4):73–75, March 2003. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/B:FISH.0000030479.47055.24.

Gen:2003:PRF

[236] Koichiro Gen, Sonoko Yamaguchi, and Hirohiko Kagawa. Physiological roles of FSH and LH in red seabream, Pagrus major. Fish Physiology and Biochemistry, 28(1-4):77–80, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/B:FISH.0000030480.97947.ba.

Rahman:2003:GGR

[237] M. A. Rahman, K. Ohta, and M. Matsuyama. Gonadotropins, gonadotropin receptors and their expressions during sexual maturation in yellowtail, a carangid fish. Fish Physiology and Biochemistry, 28(1-4):81– 83, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/B:
FISH.0000030481.94719.a0.

Nozaki:2003:IDG

[238] M. Nozaki, Y. Oshima, and S. A. Sower. Immunohistochemical detection of gonadotropin-like material in the hagfish pituitary, correlated with their gonadal conditions. Fish Physiology and Biochemistry, 28(1-4):85– 87, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/B:
FISH.0000030482.32357.cd.

Choi:2003:GMT

[239] E. J. Choi, D. H. Jin, and Y. C. Sohn. Gonadotropins in the Manchurian trout, Brachymystax lenok tsinlingensis. Fish Physiology and Biochemistry, 28(1-4):89–90, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/B:FISH.0000030483.11026.e5.

Chyb:2003:DIR

[240] J. Chyb, M. Sokolowska-Mikolajczyk, and P. Epler. Differential impact of recombinant human activin on LH secretion in male and female common carp during wintering and spawning period. Fish Physiology and Biochemistry, 28(1-4):91–92, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/B:FISH.0000030484.33347.df.

Hellqvist:2003:CSF

[241] A. Hellqvist, C. Bornestaf, and M. Schmitz. Cloning and sequencing of FSH-β and LH-β subunits and changes in their expression during the reproductive cycle in the three-spined stickleback, Gasterosteus aculeatus. Fish Physiology and Biochemistry, 28(1-4):93–94, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/B:FISH.0000030485.70984.f2.

Ishii:2003:GDE

[242] S. Ishii, Y. Yoshiura, and K. Aida. Gonadal development and expression profiles of gonadotropin genes in wild sea conger, Ariosoma meeki. Fish Physiology and Biochemistry, 28(1-4):95–96, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/B:FISH.0000030486.07523.84.

Kamei:2003:ASA

[243] H. Kamei, T. Oshira, and K. Aida. Androgen secretion activity of recombinant follicle-stimulating hormone of Japanese eel, Anguilla japonica in immature and maturing eel testes. Fish Physiology and Biochemistry, 28(1-4):97–98, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/B:FISH.0000030487.74762.c7.

Kumakura:2003:EGR

[244] N. Kumakura, K. Okuzawa, and H. Kagawa. Effects of gonadotropinreleasing hormone agonist and testosterone on gonadotropin subunit gene expression in female red seabream (Pagrus major). Fish Physiology and Biochemistry, 28(1-4):99–100, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/B:FISH.0000030488.26754.1e.

Miranda:2003:MCC

[245] L. A. Miranda, L. G. Guilgur, and G. M. Somoza. Molecular cloning of cDNAs encoding FSH-β and LH-β subunits in the pejerrey fish, Odontesthes bonariensis. Fish Physiology and Biochemistry, 28(1-4):101–102, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1023/B:FISH.
0000030489.54129.fe.

Naito:2003:SGS

[246] N. Naito, Y. Koide, and H. Kawauchi. Salmon gonadotropes secrete free form alpha (α)-subunit: quantitative analysis by the sandwich cell immunoblot assay. Fish Physiology and Biochemistry, 28(1-4):103–108, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1023/B:FISH.
0000030490.89912.f6.

Schmitz:2003:DEI

[247] M. Schmitz. Differential effect of insulin-like growth factor I on in vitro gonadotropin subunits expression in Atlantic salmon. Fish Physiology and Biochemistry, 28(1-4):105–106, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/B:FISH.0000030491.40805.96.

Shein:2003:MCG

[248] Ni Lar Shein, M. Takushima, and K. Soyano. Molecular cloning of gonadotropin cDNA in sevenband grouper, Epinephelus septemfasciatus. Fish Physiology and Biochemistry, 28(1-4):107–108, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/B:FISH.0000030492.22399.39.

Shimizu:2003:IIG

[249] A. Shimizu, H. Kagawa, and H. Tanaka. Immunocytochemical identification of gonadotrophs (FSH cells and LH cells) in various perciform fishes using antisera raised against synthetic peptides. Fish Physiology and Biochemistry, 28(1-4):109–110, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/B:FISH.0000030493.74321.6d.

Yamaguchi:2003:RGS

[250] S. Yamaguchi, K. Gen, and M. Matsuyama. Regulation of gonadotropin subunit genes expression by 11-ketotestosterone during early spermatogenesis in male red seabream, Pagrus major. Fish Physiology and Biochemistry, 28(1-4):111–112, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/B:FISH.0000030494.13499.06.

Nakamura:2003:REP

[251] Masaru Nakamura, Ramji K. Bhandari, and Mikihiko Higa. The role estrogens play in sex differentiation and sex changes of fish. Fish Physiology and Biochemistry, 28(1-4):113–117, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/B:FISH.0000030495.99487.17.

Baron:2003:GED

[252] D. Baron and Y. Guiguen. Gene expression during gonadal sex differentiation in rainbow trout (Oncorhynchus mykiss): from candidate genes studies to high throughout genomic approach. Fish Physiology and Biochemistry, 28(1-4):119–123, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/B:FISH.0000030496.21741.6c.

Maack:2003:OSD

[253] G. Maack, H. Segner, and C. R. Tyler. Ontogeny of sexual differentiation in different strains of zebrafish (Danio rerio). Fish Physiology and Biochemistry, 28(1-4):125–128, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/B:FISH.0000030497.59378.88.

Strussmann:2003:PUO

[254] C. A. Strüssmann, T. Oikawa, and S. Kasuga. Potential use of otolith microchemistry for field studies of temperature-dependent sex determination and gonadal degeneration in fish. Fish Physiology and Biochemistry, 28(1-4):129–130, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/B:FISH.0000030498.33279.69.

Du:2003:EAS

[255] Jin-Lien Du, Bih-Yun Lin, and Ching-Fong Chang. Estradiol, aromatase and steroid receptors involved in the sex change of protandrous black porgy, Acanthopagrus schlegeli. Fish Physiology and Biochemistry, 28 (1-4):131–133, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/B:FISH.0000030499.77509.23.

Matsuda:2003:SDG

[256] Masaru Matsuda, Yoshitaka Nagahama, and Mitsuru Sakaizumi. The sex determining gene of medaka: a Y-specific DM domain gene (DMY) is required for male development. Fish Physiology and Biochemistry, 28 (1-4):135–139, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/B:FISH.0000030500.29914.7a.

Bhandari:2003:TAI

[257] Ramji K. Bhandari, Mikihiko Higa, and Masaru Nakamura. Treatment with an aromatase inhibitor induces complete sex change in the protogynous honeycomb grouper (Epinephelus merra). Fish Physiology and Biochemistry, 28(1-4):141–142, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/B:FISH.0000030501.23595.2e.

Fan:2003:MCT

[258] H. G. Fan, D. S. Wang, and Y. Nagahama. Molecular cloning of the three gonadotropin subunits and early expression of FSHβ during sex differentiation in the Nile tilapia, Oreochromis niloticus. Fish Physiology and Biochemistry, 28(1-4):143–144, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/B:FISH.0000030502.48938.c0.

Fernandino:2003:MCS

[259] J. L. Fernandino, L. G. Guilgur, and G. M. Somoza. Molecular cloning of SOX9, DMRT1 and SF1 cDNA partial sequences in the pejerrey fish Odontesthes bonariensis (Atheriniformes). Fish Physiology and Biochemistry, 28(1-4):145–146, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/B:FISH.0000030503.86575.f3.

Gardner:2003:SCS

[260] L. Gardner, T. Anderson, and A. Elizur. Sex change strategy and the aromatase genes. Fish Physiology and Biochemistry, 28(1-4):147–148, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1023/B:FISH.
0000030504.68522.53.

Higa:2003:RSH

[261] Mikihiko Higa, Kei Ogasawara, and Masaru Nakamura. Role of steriod hormones in sex change of protogynous wrasse. Fish Physiology and Biochemistry, 28(1-4):149–150, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/B:FISH.0000030505.28138.d1.

Ikeuchi:2003:TCL

[262] T. Ikeuchi, T. Kobayashi, and Y. Nagahama. Transgenic cell lines which stably express progestogen receptors (PRS) α and β and the PRresponsive reporter genes. Fish Physiology and Biochemistry, 28(1-4):151, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1023/B:FISH.
0000030506.43797.fd.

Kajiura-Kobayashi:2003:EIS

[263] H. Kajiura-Kobayashi, T. Kobayashi, and Y. Nagahama. Estrogeninduced and suppressed gene expression during XY sex reversal in a teleost fish, tilapia. Fish Physiology and Biochemistry, 28(1-4):153, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1023/B:FISH.
0000030507.80335.4e.

Kato:2003:GSD

[264] K. Kato, S. Miyashita, and H. Kumai. Gonadal sex differentiation and sex control in red sea bream, Pagrus major. Fish Physiology and Biochemistry, 28(1-4):155–156, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/B:FISH.0000030508.83907.e6.

Sun:2003:DEL

[265] B. Sun, N. W. Pankhurst, and M. Watts. Development of an enzymelinked immunosorbent assay (ELISA) for vitellogenin measurement in greenback flounder Rhombosolea tapirina. Fish Physiology and Biochemistry, 29(1):13–21, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/B:FISH.0000035890.17372.75.

Lund:2003:PEO

[266] S. G. Lund and B. L. Tufts. The physiological effects of heat stress and the role of heat shock proteins in rainbow trout (Oncorhynchus f ) red blood cells. Fish Physiology and Biochemistry, 29(1):1–12, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/B:FISH.0000035891.54116.34.

Piercy:2003:MHC

[267] A. Piercy, J. Gelsleichter, and F. F. Snelson, Jr. Morphological and histological changes in the genital ducts of the male Atlantic stingray, Dasyatis sabina, during the seasonal reproductive cycle. Fish Physiology and Biochemistry, 29(1):23–35, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/B:FISH.0000035894.45256.59.

Congleton:2003:PIL

[268] J. L. Congleton, P. R. Biga, and B. C. Peterson. Plasma insulin-like growth factor-i concentrations in yearling Chinook salmon (Oncorhynchus tshawytscha) migrating from the Snake River Basin, USA. Fish Physiology and Biochemistry, 29(1):57–66, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/B:FISH.0000035895.21012.61.

Woolsey:2003:APS

[269] J. Woolsey and R. L. Ingermann. Acquisition of the potential for sperm motility in steelhead (Oncorhynchus mykiss): effect of pH on dynein ATPase. Fish Physiology and Biochemistry, 29(1):47–56, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1023/B:FISH.
0000035896.84954.f6.

Furuita:2003:BCE

[270] H. Furuita, H. Ohta, and T. Yamamoto. Biochemical composition of eggsin relation to egg quality in the Japanese eel, Anguilla japonica. Fish Physiology and Biochemistry, 29(1):37–46, March 2003. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/B:FISH.0000035897.58924.9d.

Anonymous:2003:Ca

[271] Anonymous. Corrigendum. Fish Physiology and Biochemistry, 29(1):87– 90, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/B:
FISH.0000035900.90655.e1.

Guo:2003:BHA

[272] Huarong Guo, Bing Huang, and Fei Qi. Biochemical and histochemical activities of tyrosinase in the skins of normal and albino turbot Scophthalmus maximus. Fish Physiology and Biochemistry, 29(1):67–76, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1023/B:FISH.
0000035901.74653.f4.

Kang:2003:ETT

[273] Duk-Young Kang and Robert H. Devlin. Effects of 3,5,3′-triiodo-l-thyronine (T_{3}) and 6-n-propyl-2-thiouracil (PTU) on growth of GHtransgenic Coho salmon, Oncorhynchus kitsutch. Fish Physiology and Biochemistry, 29(1):77–85, March 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/B:FISH.0000035903.77056.5c.

Tong:2003:MCI

[274] Fudan Tong, Hongyun Liu, and Qun Wang. Molecular cloning of insulinlike growth factor-i from triangular bream (Megalobrama terminalis). Fish Physiology and Biochemistry, 29(2):91–95, May 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1023/B:FISH.0000035889.36190.f7.

Achuba:2003:PIF

[275] F. I. Achuba and S. A. Osakwe. Petroleum-induced free radical toxicity in African catfish (Clarias gariepinus). Fish Physiology and Biochemistry, 29(2):97–103, May 2003. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/B:FISH.0000035905.14420.eb.

Sala-Rabanal:2003:ELT

[276] M. Sala-Rabanal, J. Sánchez, and M. A. Gallardo. Effects of low temperatures and fasting on hematology and plasma composition of gilthead sea bream (Sparus aurata). Fish Physiology and Biochemistry, 29(2):105–115, May 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1023/B:FISH.
0000035904.16686.b6.

Snyder:2003:CTH

[277] Randal J. Snyder and Todd M. Hennessey. Cold tolerance and homeoviscous adaptation in freshwater alewives (Alosa pseudoharengus). Fish Physiology and Biochemistry, 29(2):117–126, May 2003. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/B:FISH.0000035920.60817.11.

Yousefian:2003:BCE

[278] Mehdi Yousefian and Ahmet Erkan Kideys. Biochemical composition of Mnemiopsis leidyi in the southern Caspian Sea. Fish Physiology and Biochemistry, 29(2):127–131, May 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/B:FISH.0000035921.70639.fb.

Bendiksen:2003:ETF

[279] E. Å. Bendiksen and M. Jobling. Effects of temperature and feed composition on essential fatty acid (n-3 and n-6) retention in Atlantic salmon (Salmo salar L.) parr. Fish Physiology and Biochemistry, 29(2):133–140, May 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1023/B:FISH.
0000035937.68098.83.

Dayal:2003:NAA

[280] J. Syama Dayal, S. Ahamad Ali, and R. Subburaj. Nutrient and amino acid profiles of egg and larvae of Asian seabass, Lates calcarifer (Bloch). Fish Physiology and Biochemistry, 29(2):141–147, May 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/B:FISH.0000035936.93304.c2.

Selvakumar:2003:TMP

[281] S. Selvakumar and P. Geraldine. Thermal modulation of pyruvate metabolism in the freshwater prawn Macrobrachium malcolmsonii: the role of lactate dehydrogenase. Fish Physiology and Biochemistry, 29(2): 149–157, May 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/B:
FISH.0000035935.33689.ea.

Viant:2003:CBH

[282] M. R. Viant, I. Werner, and M. L. Johnson. Correlation between heat-shock protein induction and reduced metabolic condition in juvenile steelhead trout (Oncorhynchus mykiss) chronically exposed to elevated temperature. Fish Physiology and Biochemistry, 29(2):159–171, May 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1023/B:FISH.
0000035938.92027.81.

Afonso:2003:SRD

[283] L. O. B. Afonso, N. Basu, and G. K. Iwama. Sex-related differences in the organismal and cellular stress response in juvenile salmon exposed to treated bleached kraft mill effluent. Fish Physiology and Biochemistry, 29(2):173–179, May 2003. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/B:FISH.0000035939.81588.09.

Olsen:2003:ESO

[284] Rolf Erik Olsen, Bjørn Tore Dragnes, and Einar Ringø. Effect of soybean oil and soybean lecithin on intestinal lipid composition and lipid droplet accumulation of rainbow trout, Oncorhynchus mykiss Walbaum. Fish Physiology and Biochemistry, 29(3):181–192, July 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/B:FISH.0000045708.67760.43.

Tocher:2003:EDL

[285] D. R. Tocher, J. G. Bell, and J. Fonseca-Madrigal. Effects of dietary lipid level and vegetable oil on fatty acid metabolism in Atlantic salmon (Salmo salar L.) over the whole production cycle. Fish Physiology and Biochemistry, 29(3):193–209, July 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/B:FISH.0000045722.44186.ee.

Hirayama:2003:PSG

[286] M. Hirayama, M. Nakaniwa, and S. Watabe. Primary structures and gene organizations of two types of Wap65 from the pufferfish Takifugu rubripes. Fish Physiology and Biochemistry, 29(3):211–224, July 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/B:FISH.0000045723.52428.5e.

Welker:2003:RBD

[287] Thomas L. Welker and James L. Congleton. Relationship between dietary lipid source, oxidative stress, and the physiological response to stress in sub-yearling Chinook salmon (Oncorhynchus tshawytscha). Fish Physiology and Biochemistry, 29(3):225–235, July 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/B:FISH.0000045724.93156.ef.

Itoi:2003:CMF

[288] S. Itoi, S. Kawahara, and S. Watabe. Changes in mitochondrial fatty acid composition following temperature acclimation of carp and their possible effects on F_{o} F_{1}-ATPase activity. Fish Physiology and Biochemistry, 29(3):237–244, July 2003. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/B:FISH.0000045721.26261.a4.

Holen:2003:TOL

[289] E. Holen and K. Hamre. Towards obtaining long term embryonic stem cell like cultures from a marine flatfish, Scophtalmus maximus. Fish Physiology and Biochemistry, 29(3):245–252, July 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/B:FISH.0000045725.01192.44.

Anonymous:2003:Cb

[290] Anonymous. Corrigendum. Fish Physiology and Biochemistry, 29(3):253, July 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1023/B:FISH.
0000045803.90859.d5.

Horiguchi:2003:IPK

[291] R. Horiguchi, M. Yoshikuni, and T. Tokumoto. Identification of a protein kinase which phosphorylate α 4 subunit of the 26S proteasome in goldfish oocytes. Fish Physiology and Biochemistry, 29(3):255–256, July 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1023/B:FISH.
0000045709.15082.b5.

Zhou:2003:DBS

[292] L. R. Zhou, E. E. Deane, and N. Y. S. Woo. Development of a black sea bream fibroblast cell line and its potential use as an in vitro model for stress protein studies. Fish Physiology and Biochemistry, 29(4): 255–262, December 2003. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-004-5262-7.

Kim:2003:ATP

[293] B. H. Kim, M. S. Rahman, and A. Takemura. Alcohol treatment promotes conversion of testosterone to estradiol-17β in female tilapia, Oreochromis mossambicus. Fish Physiology and Biochemistry, 29(4):263–268, December 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
004-3268-9.

Manning:2003:CAO

[294] A. J. Manning and M. P. M. Burton. Cytological abnormalities in the ovaries of triploid yellowtail flounder, Limanda ferruginea (Storer). Fish Physiology and Biochemistry, 29(4):269–273, December 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-004-9273-1.

Shimizu:2003:UAI

[295] A. Shimizu, T. Sakai, and H. Honda. Universal antisera for immunocytochemical identification of two different gonadotrophs in acanthopterygian fishes. Fish Physiology and Biochemistry, 29(4):275–287, December 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-004-5287-y.

Shiozaki:2003:MEH

[296] K. Shiozaki, T. Nakano, and M. Sato. Metabolism of exogenous histamine in rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 29(4):289–295, December 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-004-9295-8.

Orozco:2003:TPO

[297] A. Orozco, B. Fenton, and C. Valverde-R. Temporal profile of the outerand inner-ring iodothyronine deiodinase pathways in the liver and skin in developing rainbow trout Oncorhynchus mykiss. Fish Physiology and Biochemistry, 29(4):297–303, December 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-004-7303-7.

Halamickova:2003:CMC

[298] A. Halamíčková, L. Vorlová, and H. Buchtová. Comparison of malondialdehyde concentrations in different muscle areas of tench Tinca tinca L. Fish Physiology and Biochemistry, 29(4):305–312, December 2003. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-004-5312-1.

Ibarz:2003:OCF

[299] A. Ibarz, J. Fernández-Borràs, and J. Sánchez. Oxygen consumption and feeding rates of gilthead sea bream (Sparus aurata) reveal lack of acclimation to cold. Fish Physiology and Biochemistry, 29(4):313–321, December 2003. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
004-3321-8.

Yang:2003:IOU

[300] Zhou Yang and Ya-Fen Chen. Induced ovulation using LHRHa in anadromous obscure puffer Takifugu obscurus cultured entirely in freshwater. Fish Physiology and Biochemistry, 29(4):323–326, December 2003. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-004-3326-3.

Anonymous:2003:VC

[301] Anonymous. Volume contents. Fish Physiology and Biochemistry, 29 (4):327–329, December 2003. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-004-7329-x.

Anonymous:2003:AI

[302] Anonymous. Author index. Fish Physiology and Biochemistry, 29 (4):330, December 2003. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-004-0330-6.

Bucking:2004:DUR

[303] Carol Bucking and Chris M. Wood. Does urea reabsorption occur via the glucose pathway in the kidney of the freshwater rainbow trout? Fish Physiology and Biochemistry, 30(1):1–12, March 2004. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-004-4154-1.

Bjorn:2004:MFC

[304] H. Björn, G. Ewald, and P. Frøyen. Metabolic fate of ^{14}C-labelled chlorinated and non-chlorinated fatty acids in goldfish (Carassius auratus). Fish Physiology and Biochemistry, 30(1):13–20, March 2004. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-004-4153-2.

Borges:2004:HSB

[305] Adriana Borges, Luciana V. Scotti, and Guillermo F. Wassermann. Hematologic and serum biochemical values for jundiá (Rhamdia quelen). Fish Physiology and Biochemistry, 30(1):21–25, March 2004. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-004-5000-1.

Swimmer:2004:MBC

[306] Y. Swimmer, L. McNaughton, and R. Brill. Metabolic biochemistry of cardiac muscle in three tuna species (bigeye, Thunnus obesus; yellowfin, T. albacares; and skipjack, Katsuwonus pelamis) with divergent ambient temperature and oxygen tolerances. Fish Physiology and Biochemistry, 30(1):27–35, March 2004. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-004-5465-y.

Strand:2004:ILD

[307] J. Strand, L. Andersen, and B. Korsgaard. Impaired larval development in broods of eelpout (Zoarces viviparus) in Danish coastal waters. Fish Physiology and Biochemistry, 30(1):37–46, March 2004. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-004-6003-7.

Imsland:2004:VGH

[308] A. K. Imsland, A. Foss, and T. M. Jonassen. Variations in growth in haemoglobin genotypes of Atlantic cod. Fish Physiology and Biochemistry, 30(1):47–55, March 2004. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-004-6787-5.

Sole:2004:VSD

[309] M. Solé, J. Potrykus, and J. Blasco. Variations on stress defences and metallothionein levels in the Senegal sole, Solea senegalensis, during early larval stages. Fish Physiology and Biochemistry, 30(1):57–66, March 2004. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-004-6786-6.

Jentoft:2004:MCE

[310] Sissel Jentoft, Are H. Aastveit, and Øivind Andersen. Molecular cloning and expression of insulin-like growth factor-i (IGF-I) in Eurasian perch (Perca fluviatilis): lack of responsiveness to growth hormone treatment. Fish Physiology and Biochemistry, 30(1):67–76, March 2004. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-005-0267-4.

Waser:2004:ODF

[311] W. P. Waser and N. Heisler. Oxygen delivery to the fish eye: blood flow in the pseudobranchial artery of rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 30(1):77–85, March 2004. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-005-0268-3.

Bayarri:2004:CMR

[312] M. J. Bayarri, R. García-Allegue, and F. J. Sánchez-Vázquez. Circadian melatonin release in vitro by European sea bass pineal. Fish Physiology and Biochemistry, 30(1):87–89, March 2004. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-004-6002-8.

Anonymous:2004:IA

[313] Anonymous. Instruction for authors. Fish Physiology and Biochemistry, 30(1):91–93, March 2004. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-005-4164-7.

Ran:2004:RGS

[314] Xueqin Ran, Wensheng Li, and Haoran Lin. Rat ghrelin stimulates GH release and GH mRNA expression in the pituitary of orange-spotted grouper, Epinephelus coioides. Fish Physiology and Biochemistry, 30 (2):95–102, June 2004. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-005-1768-x.

Toa:2004:SRJ

[315] D. G. Toa, L. O. B. Afonso, and G. K. Iwama. Stress response of juvenile rainbow trout (Oncorhynchus mykiss)to chemical cues released from stressed conspecifics. Fish Physiology and Biochemistry, 30(2):103–108, June 2004. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
005-0266-5.

Goldstein:2004:PCT

[316] J. P. Goldstein, A. Scott, and S. Currie. Are primary cultures of trout (Oncorhynchus mykiss) ventricular cardiomyocytes metabolically viable? Fish Physiology and Biochemistry, 30(2):109–117, June 2004. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-005-2492-2.

Zietara:2004:ACA

[317] M. S. Zietara, E. Slominska, and E. F. Skorkowski. ATP content and adenine nucleotide catabolism in African catfish spermatozoa stored in various energetic substrates. Fish Physiology and Biochemistry, 30(2):119–127, June 2004. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
005-2493-1.

Abi-Ayad:2004:DTB

[318] S.-M. E.-A. Abi-Ayad, Z. Boutiba, and P. Kestemont. Dynamics of total body fatty acids during early ontogeny of pikeperch (Sander lucioperca) larvae. Fish Physiology and Biochemistry, 30(2):129–136, June 2004. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-005-3417-9.

Dussault:2004:ECA

[319] È. B. Dussault, R. C. Playle, and R. S. McKinley. Effects of chronic aluminum exposure on swimming and cardiac performance in rainbow trout, Oncorhynchus mykiss. Fish Physiology and Biochemistry, 30(2):137–148, June 2004. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
005-4317-8.

Gjoen:2004:EDL

[320] T. Gjøen, A. Obach, and B. Ruyter. Effect of dietary lipids on macrophage function, stress susceptibility and disease esistance in Atlantic salmon (Salmo salar). Fish Physiology and Biochemistry, 30(2):149–161, June 2004. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
005-4318-7.

Skjaeraasen:2004:EPS

[321] J. E. Skjæraasen, A. G. V. Salvanes, and B. Norberg. The effect of photoperiod on sexual maturation, appetite and growth in wild Atlantic cod (Gadus morhua L.). Fish Physiology and Biochemistry, 30(2):163–174, June 2004. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
005-4319-6.

Ruchin:2004:ICL

[322] A. B. Ruchin. Influence of colored light on growth rate of juveniles of fish. Fish Physiology and Biochemistry, 30(2):175–178, June 2004. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-005-1263-4.

Tokushima:2004:ICB

[323] Y. Tokushima, Y. Ito, and A. Hara. Immunochemical comparison between sera from the primary and secondary circulations in a salmonid fish, Sakhalin taimen (Hucho perryi). Fish Physiology and Biochemistry, 30(2):179–188, June 2004. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-005-1264-3.

Mahmoud:2004:PHS

[324] A. B. Mahmoud and M. B. Al-Salahy. Physiological and histological studies on the effect of lupine seeds on the fish, Clarias lazera treated with glucose and alloxan. Fish Physiology and Biochemistry, 30(3-4):??, September 2004. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
005-8243-6.

Hamed:2004:GRE

[325] Ragaa R. Hamed, Tahany M. Maharem, and Rasha A. M. Guinidi. Glutathione and its related enzymes in the Nile fish. Fish Physiology and Biochemistry, 30(3-4):189–199, September 2004. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-005-3259-5.

Stekhoven:2004:EHT

[326] F. M. A. H. Schuurmans Stekhoven, S. E. Wendelaar Bonga, and G. Flik. Extranuclear histones in teleost gills: An evolutionary study. Fish Physiology and Biochemistry, 30(3-4):201–211, September 2004. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-005-7442-5.

Tiano:2004:ASI

[327] Luca Tiano, Ian Davies, and Giancarlo Falcioni. Analysis of stress-induced gene expression in trout red blood cells following tributyltinchloride exposure. Fish Physiology and Biochemistry, 30(3-4):231–240, September 2004.

CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-005-8244-5.

Bransden:2004:DAA

[328] M. P. Bransden, J. M. Cobcroft, and J. G. Bell. Dietary arachidonic acid alters tissue fatty acid profile, whole body eicosanoid production and resistance to hypersaline challenge in larvae of the temperate marine fish, striped trumpeter (Latris lineata). Fish Physiology and Biochemistry, 30(3-4):241–256, September 2004. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-005-8245-4.

Andersen:2004:EDA

[329] L. Andersen, K. Kinnberg, and P. Bjerregaard. Evaluation of a 40 day assay for testing endocrine disrupters: Effects of an anti-estrogen and an aromatase inhibitor on sex ratio and vitellogenin concentrations in juvenile zebrafish (Danio rerio). Fish Physiology and Biochemistry, 30(3-4):257–266, September 2004. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-005-8246-3.

Chiba:2004:IGR

[330] H. Chiba, M. Amano, and M. Iwata. Involvement of gonadotropinreleasing hormone in thyroxine release in three different forms of teleost fish: Barfin founder, masu salmon and goldfish. Fish Physiology and Biochemistry, 30(3-4):267–273, September 2004. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-005-8676-y.

Takle:2004:ISS

[331] H. Takle, G. Baeverfjord, and Ø. Andersen. Identification of stress sensitive genes in hyperthermic Atlantic salmon (Salmo salar) embryos by RAP-PCR. Fish Physiology and Biochemistry, 30(3-4):275–281, September 2004. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
005-0291-4.

Asturiano:2004:PCC

[332] J. F. Asturiano, L. Pérez, and M. Jover. Physio-chemical characteristics of seminal plasma and development of media and methods for the cryopreservation of European eel sperm. Fish Physiology and Biochemistry, 30(3-4): 283–293, September 2004. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-005-1553-x.

Anonymous:2004:VC

[333] Anonymous. Volume contents. Fish Physiology and Biochemistry, 30(3-4):295–297, September 2004. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-005-4325-8.

Anonymous:2004:AI

[334] Anonymous. Author index. Fish Physiology and Biochemistry, 30(3-4):299, September 2004. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-005-4326-7.

Dietrich:2005:MPR

[335] G. J. Dietrich, R. Kowalski, and A. Ciereszko. Motility parameters of rainbow trout (Oncorhynchus mykiss) spermatozoa in relation to sequential collection of milt, time of Post-mortem storage and anesthesia. Fish Physiology and Biochemistry, 31(1):1–9, January 2005. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-005-3527-4.

Rothwell:2005:AEH

[336] S. E. Rothwell and M. E. Forster. Anaesthetic effects on the hepatic portal vein and on the vascular resistance of the tail of the Chinook salmon (Oncorhynchus tshawytscha). Fish Physiology and Biochemistry, 31(1):11–21, January 2005. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
005-3847-4.

Peres:2005:PEM

[337] H. Peres and A. Oliva-Teles. Protein and energy metabolism of European seabass (Dicentrarchus labrax) juveniles and estimation of maintenance requirements. Fish Physiology and Biochemistry, 31(1):23–31, January 2005. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-005-4586-2.

Higgins:2005:SIA

[338] Angela D. Higgins, Jeffrey T. Silverstein, and Kenneth P. Blemings. Starvation induced alterations in hepatic Lysine metabolism in different families of rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 31(1):33–44, January 2005. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-005-4587-1.

Borges:2005:BCS

[339] Adriana Borges, Daniele Rodrigues Siqueira, and Guillermo Federico Wassermann. Biochemical composition of seminal plasma and annual variations in semen characteristics of Jundiá Rhamdia quelen (Quoy and Gaimard, Pimelodidae). Fish Physiology and Biochemistry, 31(1):45–53, January 2005. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
005-4742-8.

Metcalf:2005:FAT

[340] V. J. Metcalf and N. J. Gemmell. Fatty acid transport in cartilaginous fish: Absence of albumin and possible utilization of lipoproteins. Fish Physiology and Biochemistry, 31(1):55–64, January 2005. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-005-5124-y.

Valenzuela:2005:EAH

[341] A. Valenzuela, V. Silva, and A. Klempau. Effect of acute hypoxia in trout (Oncorhynchus mykiss) on immature erythrocyte release and production of oxidative radicals. Fish Physiology and Biochemistry, 31(1):65–72, January 2005. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
005-5288-5.

Waser:2005:SAB

[342] W. P. Waser, A. Schmitz, and A. Wobschall. Stereological analysis of blood space and tissue types in the pseudobranch of the rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 31(1):73–82, January 2005. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
006-0002-9.

Alvarez-Gonzalez:2005:DDE

[343] C. A. Alvarez-González, M. Cervantes-Trujano, and R. Piedrahita. Development of digestive enzymes in California halibut Paralichthys californicus larvae. Fish Physiology and Biochemistry, 31(1):83–93, January 2005. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-006-0003-8.

Matsumoto:2005:DBT

[344] T. Matsumoto, Y. Nagashima, and K. Shiomi. Difference between tetrodotoxin and saxitoxins in accumulation in puffer fish Takifugu rubripes liver tissue slices. Fish Physiology and Biochemistry, 31(1):

95–100, January 2005. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-006-0001-x.

Lal:2005:IPV

[345] B. Lal and A. K. Singh. Immunological and physiological validation of enzyme-linked immunosorbent assay for growth hormone of the Asian catfish, Clarias batrachus. Fish Physiology and Biochemistry, 31(2-3):??, April 2005. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
006-0038-x.

Rajesh:2005:TIA

[346] R. Rajesh and K. C. Majumdar. Transgene integration — an analysis in autotransgenic Labeo rohita Hamilton (Pisces: Cyprinidae). Fish Physiology and Biochemistry, 31(2-3):??, April 2005. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-006-0037-y.

Rasheeda:2005:TIA

[347] M. K. Rasheeda, G. Sreenivasulu, and B. Senthilkumaran. Thioureainduced alteration in the expression patterns of some steroidogenic enzymes in the air-breathing catfish Clarias gariepinus. Fish Physiology and Biochemistry, 31(2-3):??, April 2005. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-006-0036-z.

Jacob:2005:TIA

[348] T. N. Jacob, J. P. Pandey, and B. Senthilkumaran. Thyroxine-induced alterations in the testis and seminal vesicles of air-breathing catfish, Clarias gariepinus. Fish Physiology and Biochemistry, 31(2-3):??, April 2005. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-006-0035-0.

Supriya:2005:THM

[349] A. Supriya, K. Raghuveer, and B. Senthilkumaran. Thyroid hormone modulation of ovarian recrudescence of air-breathing catfish Clarias gariepinus. Fish Physiology and Biochemistry, 31(2-3):??, April 2005. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-006-0034-1.

Wang:2005:MCG

[350] D. S. Wang, B. Senthilkumaran, and Y. Nagahama. Molecular cloning, gene expression and characterization of the third estrogen receptor of the Nile tilapia, Oreochromis niloticus. Fish Physiology and Biochemistry, 31(2-3):??, April 2005. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-006-0033-2.

Bogerd:2005:SLB

[351] Jan Bogerd. Selective ligand-binding determinants in catfish and human gonadotropin receptors. Fish Physiology and Biochemistry, 31(2-3):??, April 2005. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
006-0032-3.

Phartyal:2005:VIV

[352] Rajendra Phartyal, L. Bibekananda Singh, and Neeta Sehgal. In vitro induction of vitellogenin by estradiol 17 β in isolated hepatocytes of catfish, Clarias gariepinus. Fish Physiology and Biochemistry, 31(2-3):??, April 2005. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
006-0031-4.

Manohar:2005:PVA

[353] D. Manohar, G. Damodar Rao, and Aparna Dutta Gupta. Purification of vitellogenin from the air breathing catfish, Clarias gariepinus. Fish Physiology and Biochemistry, 31(2-3):??, April 2005. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-006-0030-5.

Sarang:2005:EPG

[354] M. K. Sarang and B. Lal. Effect of piscine GH/IGF-I on final oocyte maturation in vitro in Heteropneustes fossilis. Fish Physiology and Biochemistry, 31(2-3):??, April 2005. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-006-0029-y.

Sreenivasulu:2005:EHD

[355] G. Sreenivasulu, I. Swapna, and B. Senthilkumaran. Expression of 20βhydroxysteroid dehydrogenase and P450 17α-hydroxylase/c17–20 lyase during hCG-induced in vitro oocyte maturation in snake head murrel Channa striatus. Fish Physiology and Biochemistry, 31(2-3):??, April 2005. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
006-0028-z.

Piferrer:2005:ADR

[356] Francesc Piferrer and Mercedes Blázquez. Aromatase distribution and regulation in fish. Fish Physiology and Biochemistry, 31(2-3):??, April 2005. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-006-0027-0.

Maitra:2005:IMO

[357] Saumen Kumar Maitra, Asamanja Chattoraj, and Sharmistha Bhattacharyya. Implication of melatonin in oocyte maturation in Indian major carp Catla catla. Fish Physiology and Biochemistry, 31(2-3):??, April 2005. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-006-0025-2.

Kagawa:2005:FSG

[358] Hirohiko Kagawa, Hideki Tanaka, and Kazuharu Nomura. The first success of glass eel production in the world: basic biology on fish reproduction advances new applied technology in aquaculture. Fish Physiology and Biochemistry, 31(2-3):??, April 2005. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-006-0024-3.

Kumar:2005:ETH

[359] K. V. Sathya Sai Kumar, I. Swapna, and B. Senthilkumaran. Effect of thyroid hormone depletion on monoamines and expression patterns of catfish GnRH in the air-breathing catfish, Clarias gariepinus. Fish Physiology and Biochemistry, 31(2-3):??, April 2005. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-006-0023-4.

Kirubagaran:2005:SDG

[360] R. Kirubagaran, B. Senthilkumaran, and K. P. Joy. Seasonal dynamics in gonadotropin secretion and E_{2}-binding in the catfish Heteropneustes fossilis. Fish Physiology and Biochemistry, 31(2-3):??, April 2005. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-006-0022-5.

Sakai:2005:ILG

[361] F. Sakai, I. Swapna, and B. Senthilkumaran. Immunocytochemical localization of gonadotropins during the development of XX and XY Nile tilapia. Fish Physiology and Biochemistry, 31(2-3):??, April 2005. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-006-0021-6.

Mazumdar:2005:GLP

[362] Minakshi Mazumdar, Praful Singru, and Nishikant Subhedar. Glucagonlike peptide-1 (GLP-1) immunoreactivity in the forebrain and pituitary of the teleost, Clarias batrachus (Linn.). Fish Physiology and Biochemistry, 31(2-3):??, April 2005. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-006-0040-3.

Subhedar:2005:RNN

[363] Nishikant Subhedar, Archana Gaikwad, and Subhash Saha. Role of neuropeptide Y (NPY) in the regulation of reproduction: study based on catfish model. Fish Physiology and Biochemistry, 31(2-3):??, April 2005. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-006-0020-7.

Sakharkar:2005:RPR

[364] Amul Sakharkar, Praful Singru, and Nishikant Subhedar. Reproduction phase-related variations in the GnRH immunoreactive fibers in the pineal of the Indian major carp Cirrhinus mrigala (Ham.). Fish Physiology and Biochemistry, 31(2-3):??, April 2005. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-006-0039-9.

Swapna:2005:SGP

[365] I. Swapna, G. Sreenivasulu, and B. Senthilkumaran. Seabream GnRH: partial cDNA cloning, localization and stage-dependent expression in the ovary of snake head murrel, Channa striatus. Fish Physiology and Biochemistry, 31(2-3):??, April 2005. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-006-0019-0.

Parhar:2005:GGL

[366] Ishwar S. Parhar. GnRH and gpcr: laser-captured single cell gene profiling. Fish Physiology and Biochemistry, 31(2-3):??, April 2005. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-006-0018-1.

Nocillado:2005:IDD

[367] J. N. Nocillado, B. Levavi-Sivan, and A. Elizur. Isolation of dopamine D_{2} receptor (D_{2}R) promoters in Mugil cephalus. Fish Physiology and Biochemistry, 31(2-3):??, April 2005. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-006-0017-2.

Raghuveer:2005:EMT

[368] K. Raghuveer, R. Garhwal, and B. Senthilkumaran. Effect of methyl testosterone- and ethynyl estradiol-induced sex differentiation on catfish, Clarias gariepinus: expression profiles of DMRT1, cytochrome P450aromatases and 3 β-hydroxysteroid dehydrogenase. Fish Physiology and Biochemistry, 31(2-3):??, April 2005. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-006-0016-3.

Sundaray:2005:ICS

[369] J. K. Sundaray, K. Ohta, and M. Matsuyama. Isolation, cloning, sequencing of brain type aromatase and its expression in male and female wrasse, Pseudolabrus sieboldi. Fish Physiology and Biochemistry, 31(2-3):??, April 2005. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
006-0015-4.

Sudhakumari:2005:OEP

[370] C. C. Sudhakumari, B. Senthilkumaran, and Y. Nagahama. Ontogenic expression patterns of several nuclear receptors and cytochrome P450 aromatases in brain and gonads of the Nile tilapia Oreochromis niloticus suggests their involvement in sex differentiation. Fish Physiology and Biochemistry, 31(2-3):??, April 2005. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-006-0014-5.

Nakamura:2005:SCC

[371] Masaru Nakamura, Yasuhisa Kobayashi, and Ramji Kumar Bhandari. Sex change in coral reef fish. Fish Physiology and Biochemistry, 31(2-3):??, April 2005. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
006-7595-x.

Nagahama:2005:MMS

[372] Y. Nagahama. Molecular mechanisms of sex determination and gonadal sex differentiation in fish. Fish Physiology and Biochemistry, 31(2-3):??, April 2005. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
006-7590-2.

Anonymous:2005:F

[373] Anonymous. Foreword. Fish Physiology and Biochemistry, 31(2-3):??, April 2005. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
006-7589-8.

Twan:2005:CAE

[374] Wen-Hung Twan, Hua-Fang Wu, and Ching-Fong Chang. Corals have already evolved the vertebrate-type hormone system in the sexual reproduction. Fish Physiology and Biochemistry, 31(2-3):111–115, April 2005. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-006-7591-1.

Kobayashi:2005:PAT

[375] Yasuhisa Kobayashi, Tomoki Sunobe, and Masaru Nakamura. Promoter analysis of two aromatase genes in the serial-sex changing gobiid fish, Trimma okinawae. Fish Physiology and Biochemistry, 31(2-3):123–127, April 2005. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
006-0013-6.

Ge:2005:GTP

[376] Wei Ge. Gonadotropins and their paracrine signaling network in the zebrafish ovary. Fish Physiology and Biochemistry, 31(2-3):209–214, April 2005. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
006-0026-1.

Zhang:2005:NSM

[377] W. M. Zhang, Y. Zhang, and G. Z. Ma. Nucleotide sequence and mRNA expression analysis of β-actin gene in the orange-spotted grouper Epinephelus coioides. Fish Physiology and Biochemistry, 31(4):??, December 2005. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
005-4743-7.

Gerwick:2005:III

[378] Lena Gerwick, Graham E. Corley-Smith, and Christopher J. Bayne. Intracranial injections induce local transcription of a gene encoding precerebellin-like protein. Fish Physiology and Biochemistry, 31(4):??, December 2005. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
005-4588-0.

Imsland:2005:IBT

[379] A. K. Imsland, A. Foss, and T. M. Jonassen. The interrelation between temperature regimes and fish size in juvenile Atlantic cod (Gadus morhua): effects on growth and feed conversion efficiency. Fish Physiology and Biochemistry, 31(4):??, December 2005. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-005-4244-8.

Eddy:2005:CFJ

[380] F. B. Eddy. Cardiac function in juvenile salmon (Salmo salar L.) in response to lipopolysaccharide (LPS) and inhibitors of inducible nitric oxide synthase (iNOS). Fish Physiology and Biochemistry, 31(4):??, December 2005. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
005-3528-3.

Benedet:2005:CTA

[381] Susana Benedet, Viktoria Johansson, and Björn Thrandur Björnsson. Cloning of two Atlantic salmon growth hormone receptor isoforms and in vitro ligand-binding response. Fish Physiology and Biochemistry, 31 (4):??, December 2005. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-005-2524-y.

Beyea:2005:HSP

[382] M. M. Beyea, T. J. Benfey, and J. D. Kieffer. Hematology and stress physiology of juvenile diploid and triploid shortnose sturgeon (Acipenser brevirostrum). Fish Physiology and Biochemistry, 31(4):??, December 2005. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-005-1552-y.

Lermen:2005:AAA

[383] C. L. Lermen, V. L. P. Vieira, and V. M. Morsch. ATP, ADP and AMP dephosphorylation in membrane fractions of Rhamdia quelen exposed to different temperatures. Fish Physiology and Biochemistry, 31(4):??, December 2005. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
005-1392-9.

Huang:2005:PDG

[384] J. W. Huang, L. X. Tian, and Y. J. Liu. Pyridoxine deficiency of grouper, Epinephelus coioides: physiological and biochemical alteration. Fish Physiology and Biochemistry, 31(4):331–337, December 2005. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-005-2832-2.

Jebali:2006:EMC

[385] J. Jebali, M. Banni, and H. Boussetta. Effects of malathion and cadmium on acetylcholinesterase activity and metallothionein levels in the fish Seriola dumerilli. Fish Physiology and Biochemistry, 32(1):??, March 2006. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-006-0041-2.

Yeh:2006:MCS

[386] Hung-Yueh Yeh, Craig A. Shoemaker, and Phillip H. Klesius. Molecular cloning and sequencing of hemoglobin-β gene of channel catfish, Ictalurus punctatus Rafinesque. Fish Physiology and Biochemistry, 32(1):??, March 2006. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
006-7658-z.

Biswas:2006:AEA

[387] Angshuman Biswas, Mitali Pramanik, and Arun Kumar Ray. Alterations in enzyme activities in vital organs of triploid female catfish Heteropneustes fossilis (Bloch). Fish Physiology and Biochemistry, 32(1):??, March 2006. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
006-7657-0.

Neer:2006:SOC

[388] J. A. Neer, J. K. Carlson, and B. A. Thompson. Standard oxygen consumption of seasonally acclimatized cownose rays, Rhinoptera bonasus (Mitchill 1815), in the northern Gulf of Mexico. Fish Physiology and Biochemistry, 32(1):??, March 2006. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-006-7312-9.

Shi:2006:CBB

[389] Xiaotao Shi, Dapeng Li, and Liangqi Long. Comparative blood biochemistry of Amur sturgeon, Acipenser schrenckii, and Chinese surgeon, Acipenser sinensis. Fish Physiology and Biochemistry, 32(1):??, March 2006. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
006-7134-9.

Yanbo:2006:ATN

[390] Wang Yanbo, Zhang Wenju, and Xu Zirong. Acute toxicity of nitrite on tilapia (Oreochromis niloticus) at different external chloride concentrations. Fish Physiology and Biochemistry, 32(1):??, March 2006. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-005-5744-2.

Asadi:2006:SBP

[391] F. Asadi, M. Masoudifard, and P. Khazraeinia. Serum biochemical parameters of Acipenser persicus. Fish Physiology and Biochemistry, 32(1):??, March 2006. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
005-5738-0.

Shi:2006:PCH

[392] Zhen-Ping Shi, Ting-Jun Fan, and Ling-Ling Yang. Purification and characterization of hatching enzyme from flounder Paralichthys olivaceus. Fish Physiology and Biochemistry, 32(1):??, March 2006. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-005-5250-6.

Li:2006:EAN

[393] Guang-Li Li, Xiao-Chun Liu, and Hao-Ran Lin. Effects of aromatizable and nonaromatizable androgens on the sex inversion of redspotted grouper (Epinephelus akaara). Fish Physiology and Biochemistry, 32(1):??, March 2006. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-005-4900-z.

Rungruangsak-Torrissen:2006:DET

[394] K. Rungruangsak-Torrissen, R. Moss, and R. Waagbø. Different expressions of trypsin and chymotrypsin in relation to growth in Atlantic salmon (Salmo salar L.). Fish Physiology and Biochemistry, 32(1):??, March 2006. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-005-0630-5.

Salem:2006:ING

[395] M. Salem, C. E. Rexroad, and J. Yao. Identification of a novel gillspecific calpain from rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 32(1):??, March 2006. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-005-0560-2.

Zhang:2006:DDT

[396] Genhua Zhang, Shaoping Deng, and Leilei Li. Distribution of different taste buds and expression of α-gustducin in the barbells of yellow catfish (Pelteobagrus fulvidraco). Fish Physiology and Biochemistry, 32(1):55–62, March 2006. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
006-6937-z.

Popovic:2006:CFT

[397] N. Topic Popovic, I. Strunjak-Perovic, and M. Jadan. Comparison of five tuna plasma analytes measured on two automated blood analyzers. Fish Physiology and Biochemistry, 32(2):99–103, June 2006. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/PL00021847.

Becker:2006:ILG

[398] A. G. Becker, J. F. Gonçalves, and B. Baldisserotto. Ion levels in the gastrointestinal tract content and plasma of four teleosts with different feeding habits. Fish Physiology and Biochemistry, 32(2):105–112, June 2006. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-006-9102-9.

Valenzuela:2006:ECL

[399] Ariel E. Valenzuela, Victor M. Silva, and Alfredo E. Klempau. Effects of constant light on haematological parameters of cultured rainbow trout (Oncorhynchus mykiss) in the Southern Hemisphere. Fish Physiology and Biochemistry, 32(2):113–120, June 2006. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-006-9103-8.

Yao:2006:ICG

[400] Kouakou Yao, Jean-Charles Gabillard, and Pierre-Yves Le Bail. Influence of circulating GH levels on GH-binding capacity measurements in the hepatic membrane of rainbow trout (Oncorhynchus mykiss): importance of normalization of results. Fish Physiology and Biochemistry, 32 (2):121–130, June 2006. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-006-0004-7.

Lahnsteiner:2006:CMV

[401] F. Lahnsteiner. Carbohydrate metabolism of vitellogenic follicles and eggs of Serranus cabrilla (Serranidae) and Mullus barbatus (Mullidae) and of embryos of Sparus aurata (Sparidae). Fish Physiology and Biochemistry, 32(2):131–139, June 2006. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-006-0005-6.

Fudan:2006:ERT

[402] Tong Fudan, Wang Qun, and Zhang Haihua. Expression of recombinant triangular bream (Megalobrama terminalis) pre-IGF-i in Escherichia coli. Fish Physiology and Biochemistry, 32(2):141–145, June 2006. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-006-0006-5.

Comabella:2006:DEA

[403] Y. Comabella, R. Mendoza, and T. García-Galano. Digestive enzyme activity during early larval development of the Cuban gar Atractosteus tristoechus. Fish Physiology and Biochemistry, 32(2):147–157, June 2006. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-006-0007-4.

Favaloro:2006:MCC

[404] Eugenia Favaloro and Antonio Mazzola. Meristic character counts and incidence of skeletal anomalies in the wild Diplodus puntazzo (Cetti, 1777) of an area of the south-eastern Mediterranean Sea. Fish Physiology and Biochemistry, 32(2):159–166, June 2006. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-006-0008-3.

Rollo:2006:LMF

[405] A. Rollo, R. Sulpizio, and O. Carnevali. Live microbial feed supplement in aquaculture for improvement of stress tolerance. Fish Physiology and Biochemistry, 32(2):167–177, June 2006. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-006-0009-2.

Wang:2006:PCS

[406] Hai ying Wang, Yue jun Wang, and Mi Sun. Purification and characterization of stomach protease from the turbot (Scophthalmus maximus L.). Fish Physiology and Biochemistry, 32(2):179–188, June 2006. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-006-0010-9.

Moriyama:2006:ILG

[407] Shunsuke Moriyama, Kiyoko Yamaguchi, and Hiroshi Kawauchi. Insulinlike growth factor I of Japanese eel, Anguilla japonica: cDNA cloning, tissue distribution, and expression after treatment with growth hormone and seawater acclimation. Fish Physiology and Biochemistry, 32(3):189–201, October 2006. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
006-0011-8.

Behl:2006:SCA

[408] Rahul Behl. Stimulation of catalase activity in carp (Cyprinus carpio) ovarian follicles by gonadotropin in vitro. Fish Physiology and Biochemistry, 32(3):203–207, October 2006. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-006-0012-7.

Kumar:2006:SDD

[409] S. Kumar, N. P. Sahu, and S. C. Mukherjee. Studies on digestibility and digestive enzyme activities in Labeo rohita (Hamilton) juveniles: effect of microbial α-amylase supplementation in non-gelatinized or gelatinized corn-based diet at two protein levels. Fish Physiology and Biochemistry, 32(3):209–220, October 2006. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-006-9002-z.

Vutukuru:2006:AEC

[410] Sesha Srinivas Vutukuru, Suma Chintada, and Y. Anjaneyulu. Acute effects of copper on superoxide dismutase, catalase and lipid peroxidation in the freshwater teleost fish, Esomus danricus. Fish Physiology and Biochemistry, 32(3):221–229, October 2006. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-006-9004-x.

Fernandino:2006:EDE

[411] J. I. Fernandino, L. G. Guilgur, and G. M. Somoza. Dmrt1 expression analysis during spermatogenesis in pejerrey, Odontesthes bonariensis. Fish Physiology and Biochemistry, 32(3):231–240, October 2006. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-006-9005-9.

Pham:2006:CBP

[412] Ky Xuan Pham, Masafumi Amano, and Kunio Yamamori. Changes in brain and pituitary GnRH levels during ovarian maturation in wild female Japanese flounder. Fish Physiology and Biochemistry, 32(3):241–248, October 2006. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
006-9006-8.

Coulibaly:2006:EEU

[413] I. Coulibaly, S. A. Gahr, and C. E. Rexroad III. Embryonic expression of UCP2 in rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 32(3):249–253, October 2006. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-006-9101-x.

Celik:2006:EET

[414] Ekrem Sanver Celik and Seyit Aydin. Effect of Trachelobdella lubrica (Hirudinea: Piscicolidae) on biochemical and haematological characteristics of black scorpion fish (Scorpaena porcus, Linnaeus 1758). Fish Physiology and Biochemistry, 32(3):255–260, October 2006. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-006-9003-y.

Werner:2006:CRT

[415] I. Werner, M. R. Viant, and M. L. Johnson. Cellular responses to temperature stress in steelhead trout (Onchorynchus mykiss) parr with different rearing histories. Fish Physiology and Biochemistry, 32(3):261–273, October 2006. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
006-9105-6.

Gaye-Siessegger:2006:EDP

[416] J. Gaye-Siessegger, U. Focken, and K. Becker. Effect of dietary protein/ carbohydrate ratio on activities of hepatic enzymes involved in the amino acid metabolism of Nile tilapia, Oreochromis niloticus (L.). Fish Physiology and Biochemistry, 32(4):275–282, December 2006. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-006-9000-1.

Brill:2006:EOC

[417] R. W. Brill and P. G. Bushnell. Effects of open- and closed-system temperature changes on blood O_{2}-binding characteristics of Atlantic bluefin tuna (Thunnus thynnus). Fish Physiology and Biochemistry, 32(4): 283–294, December 2006. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-006-9104-7.

Jun-sheng:2006:OPA

[418] Li Jun-sheng, Li Jian-lin, and Wu Ting-ting. Ontogeny of protease, amylase and lipase in the alimentary tract of hybrid juvenile tilapia (Oreochromis niloticus × Oreochromis aureus). Fish Physiology and Biochemistry, 32(4):295–303, December 2006. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-006-9106-5.

Nam:2006:CCZ

[419] Yoon Kwon Nam, Young Sun Cho, and Dong Soo Kim. Characterization of copper, zinc superoxide dismutase from a cartilaginous shark species, Scyliorhinus torazame (Carcharhiniformes). Fish Physiology and Biochemistry, 32(4):305–315, December 2006. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-006-9107-4.

Fonseca-Madrigal:2006:NER

[420] Jorge Fonseca-Madrigal, J. Gordon Bell, and Douglas R. Tocher. Nutritional and environmental regulation of the synthesis of highly unsaturated fatty acids and of fatty-acid oxidation in Atlantic salmon (Salmo salar L.) enterocytes and hepatocytes. Fish Physiology and Biochemistry, 32 (4):317–328, December 2006. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-006-9109-2.

Park:2006:ALA

[421] Kwan Ha Park, Bendik Fyhn Terjesen, and Konrad Dabrowski. α-lipoic acid-enrichment partially reverses tissue ascorbic acid depletion in pacu (Piaractus mesopotamicus) fed vitamin C-devoid diets. Fish Physiology and Biochemistry, 32(4):329–338, December 2006. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-006-9110-9.

Strumbo:2006:CAT

[422] B. Strumbo, L. Sangiorgio, and T. Simonic. Cloning and analysis of transcripts and genes encoding fish-specific proteins related to PrP. Fish Physiology and Biochemistry, 32(4):339–353, December 2006. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-006-9108-3.

Osman:2007:ELN

[423] Alaa G. M. Osman, Imam A. Mekkawy, and Frank Kirschbaum. Effects of lead nitrate on the activity of metabolic enzymes during early developmental stages of the African catfish, Clarias gariepinus (Burchell, 1822). Fish Physiology and Biochemistry, 33(1):1–13, April 2007. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-006-9111-8.

John:2007:ACB

[424] P. Joseph John. Alteration of certain blood parameters of freshwater teleost Mystus vittatus after chronic exposure to Metasystox and Sevin. Fish Physiology and Biochemistry, 33(1):15–20, April 2007. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-006-9112-7.

Gacic:2007:SSD

[425] Z. Gačić, I. Damjanović, and M. Nikčević. Spectral sensitivity of the dogfish shark (Scyliorhinus canicula). Fish Physiology and Biochemistry, 33(1):21–27, April 2007. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-006-9113-6.

Liu:2007:SSB

[426] Hongxia Liu, Meiqin Yi, and Xiwu Gao. Substrate specificity of brain acetylcholinesterase and its sensitivity to carbamate insecticides in Carassius auratus. Fish Physiology and Biochemistry, 33(1):29–34, April 2007. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-006-9114-5.

Valenzuela:2007:SCH

[427] A. E. Valenzuela, V. M. Silva, and A. E. Klempau. Some changes in the haematological parameters of rainbow trout (Oncorhynchus mykiss) exposed to three artificial photoperiod regimes. Fish Physiology and Biochemistry, 33(1):35–48, April 2007. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-006-9115-4.

Cho:2007:EDI

[428] Sung Hwoan Cho, Sang-Mok Lee, and Sung-Yong Oh. Effect of dietary inclusion of various sources of green tea on growth, body composition and blood chemistry of the juvenile olive flounder, Paralichthys olivaceus. Fish Physiology and Biochemistry, 33(1):49–57, April 2007. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-006-9116-3.

Turner:2007:DRS

[429] M. A. Turner, M. R. Viant, and M. L. Johnson. Developmental rates, structural asymmetry, and metabolic fingerprints of steelhead trout (Oncorhynchus mykiss) eggs incubated at two temperatures. Fish Physiology and Biochemistry, 33(1):59–72, April 2007. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-006-9117-2.

Chen:2007:SEZ

[430] Yau-Hung Chen, Wei-Li Chen, and Ming-Kai Chern. Spatiotemporal expression of zebrafish D-amino acid oxidase during early embryogenesis. Fish Physiology and Biochemistry, 33(1):73–80, April 2007. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-006-9118-1.

Palacios:2007:LCF

[431] Elena Palacios, Ilie S. Racotta, and Carlos A. Martínez-Palacios. Lipid classes and fatty acids during embryogenesis of captive and wild silverside (Chirostoma estor estor) from Pátzcuaro Lake. Fish Physiology and Biochemistry, 33(1):81–91, April 2007. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-006-9119-0.

Yuan:2007:AHI

[432] Chuntao Yuan, Dongmei Li, and Xiaodong Han. Administration of a herbal immunoregulation mixture enhances some immune parameters in carp (Cyprinus carpio). Fish Physiology and Biochemistry, 33(2):93–101, June 2007. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
006-9120-7.

Leknes:2007:LLP

[433] I. L. Leknes. Lipofuscin-like pigments in endocardial cells enveloping the heart muscle trabeculae in platyfish (Poeciliidae: Teleostei). Fish Physiology and Biochemistry, 33(2):103–107, June 2007. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-006-9121-6.

Tashjian:2007:IDS

[434] D. Tashjian, J. J. Cech, Jr., and S. S. O. Hung. Influence of dietary l-selenomethionine exposure on the survival and osmoregulatory capacity of white sturgeon in fresh and brackish water. Fish Physiology and Biochemistry, 33(2):109–119, June 2007. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-006-9122-5.

Hamza:2007:EWA

[435] Neila Hamza, Mohamed Mhetli, and Patrick Kestemont. Effects of weaning age and diets on ontogeny of digestive activities and structures of pikeperch (Sander lucioperca) larvae. Fish Physiology and Biochemistry, 33(2):121–133, June 2007. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-006-9123-4.

Fu:2007:SDA

[436] Shi-Jian Fu. The specific dynamic action of southern catfish, Silurus meridonalis Chen, fed diets containing either raw or precooked corn starch or glucose. Fish Physiology and Biochemistry, 33(2):135–141, June 2007. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-006-9124-3.

Wang:2007:CCN

[437] Jiaqing Wang, Lin Hou, and Jialu An. Cloning and characterization of a novel albinism-related zinc finger protein gene in Japanese flounder. Fish Physiology and Biochemistry, 33(2):143–151, June 2007. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-006-9125-2.

Unal:2007:GHS

[438] G. Ünal, V. Türkoğlu, and B. Kaptaner. Gonadal histology and some biochemical characteristics of Chalcalburnus tarichi (Pallas, 1811) having abnormal gonads. Fish Physiology and Biochemistry, 33(2):153–165, June 2007. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
006-9126-1.

Huang:2007:EDT

[439] J. W. Huang, L. X. Tian, and Y. J. Liu. Effects of dietary thiamin on the physiological status of the grouper Epinephelus coioides. Fish Physiology and Biochemistry, 33(2):167–172, June 2007. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-007-9127-8.

Salles:2007:ICP

[440] C. M. C. Salles, P. Gagliano, and S. Giovanni De-Simone. Identification and characterization of proteases from skin mucus of tambacu, a neotropical hybrid fish. Fish Physiology and Biochemistry, 33(2):173–179, June 2007. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
007-9128-7.

Qi:2007:PCC

[441] H. Qi, X. P. Dong, and B. W. Zhu. Purification and characterization of a cysteine-like protease from the body wall of the sea cucumber Stichopus japonicus. Fish Physiology and Biochemistry, 33(2):181–188, June 2007. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-007-9129-6.

Sink:2007:DWB

[442] Todd David Sink, Sathyanand Kumaran, and Rebecca T. Lochmann. Development of a whole-body cortisol extraction procedure for determination of stress in golden shiners, Notemigonus crysoleucas. Fish Physiology and Biochemistry, 33(3):189–193, September 2007. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-007-9130-0.

Shen:2007:CFL

[443] Xue-Yan Shen, Jian-Zhou Cui, and Yoshitaka Nagahama. Cloning of the full-length coding sequence and expression analysis of Sox9b in guppy (Poecilia reticulata). Fish Physiology and Biochemistry, 33(3): 195–202, September 2007. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-007-9131-z.

Ahmed:2007:ERS

[444] Imtiaz Ahmed. Effect of ration size on growth, body composition, and energy and protein maintenance requirement of fingerling Indian major carp, Labeo rohita (Hamilton). Fish Physiology and Biochemistry, 33(3): 203–212, September 2007. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-007-9132-y.

Zexia:2007:MSP

[445] Gao Zexia, Wang Weimin, and James S. Diana. Morphological studies of peripheral blood cells of the Chinese sturgeon, Acipenser sinensis. Fish Physiology and Biochemistry, 33(3):213–222, September 2007. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-007-9133-x.

delaParra:2007:PCD

[446] Ana Matus de la Parra, Antonio Rosas, and Maria Teresa Viana. Partial characterization of the digestive enzymes of Pacific bluefin tuna Thunnus orientalis under culture conditions. Fish Physiology and Biochemistry, 33 (3):223–231, September 2007. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-007-9134-9.

Yeh:2007:CCC

[447] Hung-Yueh Yeh and Phillip H. Klesius. cDNA cloning, characterization, and expression analysis of channel catfish (Ictalurus punctatus Rafinesque, 1818) peroxiredoxin 6 gene. Fish Physiology and Biochemistry, 33(3): 233–239, September 2007. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-007-9135-8.

Hasan:2007:ECL

[448] Mahmud Hasan and Amrit N. Bart. Effects of capture, loading density and transport stress on the mortality, physiological responses, bacterial density and growth of rohu Labeo rohita fingerlings. Fish Physiology and Biochemistry, 33(3):241–248, September 2007. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-007-9136-7.

Kamangar:2007:CBC

[449] Barzan Bahrami Kamangar, Mohammad Javad Rasaee, and Mahmoud Bahmani. Correlations between circulating insulin-like growth factor-i and thyroxine and cortisol hormone levels, and some biometrical traits in female brood stocks during the late stages of sex maturation and in juvenile Persian sturgeon (Acipenser persicus). Fish Physiology and Biochemistry, 33(3):249–257, September 2007. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-007-9137-6.

Shrimpton:2007:BMD

[450] J. Mark Shrimpton, Aurora M. C. Sentlinger, and Daniel D. Heath. Biochemical and molecular differences in diploid and triploid ocean-type Chinook salmon (Oncorhynchus tshawytscha) smolts. Fish Physiology and Biochemistry, 33(3):259–268, September 2007. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-007-9138-5.

Mourente:2007:DDT

[451] Gabriel Mourente, J. Gordon Bell, and Douglas R. Tocher. Does dietary tocopherol level affect fatty acid metabolism in fish? Fish Physiology and Biochemistry, 33(3):269–280, September 2007. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-007-9139-4.

Joy:2007:F

[452] K. P. Joy. Foreword. Fish Physiology and Biochemistry, 33(4):281, December 2007. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
007-9161-6.

Saha:2007:FUA

[453] N. Saha and B. K. Ratha. Functional ureogenesis and adaptation to ammonia metabolism in Indian freshwater air-breathing catfishes. Fish Physiology and Biochemistry, 33(4):283–295, December 2007. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-007-9172-3.

Singru:2007:NNO

[454] Praful S. Singru, Amul J. Sakharkar, and Nishikant Subhedar. Neuronal nitric oxide synthase immunoreactivity in forebrain, pineal, and pituitary of Oreochromis mossambicus (Tilapia). Fish Physiology and Biochemistry, 33(4):297–309, December 2007. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-007-9169-y.

Rajesh:2007:CAS

[455] R. Rajesh and K. C. Majumdar. A comparative account of the structure of the growth hormone encoding gene and genetic interrelationship in six species of the genus Labeo. Fish Physiology and Biochemistry, 33 (4):311–333, December 2007. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-007-9164-3.

Swapna:2007:THM

[456] I. Swapna and B. Senthilkumaran. Thyroid hormones modulate the hypothalamo–hypophyseal–gonadal axis in teleosts: Molecular insights. Fish Physiology and Biochemistry, 33(4):335–345, December 2007. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-007-9165-2.

Garg:2007:EOA

[457] S. K. Garg. Effect of oral administration of l-thyroxine (T_{4}) on growth performance, digestibility, and nutrient retention in Channa punctatus (Bloch) and Heteropneustes fossilis (Bloch). Fish Physiology and Biochemistry, 33(4):347–358, December 2007. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-007-9166-1.

Nath:2007:VSE

[458] P. Nath, R. Sahu, and D. Bhattacharya. Vitellogenesis with special emphasis on Indian fishes. Fish Physiology and Biochemistry, 33(4):359–366, December 2007. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
007-9167-0.

Maitra:2007:RPM

[459] S. K. Maitra and A. Chattoraj. Role of photoperiod and melatonin in the regulation of ovarian functions in Indian carp Catla catla: basic information for future application. Fish Physiology and Biochemistry, 33 (4):367–382, December 2007. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-007-9174-1.

Chowdhury:2007:SVR

[460] I. Chowdhury and K. P. Joy. Seminal vesicle and its role in the reproduction of teleosts. Fish Physiology and Biochemistry, 33(4):383–398, December 2007. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
007-9162-5.

Munilkumar:2007:APN

[461] S. Munilkumar and M. C. Nandeesha. Aquaculture practices in Northeast India: Current status and future directions. Fish Physiology and Biochemistry, 33(4):399–412, December 2007. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-007-9163-4.

Routray:2007:RAC

[462] P. Routray, D. K. Verma, and N. Sarangi. Recent advances in carp seed production and milt cryopreservation. Fish Physiology and Biochemistry, 33(4):413–427, December 2007. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-007-9159-0.

Panigrahi:2007:MIB

[463] A. Panigrahi and I. S. Azad. Microbial intervention for better fish health in aquaculture: the Indian scenario. Fish Physiology and Biochemistry, 33 (4):429–440, December 2007. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-007-9160-7.

Banerjee:2007:HRO

[464] T. K. Banerjee. Histopathology of respiratory organs of certain airbreathing fishes of India. Fish Physiology and Biochemistry, 33(4): 441–454, December 2007. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-007-9170-5.

Lal:2007:PIR

[465] B. Lal. Pesticide — induced reproductive dysfunction in Indian fishes. Fish Physiology and Biochemistry, 33(4):455–462, December 2007. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-007-9171-4.

Bhattacharya:2007:AIR

[466] S. Bhattacharya, A. Bhattacharya, and S. Roy. Arsenic-induced responses in freshwater teleosts. Fish Physiology and Biochemistry, 33(4): 463–473, December 2007. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-007-9173-2.

Lakra:2007:FGC

[467] W. S. Lakra, Vindhya Mohindra, and Kuldeep K. Lal. Fish genetics and conservation research in India: status and perspectives. Fish Physiology and Biochemistry, 33(4):475–487, December 2007. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-007-9168-z.

Stekhoven:2008:IPK

[468] F. M. A. H. Schuurmans Stekhoven, M. H. A. G. Gorissen, and G. Flik. The isoelectric point, a key to understanding a variety of biochemical problems: a minireview. Fish Physiology and Biochemistry, 34(1):1–8, March 2008. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
007-9145-6.

Niu:2008:EDP

[469] J. Niu, Y. J. Liu, and Y. Zhu. Effects of dietary phospholipid level in cobia (Rachycentron canadum) larvae: growth, survival, plasma lipids and enzymes of lipid metabolism. Fish Physiology and Biochemistry, 34(1):9–17, March 2008. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
007-9140-y.

Seear:2008:SRI

[470] P. J. Seear and G. E. Sweeney. Stability of RNA isolated from postmortem tissues of Atlantic salmon (Salmo salar L.). Fish Physiology and Biochemistry, 34(1):19–24, March 2008. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-007-9141-x.

Singh:2008:BHD

[471] Pratap B. Singh and Vandana Singh. Bioaccumulation of hexachlorocyclohexane, dichlorodiphenyltrichloroethane, and estradiol-17β in catfish and carp during the pre-monsoon season in India. Fish Physiology and Biochemistry, 34(1):25–36, March 2008. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-007-9142-9.

Chen:2008:SMB

[472] Heng Chen and Rui Mu. Seasonal morphological and biochemical changes of dahlgren cells implies a potential role of the caudal neurosecretory system (CNSS) in the reproduction cycle of teleostean fish. Fish Physiology and Biochemistry, 34(1):37–42, March 2008. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-007-9143-8.

Wu:2008:MCD

[473] Cheng Wu, Ling Wang, and Fashui Hong. Mechanism of cd^{2+} on DNA cleavage and ca^{2+} on DNA repair in liver of silver crucian carp. Fish Physiology and Biochemistry, 34(1):43–51, March 2008. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-007-9144-7.

Ates:2008:ESS

[474] Burhan Ates, Ibrahim Orun, and Ismet Yilmaz. Effects of sodium selenite on some biochemical and hematological parameters of rainbow trout (Oncorhynchus mykiss Walbaum, 1792) exposed to Pb^{2+} and Cu^{2+}. Fish Physiology and Biochemistry, 34(1):53–59, March 2008. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-007-9146-5.

Asagba:2008:BCB

[475] Samuel O. Asagba, George E. Eriyamremu, and Mabel E. Igberaese. Bioaccumulation of cadmium and its biochemical effect on selected tissues of the catfish (Clarias gariepinus). Fish Physiology and Biochemistry, 34(1):61–69, March 2008. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-007-9147-4.

Sahinoz:2008:DSP

[476] Erdinç Şahinöz, Faruk Aral, and Zafer Doğu. Determination of spermatological properties of male Liza abu (Heckel, 1843) in Atatürk Dam Lake, Şanlıurfa. Fish Physiology and Biochemistry, 34(1):71–76, March 2008. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-007-9148-3.

Garcia:2008:GLC

[477] Alexandre Sachsida Garcia, Christopher C. Parrish, and Joseph A. Brown. Growth and lipid composition of Atlantic cod (Gadus morhua) larvae in response to differently enriched Artemia franciscana. Fish Physiology and Biochemistry, 34(1):77–94, March 2008. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-007-9149-2.

Sink:2008:VUD

[478] T. D. Sink, R. T. Lochmann, and K. A. Fecteau. Validation, use, and disadvantages of enzyme-linked immunosorbent assay kits for detection of cortisol in channel catfish, largemouth bass, red pacu, and golden shiners. Fish Physiology and Biochemistry, 34(1):95–101, March 2008. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-007-9150-9.

Binner:2008:EAJ

[479] Maaike Binner, Werner Kloas, and Iris Hardewig. Energy allocation in juvenile roach and burbot under different temperature and feeding regimes. Fish Physiology and Biochemistry, 34(2):103–116, June 2008. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-007-9151-8.

Izquierdo:2008:RGF

[480] M. S. Izquierdo, Lidia Robaina, and Juan Manuel Afonso. Regulation of growth, fatty acid composition and delta 6 desaturase expression by dietary lipids in gilthead seabream larvae (Sparus aurata). Fish Physiology and Biochemistry, 34(2):117–127, June 2008. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-007-9152-7.

Aliza:2008:IWH

[481] Dwinna Aliza, Ida Shazrina Ismail, and Tengku Sifzizul Tengku Muhammad. Identification of Wap65, a human homologue of hemopexin as a copper-inducible gene in swordtail fish, Xiphophorus helleri. Fish Physiology and Biochemistry, 34(2):129–138, June 2008. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-007-9153-6.

Monzani:2008:UCE

[482] Paulo Sérgio Monzani and Gilberto Moraes. Urea cycle enzymes through the development of pacu (Piaractus mesopotamicus): the role of ornithine carbamoyl transferase. Fish Physiology and Biochemistry, 34(2):139–149, June 2008. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
007-9154-5.

Kalaimani:2008:AOS

[483] N. Kalaimani, N. Chakravarthy, and T. C. Santiago. Anti-oxidant status in embryonic, post-hatch and larval stages of Asian seabass (Lates calcarifer). Fish Physiology and Biochemistry, 34(2):151–158, June 2008. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-007-9155-4.

Valenzuela:2008:EDA

[484] Ariel E. Valenzuela, Victor M. Silva, and Alfredo E. Klempau. Effects of different artificial photoperiods and temperatures on haematological parameters of rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 34(2):159–167, June 2008. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-007-9156-3.

Remyla:2008:IZC

[485] Sweety R. Remyla, Mathan Ramesh, and Kurunthachalam Senthil Kumar. Influence of zinc on cadmium induced haematological and biochemical responses in a freshwater teleost fish Catla catla. Fish Physiology and Biochemistry, 34(2):169–174, June 2008. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-007-9157-2.

Du:2008:ISD

[486] X. Z. Du, L. Zhou, and J. F. Gui. Identical sequences but different expression patterns of Hira gene in gynogenetic and gonochoristic crucian carps. Fish Physiology and Biochemistry, 34(2):175–184, June 2008. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-007-9158-1.

An:2008:CEA

[487] Kwang Wook An, Na Na Kim, and Cheol Young Choi. Cloning and expression of aquaporin 1 and arginine vasotocin receptor mRNA from the black porgy, Acanthopagrus schlegeli: effect of freshwater acclimation. Fish Physiology and Biochemistry, 34(2):185–194, June 2008. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-007-9175-0.

Dorafshan:2008:ETC

[488] Salar Dorafshan, Mohammad Reza Kalbassi, and Sahel Soltan Karimi. Effects of triploidy on the Caspian salmon Salmo trutta caspius haematology. Fish Physiology and Biochemistry, 34(3):195–200, September 2008. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-007-9176-z.

Bouraoui:2008:AEC

[489] Z. Bouraoui, M. Banni, and H. Boussetta. Acute effects of cadmium on liver phase I and phase II enzymes and metallothionein accumulation on sea bream Sparus aurata. Fish Physiology and Biochemistry, 34(3): 201–207, September 2008. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-007-9177-y.

Li:2008:ATP

[490] Huixian Li, Hui Jiang, and Jiao Chen. Acute toxicity of the pesticide methomyl on the topmouth gudgeon (Pseudorasbora parva): mortality and effects on four biomarkers. Fish Physiology and Biochemistry, 34 (3):209–216, September 2008. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-007-9178-x.

Talas:2008:ARS

[491] Zeliha Selamoglu Talas, Ibrahim Orun, and Ismet Yılmaz. Antioxidative role of selenium against the toxic effect of heavy metals (Cd^{+2}, Cr^{+3}) on liver of rainbow trout (Oncorhynchus mykiss Walbaum 1792). Fish Physiology and Biochemistry, 34(3):217–222, September 2008. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-007-9179-9.

Polednik:2008:PRW

[492] Lukáš Poledník, Jiří Řehulka, and Hana Kazihnitková. Physiological responses of over-wintering common carp (Cyprinus carpio) to disturbance by Eurasian otter (Lutra lutra). Fish Physiology and Biochemistry, 34 (3):223–234, September 2008. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-007-9180-3.

Shen:2008:TEP

[493] Xue-Yan Shen, Jian-Zhou Cui, and Yoshitaka Nagahama. Transcript expression profiles of Takifugu rubripes spermatozoa and eggs by expressed sequence tag analysis. Fish Physiology and Biochemistry, 34(3): 235–243, September 2008. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-007-9182-1.

Fischer:2008:PFD

[494] Claes Fischer and John Fleng Steffensen. Plasma FITC-dextran exchange between the primary and secondary circulatory systems in the Atlantic cod, Gadus morhua. Fish Physiology and Biochemistry, 34(3): 245–249, September 2008. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-007-9183-0.

Tewary:2008:UVC

[495] Arup Tewary and Bidhan C. Patra. Use of vitamin C as an immunostimulant. Effect on growth, nutritional quality, and immune response of Labeo rohita (Ham.). Fish Physiology and Biochemistry, 34(3):251–259, September 2008. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
007-9184-z.

Brodte:2008:TDL

[496] Eva Brodte, M. Graeve, and H.-O. Pörtner. Temperature-dependent lipid levels and components in polar and temperate eelpout (Zoarcidae). Fish Physiology and Biochemistry, 34(3):261–274, September 2008. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-007-9185-y.

Recks:2008:VFA

[497] Melissa A. Recks and Gloria T. Seaborn. Variation in fatty acid composition among nine forage species from a southeastern US estuarine and nearshore coastal ecosystem. Fish Physiology and Biochemistry, 34(3): 275–287, September 2008. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-007-9186-x.

Vikingstad:2008:CET

[498] Erik Vikingstad, Eva Andersson, and Geir Lasse Taranger. The combined effects of temperature and GnRHa treatment on the final stages of sexual maturation in Atlantic salmon (Salmo salar L.) females. Fish Physiology and Biochemistry, 34(3):289–298, September 2008. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-007-9187-9.

Sukumaran:2008:HBR

[499] K. Sukumaran, A. K. Pal, and D. Debnath. Haemato-biochemical responses and induction of HSP70 to dietary phosphorus in Catla catla (Hamilton) fingerlings. Fish Physiology and Biochemistry, 34(4):299– 306, December 2008. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-007-9188-8.

Zhang:2008:MRI

[500] H. Zhang, J. Zhang, and Y. Zhu. Microcystin-RR induces apoptosis in fish lymphocytes by generating reactive oxygen species and causing mitochondrial damage. Fish Physiology and Biochemistry, 34(4):307–312, December 2008. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
007-9189-7.

Dussault:2008:ESW

[501] È. B. Dussault, R. C. Playle, and R. S. McKinley. Effects of soft-water acclimation on the physiology, swimming performance, and cardiac parameters of the rainbow trout, Oncorhynchus mykiss. Fish Physiology and Biochemistry, 34(4):313–322, December 2008. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-007-9190-1.

Zhang:2008:PBR

[502] Xuezhen Zhang, Ping Xie, and Zechao Shi. Plasma biochemical responses of the omnivorous crucian carp (Carassius auratus) to crude cyanobacterial extracts. Fish Physiology and Biochemistry, 34(4):323–329, December 2008. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
007-9191-0.

Kumar:2008:CLC

[503] Amit Kumar, Bechan Sharma, and Ravi Shankar Pandey. Cypermethrin and λ-cyhalothrin induced alterations in nucleic acids and protein contents in a freshwater fish, Channa punctatus. Fish Physiology and Biochemistry, 34(4):331–338, December 2008. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-007-9192-z.

Wang:2008:MCE

[504] Y. Wang, L. Zhou, and J. F. Gui. Molecular cloning and expression characterization of ApoC-I in the orange-spotted grouper. Fish Physiology and Biochemistry, 34(4):339–348, December 2008. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-007-9193-y.

Mommens:2008:SPP

[505] Maren Mommens, Mariola Wojtczak, and Igor Babiak. Seminal plasma proteins of Atlantic halibut (Hippoglossus hippoglossus L.). Fish Physiology and Biochemistry, 34(4):349–355, December 2008. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-007-9194-x.

Pham:2008:CII

[506] Ky Xuan Pham, Masafumi Amano, and Kunio Yamamori. Changes in the immunostaining intensities of follicle-stimulating hormone and luteinizing hormone during ovarian maturation in the female Japanese flounder. Fish Physiology and Biochemistry, 34(4):357–365, December 2008. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-007-9195-9.

Ciba:2008:LTC

[507] P. Ciba, S. Schicktanz, and C. Kruse. Long-term culture of a cell population from Siberian sturgeon (Acipenser baerii) head kidney. Fish Physiology and Biochemistry, 34(4):367–372, December 2008. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-007-9196-8.

Alvarez-Gonzalez:2008:DDE

[508] C. A. Alvarez-González, F. J. Moyano-López, and S. Dumas. Development of digestive enzyme activity in larvae of spotted sand bass Paralabrax maculatofasciatus. 1. Biochemical analysis. Fish Physiology and Biochemistry, 34(4):373–384, December 2008. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-007-9197-7.

Prasad:2008:MDG

[509] G. Prasad and P. H. Anvar Ali. Morphology of the diet in the gut of threatened yellow catfish Horabagrus brachysoma (Gunther 1864) at two life stages. Fish Physiology and Biochemistry, 34(4):385–389, December 2008. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-007-9198-6.

Naz:2008:CBC

[510] Mehmet Naz. The changes in the biochemical compositions and enzymatic activities of rotifer (Brachionus plicatilis, Müller) and Artemia during the enrichment and starvation periods. Fish Physiology and Biochemistry, 34 (4):391–404, December 2008. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-007-9199-5.

Baldisserotto:2008:NIF

[511] Bernardo Baldisserotto, Carlos E. Copatti, and Rodrigo Roubach. Net ion fluxes in the facultative air-breather Hoplosternum littorale (tamoata) and the obligate air-breather Arapaima gigas (pirarucu) exposed to different Amazonian waters. Fish Physiology and Biochemistry, 34(4):405–412, December 2008. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
008-9200-y.

Rajesh:2008:GHE

[512] R. Rajesh and K. C. Majumdar. The growth hormone-encoding gene isolated and characterized from Labeo rohita Hamilton is expressed in CHO cells under the control of constitutive promoters in ‘autotransgene’ constructs. Fish Physiology and Biochemistry, 34(4):413–436, December 2008. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-008-9201-x.

Henningsen:2008:SLR

[513] A. D. Henningsen, F. L. Murru, and G. C. Violetta. Serum levels of reproductive steroid hormones in captive sand tiger sharks, Carcharias taurus (Rafinesque), and comments on their relation to sexual conflicts. Fish Physiology and Biochemistry, 34(4):437–446, December 2008. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-008-9202-9.

Unal:2008:DHI

[514] G. Ünal, E. Erdoğan, and M. Elp. Determination of hormones inducing oocyte maturation in Chalcalburnus tarichi (Pallas, 1811). Fish Physiology and Biochemistry, 34(4):447–454, December 2008. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-008-9203-8.

Uysal:2008:CSC

[515] K. Uysal, M. Bülbül, and A. K. Seçkin. Changes in some components of the muscle lipids of three freshwater fish species under natural extreme cold and temperate conditions. Fish Physiology and Biochemistry, 34 (4):455–463, December 2008. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-008-9220-7.

Luo:2008:SDA

[516] Yiping Luo and Xiaojun Xie. Specific dynamic action in two body size groups of the southern catfish (Silurus meridionalis) fed diets differing in carbohydrate and lipid contents. Fish Physiology and Biochemistry, 34 (4):465–471, December 2008. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-008-9221-6.

Gharaei:2008:BEH

[517] A. Gharaei, A. Esmaili-Sari, and M. Ghaffari. Beluga (Huso huso, brandet 1869) bioenergetics under dietary methylmercury. Fish Physiology and Biochemistry, 34(4):473–482, December 2008. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-008-9224-3.

DeBusk:2008:SDE

[518] B. C. DeBusk, M. Slattery, and Daniel Schlenk. Species differences and effects of soft coral extracts from Sinnularia maximus on the expression of cytochrome P4501A and 2N in butterflyfishes (Chaetodon spp.). Fish Physiology and Biochemistry, 34(4):483–492, December 2008. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-008-9225-2.

Viveiros:2009:FSI

[519] Ana T. M. Viveiros. Foreword for the special issue Fish Research in Latin America: Biodiversity, and Basic and Applied Physiology. Fish Physiology and Biochemistry, 35(1):1, March 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-008-9246-x.

Borella:2009:IIA

[520] M. I. Borella, R. Venturieri, and J. M. Mancera. Immunocytochemical identification of adenohypophyseal cells in the pirarucu (Arapaima gigas), an Amazonian basal teleost. Fish Physiology and Biochemistry, 35(1):3–16, March 2009. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-008-9254-x.

Cerqueira:2009:RSI

[521] V. R. Cerqueira and M. Y. Tsuzuki. A review of spawning induction, larviculture, and juvenile rearing of the fat snook, Centropomus parallelus. Fish Physiology and Biochemistry, 35(1):17–28, March 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-008-9245-y.

Cussac:2009:FSS

[522] V. E. Cussac, D. A. Fernández, and H. L. López. Fishes of southern South America: a story driven by temperature. Fish Physiology and Biochemistry, 35(1):29–42, March 2009. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-008-9217-2.

Fernandez:2009:FME

[523] Daniel A. Fernández and Jorge Calvo. Fish muscle: the exceptional case of notothenioids. Fish Physiology and Biochemistry, 35(1):43–52, March 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
008-9282-6.

Garcia:2009:PCF

[524] Aurea Garcia, Salvador Tello, and Fabrice Duponchelle. Patterns of commercial fish landings in the Loreto region (Peruvian Amazon) between 1984 and 2006. Fish Physiology and Biochemistry, 35(1):53–67, March 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
008-9212-7.

Garcia-Ortega:2009:NFR

[525] Armando García-Ortega. Nutrition and feeding research in the spotted rose snapper (Lutjanus guttatus) and bullseye puffer (Sphoeroides annulatus), new species for marine aquaculture. Fish Physiology and Biochemistry, 35(1):69–80, March 2009. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-008-9226-1.

Oliveira:2009:GNF

[526] C. Oliveira, F. Foresti, and A. W. S. Hilsdorf. Genetics of neotropical fish: from chromosomes to populations. Fish Physiology and Biochemistry, 35(1):81–100, March 2009. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-008-9250-1.

Miranda:2009:ELT

[527] L. A. Miranda, C. A. Strüssmann, and G. M. Somoza. Effects of light and temperature conditions on the expression of GnRH and GtH genes and levels of plasma steroids in Odontesthes bonariensis females. Fish Physiology and Biochemistry, 35(1):101–108, March 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-008-9232-3.

Honji:2009:POD

[528] R. M. Honji, A. M. Narcizo, and R. G. Moreira. Patterns of oocyte development in natural habitat and captive Salminus hilarii Valenciennes, 1850 (Teleostei: Characidae). Fish Physiology and Biochemistry, 35(1):109–123, March 2009. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-008-9239-9.

Chu-Koo:2009:GDP

[529] F. Chu-Koo, R. Dugué, and J. Nuñez. Gender determination in the paiche or pirarucu (Arapaima gigas) using plasma vitellogenin, 17β-estradiol, and 11-ketotestosterone levels. Fish Physiology and Biochemistry, 35 (1):125–136, March 2009. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-008-9211-8.

Viveiros:2009:SQC

[530] A. T. M. Viveiros and H. P. Godinho. Sperm quality and cryopreservation of Brazilian freshwater fish species: a review. Fish Physiology and Biochemistry, 35(1):137–150, March 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-008-9240-3.

Zaniboni-Filho:2009:WPE

[531] E. Zaniboni-Filho, A. P. O. Nuñer, and R. L. Serafini. Water pH and Prochilodus lineatus larvae survival. Fish Physiology and Biochemistry, 35(1):151–155, March 2009. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-008-9235-0.

Guilgur:2009:MEG

[532] L. G. Guilgur, C. A. Strüssmann, and G. M. Somoza. mRNA expression of GnRH variants and receptors in the brain, pituitary and ovaries of pejerrey (Odontesthes bonariensis) in relation to the reproductive status. Fish Physiology and Biochemistry, 35(1):157–166, March 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-008-9215-4.

Nunez:2009:TUS

[533] Jesús Núñez and Fabrice Duponchelle. Towards a universal scale to assess sexual maturation and related life history traits in oviparous teleost fishes. Fish Physiology and Biochemistry, 35(1):167–180, March 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-008-9241-2.

Fernandez:2009:EDS

[534] D. A. Fernández, M. E. Lattuca, and J. Calvo. Energy density of sub-Antarctic fishes from the Beagle Channel. Fish Physiology and Biochemistry, 35(1):181–188, March 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-008-9234-1.

Guerrero:2009:RBF

[535] H. Y. Guerrero, E. Cardillo, and D. Marcano. Reproductive biology of freshwater fishes from the Venezuelan floodplains. Fish Physiology and Biochemistry, 35(1):189–196, March 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-008-9249-7.

Nobrega:2009:OFS

[536] R. H. Nóbrega, S. R. Batlouni, and L. R. França. An overview of functional and stereological evaluation of spermatogenesis and germ cell transplantation in fish. Fish Physiology and Biochemistry, 35(1):197–206, March 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-008-9252-z.

Batlouni:2009:CJF

[537] S. R. Batlouni, R. H. Nóbrega, and L. R. França. Cell junctions in fish seminiferous epithelium. Fish Physiology and Biochemistry, 35(1):207–217, March 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
008-9253-y.

Martins:2009:LRP

[538] Maurício L. Martins, Felipe N. Vieira, and Gilberto Pereira, Jr. Leukocyte response and phagocytic activity in Nile tilapia experimentally infected with Enterococcus sp. Fish Physiology and Biochemistry, 35(1):219–222, March 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
008-9262-x.

He:2009:OJC

[539] Xugang He, Ping Zhuang, and Congxin Xie. Osmoregulation in juvenile Chinese sturgeon (Acipenser sinensis gray) during brackish water adaptation. Fish Physiology and Biochemistry, 35(2):223–230, May 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-008-9230-5.

Mickowska:2009:PCP

[540] B. Mickowska. Purification and characterization of α_{1}-proteinase inhibitor and antithrombin III: major serpins of rainbow trout (Oncorhynchuss mykiss) and carp (Cyprinus carpio) blood plasma. Fish Physiology and Biochemistry, 35(2):231–240, May 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-008-9204-7.

Maltais:2009:PPC

[541] D. Maltais and R. L. Roy. Purification and partial characterization of vitellogenin from shorthead redhorse (Moxostoma macrolepidotum) and copper redhorse (Moxostoma hubbsi) and detection in plasma and mucus with a heterologous antibody. Fish Physiology and Biochemistry, 35(2):241–254, May 2009. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-008-9205-6.

Rungruangsak-Torrissen:2009:DEF

[542] Krisna Rungruangsak-Torrissen, Jan Sunde, and Frode Oppedal. Digestive efficiency, free amino acid pools and quality of growth performance in Atlantic salmon (Salmo salar L.) affected by light regimes and vaccine types. Fish Physiology and Biochemistry, 35(2):255–272, May 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-008-9206-5.

Molich:2009:TPR

[543] Andreas Mölich, Wolfgang Waser, and Norbert Heisler. The teleost pseudobranch: a role for preconditioning of ocular blood supply? Fish Physiology and Biochemistry, 35(2):273–286, May 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-008-9207-4.

Strunjak-Perovic:2009:SNA

[544] Ivancica Strunjak-Perovic, Rozelindra Coz-Rakovac, and Margita Jadan. Seasonality of nuclear abnormalities in gilthead sea bream Sparus aurata (L.) erythrocytes. Fish Physiology and Biochemistry, 35(2):287–291, May 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
008-9208-3.

Banni:2009:AEB

[545] M. Banni, Z. Bouraoui, and H. Boussetta. Acute effects of benzo[a]pyrene on liver phase I and II enzymes, and DNA damage on sea bream Sparus aurata. Fish Physiology and Biochemistry, 35(2):293–299, May 2009. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-008-9210-9.

Kumar:2009:MKM

[546] Shivendra Kumar, N. P. Sahu, and Kartik Baruah. Modulation of key metabolic enzyme of Labeo rohita (Hamilton) juvenile: effect of dietary starch type, protein level and exogenous α-amylase in the diet. Fish Physiology and Biochemistry, 35(2):301–315, May 2009. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-008-9213-6.

Devincenti:2009:PFE

[547] C. V. Devincenti, A. O. Díaz, and A. L. Goldemberg. Pectoral fins of Micropogonias furnieri: a histochemical and ultrastructural study. Fish Physiology and Biochemistry, 35(3):317–323, August 2009. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-008-9216-3.

Chen:2009:EPE

[548] W. Chen, M. Cao, and H. Zhao. Expression pattern of prmt5 in adult fish and embryos of medaka, Oryzias latipes. Fish Physiology and Biochemistry, 35(3):325–332, August 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-008-9233-2.

He:2009:PWP

[549] Feng He, Hai Shen Wen, and Yu Guo Zhou. Polymorphisms within promoter of Japanese flounder (Paralichthys olivaceus) ovary cytochrome P450-c19 (CYP19a) gene associated with reproductive traits. Fish Physiology and Biochemistry, 35(3):333–340, August 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/s10695-008-9237-y.

Liao:2009:CRR

[550] T. Liao, Q. L. Guo, and Y. Xu. Comparative responses in rare minnow exposed to 17β-estradiol during different life stages. Fish Physiology and Biochemistry, 35(3):341–349, August 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-008-9247-9.

Xu:2009:EPX

[551] Baohua Xu, Yanbo Wang, and Qiang Lin. Effect of prebiotic xylooligosaccharides on growth performances and digestive enzyme activities of allogynogenetic crucian carp (Carassius auratus gibelio). Fish Physiology and Biochemistry, 35(3):351–357, August 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-008-9248-8.

Freijo:2009:MHC

[552] R. O. Freijo, A. M. García, and A. O. Díaz. Morphological and histochemical characteristics of the epithelium of ovarian lamellae of Genypterus blacodes (Schneider, 1801). Fish Physiology and Biochemistry, 35(3):359–367, August 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
008-9256-8.

Shahbazi:2009:CCT

[553] Fatemeh Shahbazi, Susanne Holmgren, and Jörgen Jensen. Cod CGRP and tachykinins in coeliac artery innervation of the Atlantic cod, Gadus morhua: presence and vasoactivity. Fish Physiology and Biochemistry, 35(3):369–376, August 2009. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-008-9257-7.

Green:2009:EEM

[554] C. C. Green and A. M. Kelly. Effects of the estrogen mimic genistein as a dietary component on sex differentiation and ethoxyresorufin-o-deethylase (EROD) activity in channel catfish (Ictalurus punctatus). Fish Physiology and Biochemistry, 35(3):377–384, August 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-008-9260-z.

Shan:2009:ODD

[555] Xiu-Juan Shan, Wei Huang, and Shuo-Zeng Dou. Ontogenetic development of digestive enzymes and effect of starvation in miiuy croaker Miichthys miiuy larvae. Fish Physiology and Biochemistry, 35(3):385–398, August 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
008-9263-9.

Noffs:2009:DFO

[556] M. D. Noffs, R. C. Martino, and L. S. Takahashi. Dietary fish oil replacement with lard and soybean oil affects triacylglycerol and phospholipid muscle and liver docosahexaenoic acid content but not in the brain and eyes of surubim juveniles Pseudoplatystoma sp. Fish Physiology and Biochemistry, 35(3):399–412, August 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-008-9264-8.

Manju:2009:CAI

[557] Maniyan Manju, Thomas G. Sherin, and Oommen Vilaverthottathil Oommen. Curcumin analogue inhibits lipid peroxidation in a freshwater teleost, Anabas testudineus (Bloch) — an in vitro and in vivo study. Fish Physiology and Biochemistry, 35(3):413–420, August 2009. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-008-9266-6.

Wen:2009:EPE

[558] Aiyun Wen, Feng You, and Peijun Zhang. Expression pattern of dmrt4 from olive flounder (Paralichthys olivaceus) in adult gonads and during embryogenesis. Fish Physiology and Biochemistry, 35(3):421–433, August 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
008-9267-5.

Zhou:2009:CHS

[559] Xiaoyun Zhou, Mingyun Li, and Weimin Wang. Comparison of haematology and serum biochemistry of cultured and wild dojo loach Misgurnus anguillicaudatus. Fish Physiology and Biochemistry, 35(3):435–441, August 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
008-9268-4.

Lushchak:2009:ADG

[560] Volodymyr I. Lushchak, Viktor V. Husak, and Kenneth B. Storey. AMPdeaminase from goldfish white muscle: regulatory properties and redistribution under exposure to high environmental oxygen level. Fish Physiology and Biochemistry, 35(3):443–452, August 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-008-9270-x.

Gesto:2009:ABM

[561] Manuel Gesto, Adrián Tintos, and Jesús M. Míguez. Alterations in the brain monoaminergic neurotransmitters of rainbow trout related to naphthalene exposure at the beginning of vitellogenesis. Fish Physiology and Biochemistry, 35(3):453–465, August 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-008-9271-9.

Bandyopadhyay:2009:EPB

[562] Partha Bandyopadhyay and Pradeep K. Das Mohapatra. Effect of a probiotic bacterium Bacillus circulans PB7 in the formulated diets: on growth, nutritional quality and immunity of Catla catla (Ham.). Fish Physiology and Biochemistry, 35(3):467–478, August 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-008-9272-8.

Siddiqui:2009:EDP

[563] Tabassum Qamar Siddiqui and Mukhtar A. Khan. Effects of dietary protein levels on growth, feed utilization, protein retention efficiency and body composition of young Heteropneustes fossilis (Bloch). Fish Physiology and Biochemistry, 35(3):479–488, August 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-008-9273-7.

Chen:2009:CCE

[564] Chao Chen, Zhongsheng Li, and Jianguo He. Cloning, characterization and expression analysis of a CXCR1-like gene from mandarin fish Siniperca chuatsi. Fish Physiology and Biochemistry, 35(3):489–499, August 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-008-9283-5.

Meier:2009:IGH

[565] Karina M. Meier, Márcio A. Figueiredo, and Luis F. Marins. Increased growth hormone (GH), growth hormone receptor (GHR), and insulin-like growth factor I (IGF-I) gene transcription after hyperosmotic stress in the Brazilian flounder Paralichthys orbignyanus. Fish Physiology and Biochemistry, 35(3):501–509, August 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-008-9287-1.

Liu:2009:APH

[566] Ji-Fang Liu, Yann Guiguen, and Shao-Jun Liu. Aromatase (P450arom) and 11β-hydroxylase (P45011β) genes are differentially expressed during the sex change process of the protogynous rice field eel, Monopterus albus. Fish Physiology and Biochemistry, 35(3):511–518, August 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-008-9255-9.

Enes:2009:NRH

[567] P. Enes, S. Panserat, and A. Oliva-Teles. Nutritional regulation of hepatic glucose metabolism in fish. Fish Physiology and Biochemistry, 35 (3):519–539, August 2009. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-008-9259-5.

Khemiss:2009:EVE

[568] Fathia Khemiss, Salwa Ahmadi, and Dalila Saïdane. Effect of in vitro exposure to Vibrio vulnificus on hydroelectrolytic transport and structural changes of sea bream (Sparus aurata L.) intestine. Fish Physiology and Biochemistry, 35(3):541–549, August 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-008-9265-7.

Linhart:2009:FSI

[569] Otomar Linhart. Foreword for the special issue Research on Fish Physiology and Biochemistry in Central Europe: From Basic Sciences to Applied Approaches. Fish Physiology and Biochemistry, 35(4):??, November 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-009-9343-5.

Duran:2009:BCS

[570] Ayhan Duran and Zeliha Selamoglu Talas. Biochemical changes and sensory assessment on tissues of carp (Cyprinus carpio, Linnaeus 1758) during sale conditions. Fish Physiology and Biochemistry, 35(4):??, November 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
009-9331-9.

Imanpoor:2009:RBB

[571] Mohammad Reza Imanpoor, Sayed Hadi Alavi, and Jacky Cosson. Relationship between biological characteristics of eggs and female brood stocks of Persian sturgeon Acipenser persicus: a comparison with other fish species. Fish Physiology and Biochemistry, 35(4):??, November 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/s10695-009-9330-x.

Jeney:2009:DAR

[572] Zsigmond Jeney, Tímea Rácz, and Galina Jeney. Differences in the antibody response and survival of genetically different varieties of common carp (Cyprinus carpio L.) vaccinated with a commercial Aeromonas salmonicida/A. hydrophila vaccine and challenged with A. hydrophila. Fish Physiology and Biochemistry, 35(4):??, November 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-009-9329-3.

Samuiloviene:2009:GDD

[573] Aurelija Samuiloviene, Antanas Kontautas, and Riho Gross. Genetic diversity and differentiation of sea trout (Salmo trutta) populations in Lithuanian rivers assessed by microsatellite DNA variation. Fish Physiology and Biochemistry, 35(4):??, November 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-009-9310-1.

Coz-Rakovac:2009:CAA

[574] R. Coz-Rakovac, N. Topic Popovic, and M. Jadan. Classification accuracy of algorithms for blood chemistry data for three aquaculture-affected marine fish species. Fish Physiology and Biochemistry, 35(4):??, November 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
008-9288-0.

Mraz:2009:DBL

[575] J. Mráz and J. Pickova. Differences between lipid content and composition of different parts of fillets from crossbred farmed carp (Cyprinus carpio). Fish Physiology and Biochemistry, 35(4):??, November 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-008-9291-5.

Sudova:2009:EPA

[576] Eliška Sudová, Veronika Piačková, and Zdeňka Svobodová. The effect of praziquantel applied per os on selected haematological and biochemical indices in common carp (Cyprinus carpio L.). Fish Physiology and Biochemistry, 35(4):??, November 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-008-9269-3.

Psenicka:2009:SAC

[577] Martin Pšenička, Marek Rodina, and Otomar Linhart. Structural abnormalities of common carp Cyprinus carpio spermatozoa. Fish Physiology and Biochemistry, 35(4):??, November 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/s10695-008-9285-3.

Sales:2009:PDE

[578] James Sales. Prediction of digestible energy content across feed ingredients and fish species by linear regression. Fish Physiology and Biochemistry, 35 (4):551–565, November 2009. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-008-9286-2.

Li:2009:SPT

[579] Ping Li, Martin Hulak, and Otomar Linhart. Sperm proteins in teleostean and chondrostean (sturgeon) fishes. Fish Physiology and Biochemistry, 35 (4):567–581, November 2009. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-008-9261-y.

Velisek:2009:EBS

[580] J. Velisek, Z. Svobodova, and J. Machova. Effects of bifenthrin on some haematological, biochemical and histopathological parameters of common carp (Cyprinus carpio L.). Fish Physiology and Biochemistry, 35 (4):583–590, November 2009. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-008-9258-6.

Zietara:2009:BFS

[581] M. S. Ziȩtara, A. Biegniewska, and E. F. Skorkowski. Bioenergetics of fish spermatozoa during semen storage. Fish Physiology and Biochemistry, 35 (4):607–614, November 2009. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-009-9308-8.

Jezierska:2009:EHM

[582] Barbara Jezierska, Katarzyna Ługowska, and Małgorzata Witeska. The effects of heavy metals on embryonic development of fish (a review). Fish Physiology and Biochemistry, 35(4):625–640, November 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/s10695-008-9284-4.

Boryshpolets:2009:PSW

[583] Sergey Boryshpolets, Boris Dzyuba, and Sergey Drokin. Pre-spawning water temperature affects sperm respiration and reactivation parameters in male carps. Fish Physiology and Biochemistry, 35(4):661–668, November 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (elec-

tronic). URL https://link.springer.com/article/10.1007/s10695-
008-9292-4.

Galina:2009:UIH

[584] Jeney Galina, G. Yin, and Z. Jeney. The use of immunostimulating herbs in fish. An overview of research. Fish Physiology and Biochemistry, 35 (4):669–676, November 2009. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-009-9304-z.

Alavi:2009:SQM

[585] Sayyed Mohammad Hadi Alavi, Martin Pšenička, and Otomar Linhart. Sperm quality in male Barbus barbus L. fed different diets during the spawning season. Fish Physiology and Biochemistry, 35(4): 683–693, November 2009. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/s10695-009-9325-7.

Hatef:2009:COF

[586] Azadeh Hatef, Hamid Niksirat, and Sayyed Mohammad Hadi Alavi. Composition of ovarian fluid in endangered Caspian brown trout, Salmo trutta caspius, and its effects on spermatozoa motility and fertilizing ability compared to freshwater and a saline medium. Fish Physiology and Biochemistry, 35(4):695–700, November 2009. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/s10695-008-9302-6.